### v2

# Amplifying Main Memory-Based Timing Covert and Side Channels using Processing-in-Memory Operations

Abstract
The adoption of processing-in-memory (PiM) architectures has been gaining momentum because they provide high performance and low energy consumption by alleviating the data movement bottleneck. Yet, the security of such architectures has not been thoroughly explored. The adoption of PiM solutions provides a new way to directly access main memory, which can be potentially exploited by malicious user applications. We show that this new way to access main memory opens opportunities for high-throughput timing attack vectors that are hard-to-mitigate without significant performance overhead.
We introduce IMPACT, a set of high-throughput main memory-based timing attacks that leverage characteristics of PiM architectures to establish covert and side channels. IMPACT enables high-throughput communication and private information leakage. To achieve this, IMPACT (i) eliminates expensive cache bypassing steps required by processor-centric main memory and cache-based timing attacks and (ii) leverages the intrinsic parallelism of PiM operations. First, we showcase two covert-channel attack variants that run on the host CPU and leverage PiM architectures to gain direct and fast access to main memory and establish high-throughput communication covert channels. Second, we showcase a side-channel attack on a DNA sequence analysis application that leaks the private characteristics of a user’s sample genome by leveraging PiM operations. Our results demonstrate that (i) our covert channels achieve up to 14.16 Mb/s communication throughput, which is 5.41$\times$ faster than the state-of-the-art main memory-based covert channels, and (ii) our side-channel attack allows the attacker to determine the properties of a sample genome at a throughput of 7.5 Mb/s with 96% accuracy. We discuss and evaluate several countermeasures for IMPACT to enable secure and robust PiM architectures.
11footnotetext: K. Kanellopoulos and F. N. Bostancı are co-primary authors.
1 Introduction
Data movement between the main memory and the processor is a prevalent bottleneck present in today’s computing systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Hence, both industry and academia have invested significant efforts to bring computation closer to where data resides, creating a paradigm called processing-in-memory (PiM) [13, 23, 8, 24, 25, 26, 11, 27, 19, 28, 1, 2]. PiM solutions provide significant performance improvement and energy efficiency by alleviating data movement.
PiM solutions generally take one of two approaches: (i) processing-near-memory (PnM), and (ii) processing-using-memory (PuM). First, PnM techniques offload computation to compute units near memory arrays (e.g., in the logic layer of 3D-stacked memory or within the memory chip) [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 21, 39, 40, 41, 42]. During recent years, multiple real-world PnM systems have emerged, such as UPMEM [43], AxDIMM [41], AiM [44], HB-PNM [45] and FIMDRAM [46]. At the same time, several academic works propose using general-purpose programmable PiM devices where the user application offloads instructions to the compute units inside the memory at fine-granularity [47] (e.g., PiM-enabled instructions [48]) or coarse granularity (i.e., full program [20, 49]). Second, PuM techniques propose exploiting the analog properties of memory operations and performing specialized PiM operations. Prior works improve the performance of workloads by providing bulk data copy and initialization, bitwise operations, and arithmetic operations using memory [50, 51, 52, 53, 54, 55, 56, 49, 57, 58, 59, 60, 61, 62, 63, 64]. However, to the best of our knowledge, no prior work evaluates the security of PiM architectures against timing attacks.
Analyzing and evaluating the security of PiM architectures is critical to prevent widespread vulnerabilities as PiM architectures become widely available. After a thorough analysis of existing PiM architectures, in this work, we show that the adoption of PiM solutions creates opportunities for high-throughput main memory-based timing attacks for two main reasons: (i) PiM provides direct access to main memory , and (ii) a large set of existing practical defense mechanisms are not applicable in PiM systems.
1. Direct Access to Main Memory. Timing attacks in PiM architectures become easier to exploit with the direct memory access provided to user space applications by these architectures. The adoption of PiM techniques changes how applications and computation units access main memory: (i) due to area and thermal dissipation limitations, compute units in PnM architectures have small or no caches that allow the user space applications to bypass caches more easily (i.e., with fewer memory accesses), and (ii) PuM architectures provide the ability to offload specialized PiM instructions (e.g., bulk data copy and initialization, and bitwise/arithmetic operations) to the main memory that completely bypass the deep cache hierarchies of modern processors.
2. Hard-to-Mitigate with Practical Defenses. A fundamental solution to memory-based timing attacks is to eliminate the timing channel itself. Prior mitigations explore fundamental solutions such as memory partitioning and enforcing constant-time memory accesses. However, these mitigations incur significantly high performance overheads (as shown in Section 6). Therefore, practical mitigations and detection mechanisms restricting access to timing channels are explored. Several prior works [65, 66, 67, 68] propose cache management-based methods such as restricting the access to cache management instructions (e.g., clflush instruction in x86 systems [65]) and detecting timing attacks based on cache access patterns (e.g., a high number of cache misses can indicate a timing attack) [66, 67, 68]. These mechanisms are inapplicable to PiM-based attacks because these attacks completely bypass the cache hierarchy.
Our goal is to develop a set of high-throughput and hard-to-mitigate timing side and covert-channel attacks in PiM architectures. We expect and hope that IMPACT will bring attention to security issues in emerging PiM architectures and result in security awareness in PiM architecture designs.
In this work, we introduce IMPACT, a set of high-throughput main memory-based timing attacks that leverages PiM architectures to gain direct and fast access to main memory, establish high-throughput communication covert channels and leak secrets of concurrently running applications. IMPACT leverages the direct access, provided by PiM operations, to the shared DRAM row buffer and achieves high communication and information leakage throughput by (i) eliminating expensive cache bypassing steps required by other main memory- and cache-based timing cover-channel attacks and (ii) leveraging the intrinsic parallelism of PiM operations.
Key Challenge and Limitations of Prior Work. For successful main memory-based timing side- and covert-channel attacks, attacker processes (or sender and receiver processes for covert-channel attacks) need to access the main memory with high frequency to observe or change the states of the shared main memory structures. In traditional architectures, it is difficult to access the main memory with high frequency due to the deep cache hierarchies that filter the memory requests of attacker processes.
We analyze (Section 3.2) the effectiveness of the attack primitives that bypass caches and provide direct access to the main memory to leverage the main memory timing channel: clflush instruction, eviction sets, (remote) direct memory access, (R)DMA, requests, non-temporal memory hints (e.g., movnti instruction [69]), and PiM operations. None of these primitives used in traditional architectures provide fast and reliable direct memory access at the same time. In contrast, PiM operations provide fast and reliable direct memory access to the user space applications. Therefore, PiM architectures can be more amenable to high-throughput timing side- and covert-channel attacks compared to traditional architectures.
IMPACT Covert-Channel Attack Variants. We showcase two different timing covert-channel attack variants by exploiting PnM and PuM techniques. First, we build a PnM-based covert channel, IMPACT-PnM, that uses PiM-enabled instructions [48] to directly access the main memory. The sender bypasses the locality monitor employed by the system and offloads PiM-enabled instructions for computation directly inside the main memory. The receiver measures the time it takes to access a specific set of rows to retrieve the message encoded with row buffer hits and conflicts. IMPACT-PnM achieves 12.87 Mb/s communication throughput, which is up to 5.79$\times$ faster than the state-of-the-art main memory-based covert-channel attack [70]. Second, we present IMPACT-PuM, a PuM-based covert channel that leverages the inherent parallelism in main memory using bulk copy and initialization techniques, specifically RowClone [55], to quickly transmit messages through the shared row buffer states. This way, IMPACT-PuM transmits 16-bit large messages 14$\times$ faster than IMPACT-PnM and achieves communication throughput of 14.16 Mb/s (up to 6.38$\times$ faster than the state-of-the-art [70]).
IMPACT Side-Channel Attack. We showcase a side-channel attack on a DNA sequence analysis application: DNA read mapping. Preserving the privacy of DNA sequences is crucial, especially when the sequences are used for medical purposes. DNA read mapping finds the matching locations of short DNA fragments of the sample genome in the reference genome, which acts as a template genome to construct the complete genome of a given sample. Our side-channel attack leverages PiM operations and leaks the private characteristics of a user’s sample genome by observing the memory access patterns of the DNA read mapping application. Our evaluation demonstrates that our side-channel attack enables determining the properties of a sample genome at a throughput of 7.5 Mb/s with 96% accuracy.
Defense Mechanisms. Based on our security analysis, we discuss three and evaluate two potential defense mechanisms that can mitigate IMPACT attacks by eliminating the timing channel. We discuss the performance versus security trade-offs of these countermeasures. We demonstrate that the two best-performing defense mechanisms lead to high performance overheads of 25% and 16% on average across five data-intensive workloads.
This paper makes the following contributions:
  • We present IMPACT, the first timing attack in PiM systems. We leverage the direct memory access enabled by PiM operations to establish high-throughput and reliable timing covert channels.
  • We showcase two IMPACT variants by exploiting PnM and PuM techniques: IMPACT-PnM and IMPACT-PuM. IMPACT-PnM allows the sender and the receiver to exchange messages by using PiM instructions. IMPACT-PuM further leverages the parallelism provided by bulk copy PuM techniques to improve the throughput. We extensively evaluate both IMPACT-PnM and IMPACT-PuM and show that they significantly improve communication throughput compared to the state-of-the-art main-memory-based covert channel.
  • We showcase an IMPACT side-channel attack on DNA read mapping by leveraging PnM instructions. We show that our side-channel attack enables the attacker to leak sensitive information on the victim’s sample genome with high throughput and a low error rate.
  • We evaluate two defense mechanisms to protect the system against IMPACT attacks caused by the adoption of PiM solutions. We discuss the trade-offs of these mechanisms to aid the adoption of robust and secure PiM architectures.
2 Background
2.1 DRAM Organization and Operation
Organization. Figure 1 illustrates a DRAM module’s hierarchical organization. A DRAM chip is organized into multiple DRAM banks, each of which shares a common I/O circuitry in a DRAM bank [71, 72, 73, 74, 75]. DRAM cells are organized as a two-dimensional array to form a bank and are addressed by its row and column [76, 77, 22, 78]. The DRAM cell is connected to a bitline that is used to access the data stored in the cell and a wordline that controls access to the cell. Each DRAM cell in a DRAM row is connected to a common wordline via its access transistors. A bitline connects a column of DRAM cells to a DRAM sense amplifier to read or write data. A row of sense amplifiers is called a row buffer [56, 79, 80, 81, 82, 55]. At the lowest level of the hierarchy, a single DRAM cell comprises 1) a storage capacitor that stores a single bit of data encoded using the charge level in the capacitor and 2) an access transistor that is used to read from and write to the storage capacitor [83, 84, 85, 86, 87].
Figure 1: Organization of a typical modern DRAM module.
Operation. The memory controller serves memory access requests. To serve memory requests, the memory controller issues DRAM commands, e.g., row activation ($ACT$), bank precharge ($PRE$), data read ($RD$), data write ($WR$), and refresh ($REF$). Between consecutive DRAM commands, the memory controller must obey standardized timings to ensure correct DRAM operation. For read and write requests, the memory controller first issues an $ACT$ command to the corresponding bank, which opens the corresponding row. The data in the row is fetched to the row buffer, where it is temporarily stored. The data in the row buffer becomes accessible by a $RD$/$WR$ command when the bitline and cell voltage are restored to a certain level. Hence, the memory controller should wait at least the activation latency ($t_{RCD}$) between an $ACT$ and a $RD$/$WR$ command. All changes are propagated to the DRAM cells in the open row. Subsequent accesses to the same row can be served quickly without fetching the data to the row buffer, creating a row hit [88]. However, in the case where the target row is not in the row buffer, the memory controller issues a $PRE$ command to close the already open row and issues an $ACT$ command to fetch the target row, creating a row conflict/miss. The time interval for issuing 1) a $PRE$ after an $ACT$ must be at least restoration latency ($t_{RAS}$) and 2) an $ACT$ after a $PRE$ must be at least the precharge latency ($t_{RP}$).
2.2 Processing-in-Memory
Processing-in-memory (PiM) is a promising paradigm that moves computation closer to where the data resides, alleviating the data movement bottleneck between memory and the processor [13, 23, 8, 24, 25, 26, 11, 27, 19, 28, 1, 2]. There are two main approaches to PiM: 1) processing-near-memory (PnM) and 2) processing-using-memory (PuM).
Processing Near Memory (PnM). In processing-near-memory (PnM) computation is placed near the memory arrays. PnM takes advantage of the computation capability of conventional memory controllers or the logic layer(s) of the 3D-stacked memory technologies [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 21, 39, 40, 41, 42]. PnM logic can execute portions of applications (e.g., instructions [48, 32], functions [28], whole application threads [20, 49]) depending on the PnM architecture. PnM computation units typically employ small caches due to 1) area constraints in the PnM logic and 2) the ineffectiveness of caches in improving the performance of a wide range of memory-intensive applications. Therefore, various PnM systems [32, 48, 28] execute the memory-intensive regions of applications on the PnM computation units, while executing the more cache-friendly regions of applications on the host CPU with a deeper cache hierarchy.
Processing Using Memory (PuM). Processing Using Memory (PuM) expands the functionality of the memory arrays with minimal changes so that they can perform computation [50, 51, 52, 53, 54, 55, 56, 49, 57, 58, 59, 60, 61, 62, 63, 64]. This approach leverages the properties and operational principles of memory cells and arrays to perform computation, without the need for dedicated logic unit added to the memory system. PuM architectures take advantage of the high bandwidth of the memory arrays.
2.3 Microarchitectural Timing Attacks
Microarchitectural timing attacks exploit the observable state changes in microarchitecture to build covert channel attacks that can propagate messages between processes and leak secrets of concurrently running processes. These attacks measure the time to complete an operation that depends on the shared structures in the system (e.g., accessing a memory location by going through the shared caches) to infer the states of these shared structures.
Main Memory Timing Attacks. Existing main memory timing attacks exploit several shared structures in the main memory to leak data, and one such structure is the row buffer. The row buffer acts as an internal cache inside DRAM banks. Thus, an attacker can understand if a row is recently accessed (i.e., already in the row buffer) or not by measuring access to that row. Based on this observation, prior works propose several attacks, revealing the DRAM address mapping and leaking keystrokes. DRAMA [70] (i) builds a covert channel that exchanges messages between processes using the row buffer conflicts and hits, and (ii) showcases a side-channel attack that can leak keystrokes in a real system. Similarly, many works [70, 89, 90] exploit row buffer-based timing differences to determine DRAM address mappings. Other attacks based on the main memory exploit either the memory bus [91], or observations of latencies incurred by other memory-related operations such as memory deduplication [92, 93], and memory compression [94, 95, 96, 97, 98, 99, 100, 101] and decompression [102].
3 A Case for PiM-based Timing Attacks
The adoption of processing-in-memory (PiM) architectures has been gaining momentum in recent years because they provide high performance and low energy consumption by alleviating the data movement bottleneck. Therefore, it is timely and critical to analyze and evaluate the security of PiM architectures to prevent any of their vulnerabilities from causing widespread susceptibility as PiM architectures become widely available. Unfortunately, no prior work investigates and evaluates the security challenges of new and emerging PiM architectures. After a thorough analysis of existing PiM architectures, we show that the adoption of PiM solutions creates opportunities for high-throughput main memory-based timing attacks for two main reasons. First, the adoption of PiM solutions in the forms of processing-near-memory (PnM) and processing-using-memory (PuM) exposes the internal states of the main memory (e.g., row buffer states of DRAM banks) by providing direct access to the user space applications. Second, timing attacks in PiM architectures can be more difficult to detect and mitigate due to commonly employed mechanisms, like cache monitoring mechanisms that track the cache miss ratio during runtime to detect anomalies [66, 67, 68], being either inapplicable or performance-inefficient in PiM architectures.
3.1 Row Buffer Timing Channel
A DRAM row buffer acts as a direct-mapped cache inside a DRAM bank and is shared across applications that can access main memory. Accessing a row that is already cached in the row buffer results in a significantly lower latency because the memory controller does not need to issue additional commands, $PRE$ to precharge the open row and $ACT$ to fetch the target row to the row buffer. Therefore, it is possible for a user space application to infer whether a row is accessed recently by measuring the time it takes to access the row. To demonstrate the latency gap between different states of the row buffer, we evaluate microbenchmarks that create memory accesses that result in row buffer hits and conflicts, and we measure the latencies observed by the user space application. Figure 2 shows the breakdown of memory access latencies categorized into two classes: row buffer hits and conflicts.
Figure 2: The distribution of memory access latencies.
We observe that the average latency gap between a row buffer conflict and a row buffer hit is 74 CPU cycles. Since this latency gap is large enough to detect, it is possible for user space applications that share the main memory to exploit the row buffer as a covert channel to exchange messages. Our results align with the observations of prior works [70].
3.2 Existing Main Memory Attack Primitives
Main memory timing attacks are powerful as they can be used to gather information on applications that share main memory and establish covert channels to exchange messages between different processes. However, these covert-channel attacks are limited in throughput in traditional architectures for two main reasons: (i) accessing a memory location in main memory requires expensive cache bypassing techniques due to frequently accessed data being cached to improve performance, and (ii) even when bypassing the cache hierarchy, deep cache hierarchies employed in modern processors incur an additional high lookup latency on memory requests that are eventually sent to main memory.
There are several ways to bypass the cache hierarchy in traditional systems by using one or more of four attack primitives: (i) using specialized instructions to flush a cache line (e.g., clflush), (ii) using cache eviction sets (i.e., by creating many conflicting memory requests to replace the cache line with newly-fetched cache lines), (iii) memory access requests generated by direct memory access (DMA) and remote direct memory access (R-DMA) engines, and (iv) non-temporal memory hints (e.g., movnti). However, when used in a main memory-based timing covert channel attack, these methods do not provide fast and reliable communication.
Table 1: Efficiency and Effectiveness of Attack Primitives
Low Latency Effectiveness
No Cache
Lookup
No Excessive
Memory Accesses
Timing Difference
Detectability
ISA
Guarantees
Specialized
Instructions
Eviction Sets
DMA/R-DMA N/A
Non-temporal
Memory Hints
PiM Operations
We identify four properties of an efficient (i.e., low-latency) and effective attack primitive: 1) avoiding cache lookup overhead, 2) avoiding high-latency memory accesses, 3) creating an easily detectable timing difference, and 4) reliable functionality (e.g., guaranteed to work by the ISA). Table 1 shows a comprehensive comparison of these four attack primitives in terms of these properties.
First, specialized cache line flush instructions have two main disadvantages that hurt the efficiency and effectiveness of attacks that use them: 1) they put the latency of writing back a cache line to the main memory on the critical path to access a memory location, and 2) in some modern systems, these instructions are privileged and cannot be used by user space applications.
Second, eviction sets require a large number of load requests proportional to the number of ways in the caches, which scales with modern designs. Therefore, the attacks that use them have limited throughput. While the ISA guarantees the instructions used to create the eviction sets, the eviction set is not guaranteed to evict the target cache line (due to cache characteristics such as cache replacement policy, cache prefetchers etc.).
Third, DMA and R-DMA engines incur high overheads due to deep software stacks. As a result, attacks using this primitive need to utilize states with larger timing differences to be distinguishable. An example attack using this primitive[103] showcases a covert channel attack based on memory contention. However, the basis of this type of covert channel is not as fine-grained as row buffer-based timing differences. Therefore, the communication throughput these covert channels provide is limited (in the order of Kb/s).
Fourth, the implementation of non-temporal memory hints varies across systems, as system designers are not bound to a specific approach. Consequently, the ISA does not guarantee that memory accesses related to non-temporal hints will consistently bypass caches.
In contrast, PiM architectures provide fast and reliable direct memory access to user space applications that can be used to establish a covert-channel attack.
3.3 High Throughput Main Memory-based Covert Channels
To demonstrate the impact of the latencies induced by deep cache hierarchies (i.e., cache access and cache eviction latencies) on communication throughput, we simulate a realistic system (described in Section 5.1) that establishes a state-of-the-art covert channel [70].
Attack Scenarios. We simulate two attacks. First, we evaluate a baseline attack that relies on cache bypassing using cache eviction sets. To transmit a bit through the row buffer, the baseline attack first evicts the corresponding cache line and then, it fetches the target row into the row buffer using an additional memory request. The baseline attack accomplishes cache line eviction by issuing N memory requests to the target cache set in an N-way cache. Note that the actual eviction latency can be much higher due to a larger number of memory accesses required to bypass more complex replacement policies employed in today’s systems or due to address translation overheads. Second, we evaluate a direct memory access attack that can directly access the main memory. The direct memory access attack issues a memory request to fetch the data to the row buffer and does not need any cache access or eviction steps.
For both attacks, we measure the time it takes to send a bit through the row buffer including cache access and cache eviction latencies, and with encoding logic-1 values as row buffer hits and logic-0 as row buffer conflicts. We calculate the communication throughput based on the average time needed to propagate bits from the sender to receiver. We analyze the impact of (i) the LLC size, and (ii) the number of LLC ways on communication throughput to show the limitations in current and future systems with different cache designs.
Impact of LLC Size on Communication Throughput. Figure 3 shows the throughput of baseline and direct memory access attacks. The x-axis shows the cache configuration with increasing last-level cache (LLC) sizes and a fixed number of ways (i.e., 16). To calculate the cache access latency with increasing LLC sizes, we followed the same methodology used in prior works [104]. To better understand the throughput results, Figure 4 shows the latency of a cache line eviction of caches with corresponding sizes.
Figure 3: Impact of the LLC size on the data leakage throughput of covert channels.
Figure 4: Latency of a cache line eviction in CPU cycles with increasing LLC size.
We make two key observations. First, the communication throughput of the direct memory access attack is $11.27Mb/s$ across all LLC sizes, which is significantly greater than the throughput of up to $2.29Mb/s$ achieved by the baseline attack. Second, the baseline attack’s throughput decreases with increasing LLC sizes due to increasingly expensive cache access and eviction latencies. We observe that with increasing LLC size, the eviction latency grows significantly due to the increase in access latency. We conclude that with increasing LLC sizes, the cache access and eviction latencies grow significantly and significantly hinder the communication throughput of the baseline attack compared to the direct memory access attack which is independent of the LLC size.
Impact of LLC Associativity on Communication Throughput. Figure 5 shows the communication throughput of baseline and direct memory access attacks. The x-axis shows the increasing number of LLC ways (from 2 to 128) while the total LLC size is fixed at 16 MB (modern processors typically employ 8 cores and a 2MB LLC slice per core). Increasing the number of LLC ways increases the number of memory requests required to evict a cache line from the cache. To better understand this effect, Figure 6 shows the latency of a cache line eviction across increasing numbers of LLC ways. We observe that the baseline attack’s communication throughput decreases significantly with increasing number of LLC ways due to the increased cache eviction latency.
Figure 5: Impact of the number of LLC ways on the data leakage throughput of covert channels.
Figure 6: Latency of evicting a set in CPU cycles with increasing number of LLC ways.
We conclude that attacks that communicate with main memory through caches are greatly limited in throughput due to expensive cache lookup and eviction latencies, which worsen with increasing size and associativity. However, a covert channel based on direct memory access can achieve high throughput independent from cache configurations as it does not require bypassing the cache hierarchy.
Adoption of PiM solutions can fundamentally enable high-throughput timing attacks through main memory by providing direct access to the main memory to user space applications. Therefore, it is crucial to carefully consider the security implications of PiM architectures and to develop robust security measures to protect against potential attacks. Our goal is to develop a set of high-throughput and hard-to-mitigate timing side and covert-channel attacks in PiM architectures. We expect and hope that IMPACT will bring attention to security issues in emerging PiM architectures and result in security awareness in PiM architecture designs.
4 IMPACT: A High-Throughput Main Memory Timing Attack
This section introduces IMPACT, a set of high-throughput main memory timing attacks that takes advantage of the direct memory access enabled by the emerging processing-in-memory paradigm. IMPACT (i) operates independently of specialized instructions that flush data from the cache hierarchy and (ii) avoids the resource-intensive cache eviction steps typically required for main memory and cache-based timing attacks, achieving higher data leakage throughput than previous memory-based attacks. In this work, we present (1) two variants of IMPACT covert channel attack and (2) an IMPACT side-channel attack on DNA read mapping applications enabled by processing-near-memory (PnM) and processing-using-memory (PuM) solutions. We evaluate the communication throughput of IMPACT covert channels in Section 5.3 and the leakage throughput of IMPACT side channel in Section 5.4.
4.1 IMPACT Covert-Channel Attacks
4.1.1 PnM-based Covert-Channel Attack
Baseline PnM Architecture. In IMPACT-PnM, we consider PiM-Enabled Instructions (PEI)  [48] as the baseline PnM architecture. PEI expands the instruction set architecture (ISA) of the host processor with PIM-enabled instructions (PEIs), while keeping the existing programming model intact. A PEI is an instruction that can be executed either on the host processor or on the in-memory computation logic which is integrated near each DRAM bank of an HBM stack. PEI can also easily be adopted in other memory technologies since its design does not rely on specific features of the memory technology. We chose PEI as the baseline PnM architecture as it introduces a simple and effective model (i.e., offloading computation at the instruction granularity) for PnM, and our attack vector can be generalized for other PnM architectures with similar design components (e.g., FIMDRAM [46]).
Figure 7 depicts the organization of PEI. PEI architecture consists of two key components. First, PEI integrates a PEI Computation Unit (PCU) near each DRAM bank  1 and inside the CPU  2. In-memory PCUs are shared among all the cores of the host processor and are responsible for executing PEIs near the respective DRAM bank. Second, PEI adds a PEI Management Unit (PMU)  3 integrated with the last-level cache of the host system to (i) monitor the data locality of application regions and (ii) map each region to host-side or memory-side PCUs based on their data locality. This way a PEI can be selectively executed on the host processor PCU (instead of the in-memory PCU) when large on-chip caches are beneficial for its execution. As mentioned in [105], if a locality monitor entry is allocated by a PIM operation, it is often too aggressive to consider it as having high locality on its first hit in the locality monitor. Therefore, they propose adding an ignore flag per entry to ignore the first hit to an entry allocated by a PIM operation. Such a design allows us to easily bypass the locality monitor even with a small virtual address range in our disposal.
Figure 7: Overview of PEI’s architecture.
Attack Overview. The PnM-based covert channel (IMPACT-PnM) leverages the direct memory access provided by PiM-enabled instructions, which bypasses the cache hierarchy. IMPACT-PnM exploits the timing difference between the latency of a DRAM row buffer hit and a conflict to exchange messages between the sender and the receiver. The sender transmits bits of the message to the receiver in multiple batches of M bits, where the value of each bit is encoded as interference (i.e., row buffer hits or conflicts) in the row buffer of a DRAM bank. We assume the sender and the receiver co-locate data in the same DRAM banks. [^1] If PiM is the main memory, then the sender can co-locate its data with the receiver process’s data by using memory massaging techniques, as used in many prior works [90, 70, 106, 107]. Two processes can share data itself in various cases, such as databases and websites running on browsers. In this attack, we encode logic-1 as interference in the DRAM bank, and logic-0 as no interference.
Figure 8 and Listing 1 show the end-to-end flow of IMPACT-PnM and all the steps involved in establishing the covert channel between the sender and the receiver.
Step 1. The receiver issues PEI operations to initialize each DRAM bank by activating a predetermined row (Line 18). It is necessary that the cache lines of the activated row are not used before in order to bypass the locality monitor (PMU) that is employed in PEI-based systems. Then, the sender and the receiver synchronize on a barrier (Lines 5&20) which ensures that they start the communication turn at the same time.
Step 2. The sender starts the communication turn by issuing PEI instructions to activate a row in each DRAM bank (Lines 7-12) or not based on the bit value that will be transmitted. To transmit a 0, as shown in Fig. 8 ( 1) the sender does not activate any row in the corresponding DRAM bank to avoid causing interference in the row buffer (e.g., Row 0xB remains activated in the row buffer  2). To transmit a 1, as shown in Fig. 8 ( 3) the sender activates a row that is different from the one activated by the receiver to cause interference in the corresponding DRAM bank (i.e., a row buffer conflict between Row 0xD and Row 0xE  4). At the end of the communication turn (after transmitting M bits), the sender first executes a memory fence (Line 14) to ensure that all the PEI instructions have been executed before the receiver starts probing the DRAM banks and then signals the receiver to start probing the DRAM banks (Line 15).
Step 3. The receiver starts probing the DRAM banks by issuing PEI instructions to activate the same row as the one activated during initialization (Line 25). Then, the receiver measures the time it takes to complete the request (Line 27). If the latency of executing the PEI exceeds a predetermined hit threshold, the receiver senses the interference in the row buffer and decodes the bit as a 1 (Lines 28-29  5, 6). Otherwise, the access is considered a row buffer hit and the receiver decodes the bit as a 0 (Lines 30-31  7, 8). The sender and the receiver repeat the same steps for each batch until the entire message is transmitted.
1 Sender:
2 message = [0: N-1]; //N banks and N-bit message
3 batch_size = M; //M-bits per batch where M<N
4 init_semaphore(0);
5 barrier(); // synchronize with receiver
6 bank=0;
7 for every batch:
8 for each bit in batch_size;
9 if message[bank] == 1:
10 pim_add(row[bank]);
11 else:
12 NOP();
13 bank++;
14 memory_fence(); // signal receiver
15 increment_semaphore(); // to start probing
16
17 Receiver:
18 init_row = initialize_DRAM_rows_with_PEIs();
19 bank = 0;
20 barrier(); // synchronize with sender
21 for every batch:
22 block_on_semaphore();//if semaphore==0, wait
23 for each bit in batch_size;
24 timer(start);
25 pim_add(init_row[bank]);
26 timer(end);
27 latency = end - start;
28 if latency > threshold:
29 result = ’1’;
30 else:
31 result = ’0’;
32 bank++;
Listing 1: Pseudocode of PnM-based Covert-Channel Attack.
Figure 8: Flow of the PnM-based Covert Channel Attack.
Sender-Receiver Synchronization. The sender and the receiver synchronize in order to ensure that the sender’s and receiver’s actions (i) are accurately aligned so that they do not interfere with each other and fail to communicate the message and (ii) to overlap the sender and receiver operations to increase the throughput of the attack. To achieve this, the sender and the receiver communicate using semaphores. The value of the semaphore indicates the number of batches that the sender has completed transmitting but the receiver has not yet completed probing. When the sender completes the transmission of a batch of bits, they increment the semaphore to signal the receiver to start probing the DRAM banks. This way, the sender overlaps the transmission of multiple batches of bits with the receiver’s probing operations without the need to wait for the receiver to complete the probing of the DRAM banks. Doing so, the sender can freely proceed to future batches without relying on synchronizing with the receiver. The receiver blocks on the semaphore until the sender increments it. When the semaphore is incremented, the receiver decrements it and starts probing the DRAM banks. When the receiver completes the probing of the DRAM banks, they read the semaphore again and the process repeats until the entire message is transmitted. In our attack implementation (see §5.1) we use POSIX semaphores to implement the sender-receiver synchronization and use a batch size of 4 bits.
4.1.2 PuM-based Covert Channel
Baseline PuM Architecture. We assume a PuM architecture that provides user applications with RowClone [55] that enables bulk data movement operations (e.g., bulk data copy and initialization) within DRAM’s subarray boundaries by exploiting the internal operations of DRAM. [^2] There could be near-memory RowClone variants that move data without using operational principles of DRAM as well. However, in this attack, we assume a PuM-based RowClone operation. For our attack, we focus on the Fast Parallel Mode in RowClone, which copies the data of a row to another row within a DRAM subarray.
Figure 9 demonstrates how the content of the source row, i.e., src, is copied to the destination row, i.e., dst using RowClone’s Fast Parallel Mode. Initially, the rows and the row buffer are closed ( 1). RowClone operation consists of two key steps. First, the src is activated to store the entire row’s data in the row buffer ( 2). Second, the dst is activated, which leads to copying the contents of the row buffer in the dst ( 3). RowClone significantly reduces the latency and energy consumption of the bulk data copy operation from one row to another as it enables data movement without transferring data between the DRAM and the processing units. To execute RowClone, we assume that the user space application specifies a source and a destination virtual address range and the corresponding mask to copy only specific parts of the source to the destination. This way, the memory controller receives one RowClone request and breaks it into parallel RowClone requests, one for each set bit of the mask.
Figure 9: Overview of RowClone’s Fast Parallel Mode.
Attack Overview. PuM-based covert channel attack (IMPACT-PuM) exploits in-DRAM bulk copy operations to establish a covert channel between a sender and a receiver. The sender transmits a message by executing a RowClone operation with three operands: (i) a source address range, (ii) a destination address range, and (iii) a mask to select which parts of the ranges will be copied. Source and address ranges correspond to physical pages spanning across all DRAM banks and by using the mask, the sender can encode the transmitted message by causing contention only in selected DRAM banks. The receiver repeatedly issues RowClone operations and measures the time it takes to complete these operations. This way, it can detect contention in a DRAM bank when another process accesses a different row in the same DRAM bank. If the sender issues a RowClone operation to a DRAM bank, it induces contention and the receiver detects a higher latency for its RowClone operation. This latency indicates that the corresponding bit is a logic-1. Otherwise, if the sender does not induce contention to a DRAM bank, the receiver detects a lower latency for the RowClone operation, indicating that the corresponding bit is a logic-0.
Figure 10 and Listing 2 show the end-to-end flow of IMPACT-PuM and all the steps involved to establish the covert channel between the sender and the receiver.
Step 1. First, the receiver initializes all DRAM banks by issuing a RowClone operation to copy the contents of a source row to a destination row (Line 16). Then, the sender and the receiver synchronize on a barrier (Lines 10&17) which ensures that they start the communication turn at the same time.
Step 2. The sender uses a single RowClone operation to transmit in parallel an N-bit message (where N is the number of banks) to the receiver and encodes the message in the mask operand of the RowClone operation (Lines 11-12). The sender encodes the message by setting the corresponding bits in the mask operand to 1 or 0 ( 1, 2). Doing so, the sender causes contention in the corresponding DRAM banks and activates a different row in the row buffer. To send a 1, the sender triggers RowClone in the corresponding DRAM bank and induces contention by altering the activated row in the row buffer ( 3). To send a 0, the sender does not trigger RowClone in the corresponding DRAM bank and does not induce contention ( 4). The sender synchronizes with the receiver using a barrier operation (Lines 13&18).
Figure 10: End-to-End Flow of the Processing-using-Memory Covert Channel Attack.
1 Memory Controller:
2 row_clone(src, dst, mask):
3 for each DRAM bank do:
4 if mask[bank] == 1 then:
5 row_clone(bank, src, dst);
6 else:
7 NOP();
8
9 Sender:
10 barrier_1();
11 mask = init(message);
12 row_clone(src, dst, mask);
13 barrier_2();
14
15 Receiver:
16 init_DRAM_rows_with_RowClone();
17 barrier_1();
18 barrier_2();
19 for each DRAM bank do:
20 mask = 1 << bank;
21 timer(start);
22 row_clone(src, dst, mask);
23 timer(end);
24 latency = end - start;
25 if latency > threshold:
26 result = ’1’;
27 else:
28 result = ’0’;
Listing 2: Pseudocode of PuM-based Covert-Channel Attack.
Step 3. The receiver in IMPACT-PuM detects contention in DRAM banks by issuing RowClone operations and measuring their latency. The receiver continuously issues one RowClone operation for each DRAM bank (in contrast to the sender who issues only a single RowClone operation), to the initial virtual address range (i.e., rows initialized in Step 1) and measures the latency of each operation ( 2, Line 24). If the sender induces contention to a DRAM bank, the receiver high latency for the RowClone operation which indicates that the corresponding bit in the message is set to 1 ( 3, Lines 25-26). Otherwise, the receiver detects a fast RowClone operation which indicates that the bit is set to 0 ( 4, Lines 27-28).
Advantages over PnM-based Covert Channel. IMPACT-PuM improves upon the PnM-based covert-channel attack in a major way. Using a single RowClone operation, the sender can transmit in parallel an N-bit message to the receiver, where N is the number of DRAM banks which requires less computational resources to transmit the same amount of information compared to the PnM-based covert-channel attack. For example, using a single thread, the sender can activate all N DRAM banks and transmit N bits of information in parallel while in the PnM-based covert-channel attack, the sender needs to separately execute a PEI instruction to activate each DRAM bank.
4.2 PnM-based Side-Channel Attack on DNA Read Mapping
DNA Read Mapping. DNA is one of the most unique identifiers of an individual, such that even a small fraction of DNA can include sufficient information for linking to the owner’s identity[108]. Therefore, preserving DNA privacy carries importance. DNA sequencing of a biological sample produces short DNA fragments called reads. Read mapping is a fundamental operation to find the matching locations of reads in the reference genome that acts as a template genome to construct the complete genome of the sample.
Figure 11 shows a high-level overview of a DNA read mapping (DNA-RM) workflow. DNA-RM has two major key steps: (i) seeding and (ii) alignment. During the seeding step, we identify the possible locations of the reads on the reference genome based on sequence similarities between the reads and the reference genome. Seeding involves hashing small segments of the DNA sequences (i.e., seeds) and probing a hash table that is built from the reference genome. After seeding, we align the reads with their candidate regions on the reference (i.e., potential reference segments) using dynamic programming-based methods to find each read’s best matching location among all possible locations in the alignment step.
Figure 11: A simple overview of DNA read mapping workflow: (a) seeding and (b) alignment
DNA-RM requires a large amount of data movement to access different parts of the hash table and then corresponding regions on the reference genome. Hence, several prior works aim to improve the performance of DNA-RM by offloading seeding and alignment to PiM-enabled systems to reduce the data movement overheads [33, 109, 110, 111, 112, 113, 114, 115]. In this work, we demonstrate that IMPACT can be used to launch a high-throughput side-channel attack against DNA-RM and quickly leak a large amount of information about the sample genome.
Attack Overview. We assume that the victim is running a DNA-RM application on a PiM-enabled system and the attacker is able to run a malicious application on the same system. The attacker and the victim are using the same read mapping tool (e.g., minimap2 [116]), and the hash table is shared between the attacker and the victim. [^3] The hash table that contains information about the location of seeds in the reference genome is created once (acts as a database), and each user (attacker and victim in our case) of minimap2 probes it with queries. The attacker’s goal is to leak information about the test genome by observing the DNA-RM application’s memory access patterns.
Figure 12 shows the high-level overview of the attack. The victim application hashes each seed of the test genome and offloads the corresponding seeding and alignment step to the PiM-enabled system  1. During the seeding step, the corresponding DRAM row gets activated (i.e., Index 0-5 of the hash table)  2. The attacker continuously uses PnM instructions to probe the shared hash table in an attempt to identify if the victim application is accessing a specific index of the hash table  3. If the attacker observes that the victim application is accessing a specific index, it retrieves partial or exact information about the potential locations of the test genome in the reference genome. As shown in previous works on DNA imputation [117, 118], the attacker can use the leaked information to infer properties about some regions of the private sample genome  4.
Figure 12: High-level overview of IMPACT attack model on PiM-enabled DNA-RM
5 Evaluation
5.1 Methodology
Covert Channel Threat Model. Our threat model for the covert-channel attack envisions a scenario where a sender and a receiver operate on the same PiM-enabled system to exchange information. PiM instructions are available to user space programs without any restrictions and are not privileged instructions. The receiver is assumed to have access to cpuid and rdtscp instructions, enabling high-precision measurement of memory access times. For IMPACT-PnM, the sender and the receiver need to colocate data in the same DRAM banks. For IMPACT-PuM, (i) the sender allocates two virtual address ranges whose corresponding physical pages span across all DRAM banks, and (ii) RowClone guarantees atomicity - when the sender issues a RowClone operations that span across multiple banks, all individual bank-level RowClone operations need to be completed before another process accesses DRAM.
Side Channel Threat Model. Our threat model for the side-channel attack envisions a scenario where an attacker process operates on the same PiM-enabled system as a victim process and leaks information about the victim process. We use a DNA-RM implementation based on minimap2 [116] and we modify the implementation to offload the seeding and alignment step to the PiM-enabled system. We assume that the alignment step includes chaining. We use the human genome as the reference and compare it against synthetic sample genomes. We assume that the hash table is distributed across each bank (i.e., DRAM employs a page interleaving mapping). We experiment with multiple different seed sizes, and we report the best-performing selection.
Proof-of-Concept. We construct a proof-of-concept of all IMPACT variants using the Sniper Multicore Simulator [119]. We choose to construct the PoC using Sniper for two reasons. First, it is validated against modern Intel processors [120]. Second, there is no commercially available system that follows a similar architecture as PEI or any modern system that enables end-to-end RowClone operations. We extend Sniper to (i) accurately model all internal DRAM operations, (ii) emulate and simulate the PiM interface and PiM operations, and (iii) emulate the functionality of the cpuid and the rdtscp instructions in order to measure execution cycles from the user space.
Table 2 shows the configuration our simulation setup. Before launching the attack, the sender and the receiver threads warm up by executing both attacks in order to avoid compulsory DRAM accesses for instructions or page table walks (i.e., after an L2 TLB miss, the page table walker might issue requests to DRAM) and reduce the noise of the attack. For IMPACT-PnM, we assume that both the attacker and the receiver use PEIs to perform an addition operation that takes $\sim 3$ cycles to execute (in addition to the actual DRAM access).
Comparison Points for Covert-Channel Attacks. To showcase the effectiveness of IMPACT, we evaluate the leakage throughput of five different covert-channel attacks.
(i) DRAMA-clflush: The main memory covert-channel attack proposed in [70] which uses clflush instructions to send memory requests directly to main memory. clflush only probes the LLC to flush the cache line.
(ii) DRAMA-Eviction: The main memory covert channel attack proposed in [70] which uses cache eviction sets to send memory requests directly to main memory. In both DRAMA attacks, the sender creates contention in the DRAM banks by causing row buffer conflicts and the receiver detects the contention to decode the message.
(iii) Streamline: A state-of-the-art cache covert channel proposed in [121] which enables a flush-less collusion between the sender and the receiver.
(iv) DMA Engine: A main memory attack that uses the DMA engine to bypass the caches and send memory requests directly to main memory. In this scenario, we assume a powerful attacker can infer the DRAM access time directly from the user space by disregarding context switch and OS-related latencies.
(v) PnM-OffChip: A PnM-based attack that probes the locality monitor and a state-of-the-art off-chip predictor [122] before offloading PEI instructions to compute units near DRAM banks. If the off-chip predictor predicts that data experience high locality, the computation is performed by the CPU instead of the PnM, leading to reduced throughput.
(iv) IMPACT-PnM: Our proposed attack that bypasses the locality monitor and offloads PEI instructions directly to compute units near DRAM banks in order to establish a covert channel between the sender and the receiver.
(v) IMPACT-PuM: Our proposed attack that makes use of RowClone to offload bulk-copy operations directly to main memory and establish a covert channel between the sender and the receiver.
For a fair comparison against DRAMA and Streamline, we showcase the upper bound of the communication throughput achieved by each attack. To evaluate the throughput of the DRAMA and Streamline-based attacks, we use our simulation infrastructure to extract parameters such as the LLC hit latency, average LLC miss latency, cache lookup latency, cache hit/miss ratio, and feed them in an analytical model that calculates the maximum throughput of the attack. We validated the analytical model against the throughput achieved by DRAMA and Streamline in a real system. For example, in Streamline, the authors report a maximum communication throughput of 1.8M Mb/s and our model estimates that the upper bound of the throughput is 2.7 Mb/s for the smallest LLC size.
Noise Sources in Simulation Setup and Error Rate. There are two main microarchitectural structures that generate noise in the simulated system: 1) hardware prefetchers and 2) page table walkers. To observe the impact of application-induced noise on the throughput of IMPACT attacks, we run applications from GraphBIG [123] benchmark suites concurrently with attacker and victim (or sender and receiver) processes. We measure the throughput of each attack by incorporating the error rate in the measurement: we measure the throughput only based on successful leakage.
Table 2: Simulation Configurations
CPU 4-core, 4-way, OoO x86 core, 2.6GHz
MMU L1 I-TLB: 128-entry, L1 DTLB (4KB): 64-entry, 4-way assoc, 1-cycle access, L1 DTLB (2MB): 32-entry, 4-way assoc, 1-cycle access, L2 TLB: 1536-entry, 12-way assoc, 12-cycle access, 3 Split PWCs: 32-entry, 2-cycle access
L1 Cache I-cache: 32 KB; L1D Cache: 32 KB, 8-way assoc, 4-cycle access latency, LRU, IP-stride prefetcher [124]
L2 Cache 1 MB, 16-way assoc, 12-cycle latency, SRRIP [125], Streamer [126]
L3 Cache 2 MB/core, 16-way assoc, 32-cycle latency, SRRIP [125]
Main Memory DDR4-2400, 16 Banks, 4 Ranks, 1 Channel, Row size = 8192 bytes, $t_{RCD}$ = 13.5 ns, $t_{RP}$ = 13.5 ns , $t_{RC}$ = 13.5 ns, Open Row policy, Row Timeout = 100 ns
5.2 Validation of Proof-of-Concept
We first demonstrate the PoC results and how we successfully mount the attack in the simulated system. Figures 13 and 14 show the latency (in cycles) measured by the receiver after executing a PiM instruction, one for each of the 16 banks (we assume that the message is 16-bits large), in IMPACT-PnM and IMPACT-PuM, respectively. As we observe, the receiver is able to successfully distinguish whether the sender causes contention in the bank in both IMPACT-PnM and IMPACT-PuM. As shown in the figure, given a latency threshold of 150 cycles, the receiver can decode the complete message: if the latency is less than 150 cycles (i.e., row buffer hit), the receiver decodes the message as a logic-0, otherwise (i.e., row buffer conflict) it decodes the message as a logic-1.
Figure 13: Measured latency (in cycles) by the receiver to execute a PEI instruction for each one of the 16 banks in order to decode a 16-bit message.
Figure 14: Measured latency (in cycles) by the receiver to execute a RowClone instruction for each one of the 16 banks in order to decode a 16-bit message.
5.3 Throughput Analysis of IMPACT Covert-Channel Attacks
We examine the throughput of all different attacks to understand the potential of IMPACT in establishing a high throughput covert channel between the sender and the receiver. Figure 15 shows the throughput achieved by all different attacks when one sender and one receiver thread run concurrently across systems with increasing LLC sizes. We make four key observations. First, IMPACT-PnM and IMPACT-PuM achieve significantly higher throughput than all other attack vectors. For example, IMPACT-PnM achieves a throughput of 12.87 Mb/s irrespective of the cache size, which is 2.29$\times$ (up to 4.91$\times$ with an LLC size of 128MB) higher than the throughput achieved by the second-best performing attack, DRAMA-clflush. Second, DRAMA and Streamline achieve decreasingly lower throughput as the LLC size increases. This is because the larger the LLC size, the higher the latency it takes to access the LLC. Third, IMPACT-PuM achieves 10% higher throughput than IMPACT-PnM since the sender can issue all RowClone requests in parallel, whereas in IMPACT-PnM the sender has to send the PiM instructions in a sequential manner. Fourth, the DMA engine attack achieves a throughput of 5.27 Mb/s irrespective of the cache size, but is almost 2.4x slower than the PnM-based one due to the OS-related overheads associated with it (e.g., context switch overhead, OS instructions, etc.). Fifth, the PnM-OffChip attack leads to high throughput, up to 12.64 Mb/s, but the throughput gets reduced down to 10.64 Mb/s as the cache size increases since the off-chip predictor consults more in favor of caching data when the LLC is large, essentially reducing the attack throughput. We conclude that IMPACT-based attacks can achieve significantly higher throughput than the existing state-of-the-art cache covert channels.
Figure 15: Communication throughput comparison across increasing LLC sizes.
Breakdown of Throughput Improvements. The impact of the fundamental difference between PiM architectures and traditional architectures can be seen as the difference between PnM-OffChip and DRAMA-clflush attack throughputs in Fig. 15. The impact of bypassing the locality monitor can be seen as the difference between IMPACT-PnM and PnM-OffChip attack throughputs. Finally, the impact of the parallelism provided by RowClone can be seen as the difference between IMPACT-PuM and IMPACT-PnM attack throughputs.
To better understand why IMPACT-PuM achieves higher throughput than IMPACT-PnM, we evaluate the time it takes for the sender and the receiver to perform their respective operations in both attacks. Figure 16 shows the time (in cycles) it takes for the sender to send the message and for the receiver to read the message in IMPACT-PnM and IMPACT-PuM. As we observe, in IMPACT-PuM, the sender can send the message 14$\times$ faster than the sender in IMPACT-PuM since it can issue all RowClone requests in parallel. In IMPACT-PnM, even though the sender is significantly slower than the sender of IMPACT-PuM, the sender and the receiver overlap, at the cost of synchronization, in order to hide the latency of the sender. This is why IMPACT-PnM achieves only 10% lower throughput than IMPACT-PuM.
Figure 16: Time (in cycles) it takes (i) for the sender to send the message and (ii) for the receiver to read the message in IMPACT-PnM and IMPACT-PuM.
5.4 Throughput Analysis of IMPACT Side-Channel Attack
Figure 17 shows the throughput (taking into account only correct identification of activated hash table entries), left y-axis, and the error rate (i.e., wrong identification of activated hash table entries), right y-axis, of the side-channel attack against DNA-RM across different number of DRAM banks. We make two key observations with respect to the throughput of the attack. First, when using a PiM-enabled DRAM device with 1024 banks, the attacker is able to leak information about the sample genome at a throughput of 7.57 Mb/s. Second, as the number of banks increases, even though the throughput of the attack decreases down to 2.56 Mb/s (since the attacker has to probe more banks and the error rate also increases), the leaked information provides more exact information about the sample genome since the amount of hash table data per bank decreases, reducing the number of potential locations of the reads in the reference genome.
We make two key observations with respect to the error rate of the attack. First, for the 1024 bank configuration, the error rate of the attack is less than 5%. Second, we observe that the error rate of the attack is less than 15% across all different DRAM bank configurations. As the number of banks increases, the error rate of the attack increases since the attacker has to probe more banks, which makes the attack more prone to noise and reduces its bandwidth(i.e., DRAM command delay, memory controller scheduling, etc.).
Even though the error rate of the attack increases as the number of banks increases, the attacker is able to leak information about the test genome with higher accuracy. For example, when using a PiM-enabled DRAM device with 1024 banks, the victim activates a row with 16 hash table entries, and the attacker is able to identify if the sample read is present in any of the locations pointed by the 20 hash table entries with an accuracy of 96%. When using a PiM-enabled DRAM device with 2048 banks, the victim activates a row with 8 hash table entries, and the attacker is able to identify if the sample read is present in any of the locations pointed by the 8 hash table entries with an accuracy of 91%. The exact end-to-end accuracy of identifying the properties of the sample genome depends on the chaining mechanism and the imputation algorithm [117, 118] used by the attacker, and it is out of the scope of this work.
Figure 17: Leakage throughput and error rate of the side-channel attack against DNA-RM across different numbers of DRAM banks.
6 Mitigations for IMPACT
This section discusses the effectiveness of three different techniques to mitigate IMPACT: (i) An address mapping scheme that partitions memory banks across different processes, (ii) a memory controller policy that enforces closing the row buffer after every memory access, and (iii) a memory controller policy that enforces constant-time DRAM accesses to hide the timing differences from the attackers.
Memory Partitioning (MPR). Memory bank partitioning can serve as an effective countermeasure against main-memory-based covert channel attacks. By dividing the memory into separate banks, each with its own dedicated resources and access control, memory bank partitioning restricts the interactions between different memory regions. This partitioning prevents unauthorized communication channels from crossing the boundaries of the designated banks. In the context of a IMPACT, memory bank partitioning limits the sender’s ability to transmit information across different memory banks and the receiver’s ability to observe contention in different memory banks. By segregating the memory into distinct banks, the covert channel’s ability to exploit timing differences is significantly constrained. However, statically partitioning DRAM resources across applications disables data sharing, which can significantly (i) increase data duplication, (ii) increase memory underutilization, and (iii) lead to frequent swapping from the disk when the memory space is full.
Closed Row Policy (CRP). Enforcing a memory controller policy that closes a DRAM row after each operation can be an effective measure to mitigate and eventually stop timing covert channel attacks. By closing the DRAM row after each operation, the memory controller ensures that subsequent memory accesses, whether legitimate or covert, cannot access the same row without incurring the overhead of row activation. In the context of IMPACT, which heavily relies on the difference in memory access latency to transmit information, closing the DRAM row after each operation disrupts the timing patterns exploited by the attacker. However, as we show in Section 6, enforcing a closed row policy leads to significant performance overheads, even for workloads with low data locality, as every DRAM access will lead to a DRAM row buffer miss.
Constant-Time DRAM Accesses (CTD). One way to mitigate main-memory-based covert channel attacks is to enforce a constant-time DRAM access policy. When a memory access is issued, the memory controller waits for a fixed amount of time before returning the data to the processor. This fixed amount of time is equal to the worst-case DRAM access latency. By enforcing a constant-time DRAM access policy, the memory controller ensures that all memory accesses take the same amount of time, regardless of the accessed memory location. In the context of IMPACT, where the communication relies on the timing or side effects of memory accesses, enforcing a constant-time DRAM access policy obfuscates the timing patterns exploited by the attacker. However, as we show in Section 6, enforcing a constant-time DRAM access policy leads to significant performance overheads as every DRAM access costs the worst-case DRAM access latency.
Performance Evaluation. To better understand the effectiveness of the countermeasures, we evaluate the performance of CRP and CTD compared to the baseline open-row policy for 5 multi-programmed graph applications which (i) use and share the input graph (i.e., accesses to the same DRAM bank) and (ii) experience a mix of low and high data locality, i.e., the number of last-level cache misses per kilo instructions. We use a 2-core system where each core runs a different instance of the same application (e.g., core 0 runs an instance of PageRank and core 1 the same instance of PageRank using the same input). We evaluate five workloads from GraphBIG benchmark suite [123]: Betweenness Centrality (BC; Data set size: 7.8 GB; LLC MPKI: 0.57), Breadth-First search (BFS; 7.6 GB; 38.59), Connected components (CC; 7.9 GB; 45.2), Triangle counting (TC; 8.2 GB; 5.08), PageRank (PR; 7.0 GB; 1.86).
Figure 18 shows the performance overhead of CRP and CTD compared to the open-row policy for all 5 workloads. We make two key observations. First, CTD leads to the highest performance degration, 26% on average, compared to the open row policy. Especially, in workloads like PR and BC, CTD leads to a performance degradation of 37% and 45% respectively. Second, CRP leads to a performance degradation of 15% on average compared to the open row policy but leads to low performance overheads in workloads that do not benefit from the open row policy like TC, CC and BFS. We conclude that existing countermeasures lead to high performance overheads and there is a dire need to come up with techniques that improve security with low performance overhead.
Figure 18: Performance overhead of Closed-Row policy and Constant-Time policy compared to Open-Row policy.
7 Related Work
To the best of our knowledge, this is the first work that evaluates the implications of PiM on timing attacks.
Security Implications of PiM Systems. A set of prior works [127, 128, 129] discuss specific security vulnerabilities in PiM systems. Arafin et al. [127] survey and theoretically analyze security vulnerabilities in PiM systems, with a focus on five specific attack scenarios. In this work, we quantitatively and qualitatively analyze and evaluate main memory-based covert and side channels in detail. Wang et al. [129] extracts neural network architectural information from an RRAM accelerator using power side channels. Similarly, Ensan et al. [128] proposes an attack revealing implemented intellectual property in RRAM-based PiM architectures using power/timing side channels. In contrast, in this work, we demonstrate how to establish high-throughput timing covert channels using PiM operations through software.
Timing Attacks. A significant body of work has focused on exploring timing attacks [130, 131]. These attacks involve exploiting variations in the execution time of memory and processor core operations to infer sensitive information. Researchers have investigated various timing attack techniques, such as cache-based attacks [132, 133, 134, 121, 135, 136, 137, 138, 139] to uncover vulnerabilities in traditional systems. There is an extensive body of research related to traditional covert-channel attacks, including Flush+Reload [136], where the sender flushes a shared cache line and monitors the access time by the receiver to infer transmitted data. Flush+Flush [135] utilizes cache flush operations to probe for the presence of specific data patterns. Prime+Probe [140] exploits cache set conflicts to communicate information between sender and receiver. Streamline [121] is a flush-less cache covert channel attack that enables asynchronous communication. We compare IMPACT to Streamline in Section 5.3 and show that IMPACT achieves significantly higher throughput.
Other Main Memory-based Timing Attack Mitigations. We already extensively discuss and evaluate mitigations that fundamentally eliminate the timing channel in Section 6. We also discuss practical solutions that restrict access to the timing channel through cache management methods and show that they are not applicable to PiM-based attacks. Another mitigation method is to restrict the usage of fine-grained timers to prevent attackers from exploiting the timing channel. A set of modern processors limits access to fine-grained timers such as Apple M1 cores [141]. Fine-grained timers are used in a wide range of user space applications, such as performance and power consumption monitoring, multithreading, synchronization and locking, optimizations on protocols such as TCP. Restricting access to fine-grained timers can also disable these applications or degrade their effects.
DAGguise [142] is a defense mechanism against memory timing side-channel attacks. DAGguise utilizes a memory access pattern representation obtained through offline profiling and shapes an application’s memory access pattern to eliminate information leakage. However, DAGguise does not mitigate row buffer-based timing channels as is and requires the memory controller to employ a closed row policy (evaluated in §6).
Enhancing System Security using PiM. A set of prior works proposes leveraging PIM to enhance the security of modern computing systems. SecNDP [143] proposes an encryption and verification mechanism for untrusted PiM devices that enables computation over encrypted data. SecNDP only focuses on protecting confidentiality and does not consider or analyze potential timing channels in contrast to IMPACT. Nelson et al.[144] propose offloading security critical tasks to near-memory compute units to avoid storing sensitive information in caches and eliminate cache-based side-channel attacks leaking them. In contrast, in this work, we demonstrate and evaluate the implications of PiM on main memory-based timing attacks. InvisiMem. [145] and ObfusMem [146] propose leveraging 3D-stacked memory to design secure processors that provide oblivious RAM equivalent guarantees.
8 Conclusion
The adoption of processing-in-memory (PiM) architectures has gained significant attention due to their ability to provide high performance and low energy consumption by reducing data movement bottlenecks. In this work, we introduce IMPACT, a new high-throughput main memory timing attack that takes advantage of the direct memory access provided by PiM architectures. IMPACT exploits the shared DRAM row buffer to 1) establish covert channels to communicate with other processes at high throughput and 2) leak security-critical information of victim processes based on their memory accesses. IMPACT covert channels leverage parallelism and eliminate cache bypassing steps, resulting in increased throughput compared to existing main memory-based covert channels. IMPACT side-channel attack targets a genomics application (i.e., DNA read mapping) and leaks the private characteristics of a user’s sample genome by leveraging PiM operations. The proposed countermeasures for IMPACT and the discussion on secure and robust PiM architecture designs highlight the need for addressing security concerns in PiM systems.

### v3

@@
 Yet, the security of such architectures has not been thoroughly explored.
-The adoption of PiM solutions provides a new way to directly access main memory, which can be potentially exploited by malicious user applications.
+The adoption of PiM solutions provides a new way to directly access main memory, which malicious user applications can potentially exploit.
 We show that this new way to access main memory opens opportunities for high-throughput timing attack vectors that are hard-to-mitigate without significant performance overhead.
 We introduce IMPACT, a set of high-throughput main memory-based timing attacks that leverage characteristics of PiM architectures to establish covert and side channels.
-IMPACT enables high-throughput communication and private information leakage.
-To achieve this, IMPACT (i) eliminates expensive cache bypassing steps required by processor-centric main memory and cache-based timing attacks and (ii) leverages the intrinsic parallelism of PiM operations.
-First, we showcase two covert-channel attack variants that run on the host CPU and leverage PiM architectures to gain direct and fast access to main memory and establish high-throughput communication covert channels.
-Second, we showcase a side-channel attack on a DNA sequence analysis application that leaks the private characteristics of a user’s sample genome by leveraging PiM operations.
-Our results demonstrate that (i) our covert channels achieve up to 14.16 Mb/s communication throughput, which is 5.41$\times$ faster than the state-of-the-art main memory-based covert channels, and (ii) our side-channel attack allows the attacker to determine the properties of a sample genome at a throughput of 7.5 Mb/s with 96% accuracy.
-We discuss and evaluate several countermeasures for IMPACT to enable secure and robust PiM architectures.
+IMPACT enables high-throughput communication and private information leakage by exploiting the shared DRAM row buffer.
+To achieve high throughput, IMPACT (i) eliminates expensive cache bypassing steps required by processor-centric main memory and cache-based timing covert and side-channel attacks and (ii) leverages the intrinsic parallelism of PiM operations.
+We showcase two applications of IMPACT.
+First, we build two covert-channel attack variants that run on the host CPU and leverage different PiM approaches (i.e., processing-near-memory and processing-using-memory) to gain direct and fast access to main memory and establish high-throughput communication covert channels.
+Second, we showcase a side-channel attack that leaks private information of concurrently running victim applications that are accelerated with PiM.
+Our results demonstrate that (i) our covert channels achieve 12.87 Mb/s and 14.16 Mb/s communication throughput, respectively, which is up to 4.91$\times$ and 5.41$\times$ faster than the state-of-the-art main memory-based covert channels, and (ii) our side-channel attack allows the attacker to leak secrets with a low error rate.
+To avoid such covert and side channels in emerging PiM systems, we propose and evaluate three defenses that fundamentally eliminate the timing channel.
 11footnotetext: K. Kanellopoulos and F. N. Bostancı are co-primary authors.
 1 Introduction
-Data movement between the main memory and the processor is a prevalent bottleneck present in today’s computing systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
+Data movement between the main memory and the processor is a critical bottleneck present in today’s computing systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
 Hence, both industry and academia have invested significant efforts to bring computation closer to where data resides, creating a paradigm called processing-in-memory (PiM)  [13, 23, 8, 24, 25, 26, 11, 27, 19, 28, 1, 2].
-PiM solutions provide significant performance improvement and energy efficiency by alleviating data movement.
-PiM solutions generally take one of two approaches: (i) processing-near-memory (PnM), and (ii) processing-using-memory (PuM).
-First, PnM techniques offload computation to compute units near memory arrays (e.g., in the logic layer of 3D-stacked memory or within the memory chip) [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 21, 39, 40, 41, 42].
-During recent years, multiple real-world PnM systems have emerged, such as UPMEM [43], AxDIMM [41], AiM [44], HB-PNM [45] and FIMDRAM [46].
-At the same time, several academic works propose using general-purpose programmable PiM devices where the user application offloads instructions to the compute units inside the memory at fine-granularity [47] (e.g., PiM-enabled instructions [48]) or coarse granularity (i.e., full program [20, 49]).
-Second, PuM techniques propose exploiting the analog properties of memory operations and performing specialized PiM operations.
-Prior works improve the performance of workloads by providing bulk data copy and initialization, bitwise operations, and arithmetic operations using memory [50, 51, 52, 53, 54, 55, 56, 49, 57, 58, 59, 60, 61, 62, 63, 64].
-However, to the best of our knowledge, no prior work evaluates the security of PiM architectures against timing attacks.
-Analyzing and evaluating the security of PiM architectures is critical to prevent widespread vulnerabilities as PiM architectures become widely available.
-After a thorough analysis of existing PiM architectures, in this work, we show that the adoption of PiM solutions creates opportunities for high-throughput main memory-based timing attacks for two main reasons: (i) PiM provides direct access to main memory , and (ii) a large set of existing practical defense mechanisms are not applicable in PiM systems.
+PiM solutions provide significant performance improvement and energy efficiency by reducing data movement.
+PiM solutions take either of two approaches: (i) processing-near-memory (PnM), where the computation happens in a computation unit placed on the same chip as the memory array [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 21, 39, 40, 41, 42], which also appears in real-world products (e.g., UPMEM [43], AxDIMM [41], AiM [44], HB-PNM [45] and FIMDRAM [46]) and (ii) processing-using-memory (PuM) where the memory array performs the computation by leveraging the analog computation capability of DRAM cells [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62].
+Prior works on PnM and PuM clearly demonstrate that these emerging architectures promise great performance and energy efficiency benefits.
+In addition, the existence of real products that implement such techniques [43, 41, 44, 45, 46] indicates that these techniques are likely to be used in many real systems in the near future.
+Therefore, it is important to investigate the security aspect of such techniques and implement countermeasures to avoid potential widespread security vulnerabilities that their deployment can cause.
+Unfortunately, no prior work analyzes and evaluates the security of emerging PiM architectures against timing covert and side-channel attacks.
+In this work, we analyze existing PiM architectures and show that the adoption of PiM solutions creates opportunities for high-throughput main memory-based attacks that are hard-to-mitigate without significant performance overheads.
+This is possible due to two main reasons: (i) PiM architectures provide direct access to main memory, which is essential for their performance and energy efficiency benefits, and (ii) a large set of existing practical defense mechanisms are not applicable in these PiM architectures.
 1.
 Direct Access to Main Memory.
-Timing attacks in PiM architectures become easier to exploit with the direct memory access provided to user space applications by these architectures.
+High-throughput main memory-based timing side- and covert-channel attacks need to access main memory with high frequency to observe or change the states of shared main memory structures (e.g., DRAM row buffers).
+In today’s systems, it is difficult to access main memory with high frequency due to the deep cache hierarchies that filter the memory requests of attacker processes.
+[^1] We provide analysis as to why prior attack primitives that bypass caches do not provide fast and reliable direct main memory access in §3.2.
 The adoption of PiM techniques changes how applications and computation units access main memory: (i) due to area and thermal dissipation limitations, compute units in PnM architectures have small or no caches that allow the user space applications to bypass caches more easily (i.e., with fewer memory accesses), and (ii) PuM architectures provide the ability to offload specialized PiM instructions (e.g., bulk data copy and initialization, and bitwise/arithmetic operations) to the main memory that completely bypass the deep cache hierarchies of modern processors.
+As a result, PiM architectures provide user space applications with fast and reliable direct main memory access that can be exploited by attackers for high-throughput timing covert- and side-channel attacks.
 2.
@@
 A fundamental solution to memory-based timing attacks is to eliminate the timing channel itself.
-Prior mitigations explore fundamental solutions such as memory partitioning and enforcing constant-time memory accesses.
-However, these mitigations incur significantly high performance overheads (as shown in Section 6).
-Therefore, practical mitigations and detection mechanisms restricting access to timing channels are explored.
-Several prior works [65, 66, 67, 68] propose cache management-based methods such as restricting the access to cache management instructions (e.g., clflush instruction in x86 systems [65]) and detecting timing attacks based on cache access patterns (e.g., a high number of cache misses can indicate a timing attack) [66, 67, 68].
+Prior defenses explore fundamental solutions such as memory partitioning and enforcing constant-time memory accesses.
+However, these defenses incur significantly high performance overheads (as shown in Section 6).
+Therefore, several prior works [63, 64, 65, 66] explore practical defenses and detection mechanisms that restrict access to timing channels with cache management-based methods, such as restricting the access to cache management instructions (e.g., clflush instruction in x86 systems [63]) and detecting timing attacks based on cache access patterns (e.g., a high number of cache misses can indicate a timing attack) [64, 65, 66].
 These mechanisms are inapplicable to PiM-based attacks because these attacks completely bypass the cache hierarchy.
-Our goal is to develop a set of high-throughput and hard-to-mitigate timing side and covert-channel attacks in PiM architectures.
-We expect and hope that IMPACT will bring attention to security issues in emerging PiM architectures and result in security awareness in PiM architecture designs.
-In this work, we introduce IMPACT, a set of high-throughput main memory-based timing attacks that leverages PiM architectures to gain direct and fast access to main memory, establish high-throughput communication covert channels and leak secrets of concurrently running applications.
-IMPACT leverages the direct access, provided by PiM operations, to the shared DRAM row buffer and achieves high communication and information leakage throughput by (i) eliminating expensive cache bypassing steps required by other main memory- and cache-based timing cover-channel attacks and (ii) leveraging the intrinsic parallelism of PiM operations.
-Key Challenge and Limitations of Prior Work.
-For successful main memory-based timing side- and covert-channel attacks, attacker processes (or sender and receiver processes for covert-channel attacks) need to access the main memory with high frequency to observe or change the states of the shared main memory structures.
-In traditional architectures, it is difficult to access the main memory with high frequency due to the deep cache hierarchies that filter the memory requests of attacker processes.
-We analyze (Section 3.2) the effectiveness of the attack primitives that bypass caches and provide direct access to the main memory to leverage the main memory timing channel: clflush instruction, eviction sets, (remote) direct memory access, (R)DMA, requests, non-temporal memory hints (e.g., movnti instruction [69]), and PiM operations.
-None of these primitives used in traditional architectures provide fast and reliable direct memory access at the same time.
-In contrast, PiM operations provide fast and reliable direct memory access to the user space applications.
-Therefore, PiM architectures can be more amenable to high-throughput timing side- and covert-channel attacks compared to traditional architectures.
-IMPACT Covert-Channel Attack Variants.
-We showcase two different timing covert-channel attack variants by exploiting PnM and PuM techniques.
-First, we build a PnM-based covert channel, IMPACT-PnM, that uses PiM-enabled instructions [48] to directly access the main memory.
+Based on our analysis, we introduce IMPACT, a set of high-throughput In-Memory Processing-based timing Attacks that leverage the direct and fast main memory accesses enabled by PiM architectures.
+Using the direct access via PiM operations, IMPACT observes and exploits the shared DRAM row buffer.
+We demonstrate that IMPACT achieves high communication and information leakage throughput by (i) eliminating expensive cache bypassing steps required by other main memory- and cache-based timing covert- and side-channel attacks and (ii) leveraging the intrinsic parallelism of PiM operations.
+We showcase this finding on two covert- and one side-channel attacks.
+To protect future systems against IMPACT covert and side channels, we propose three defense mechanisms that fundamentally eliminate the channel itself.
+IMPACT-PnM Covert Channel Variant.
+We build a PnM-based covert channel that uses PiM-enabled instructions [67] to directly access the main memory.
 The sender bypasses the locality monitor employed by the system and offloads PiM-enabled instructions for computation directly inside the main memory.
 The receiver measures the time it takes to access a specific set of rows to retrieve the message encoded with row buffer hits and conflicts.
-IMPACT-PnM achieves 12.87 Mb/s communication throughput, which is up to 5.79$\times$ faster than the state-of-the-art main memory-based covert-channel attack [70].
-Second, we present IMPACT-PuM, a PuM-based covert channel that leverages the inherent parallelism in main memory using bulk copy and initialization techniques, specifically RowClone [55], to quickly transmit messages through the shared row buffer states.
-This way, IMPACT-PuM transmits 16-bit large messages 14$\times$ faster than IMPACT-PnM and achieves communication throughput of 14.16 Mb/s (up to 6.38$\times$ faster than the state-of-the-art [70]).
+IMPACT-PnM achieves 12.87 Mb/s communication throughput, which is up to 4.91$\times$ faster than the state-of-the-art main memory-based attack [68].
+IMPACT-PuM Covert Channel Variant.
+We build a PuM-based covert channel that leverages the inherent parallelism in main memory using bulk copy and initialization techniques, specifically RowClone [52], to quickly transmit messages through the shared row buffer states.
+This way, IMPACT-PuM transmits 16-bit large messages 14$\times$ faster than IMPACT-PnM and achieves communication throughput of 14.16 Mb/s (up to 5.41$\times$ faster than the state-of-the-art [68]).
 IMPACT Side-Channel Attack.
-We showcase a side-channel attack on a DNA sequence analysis application: DNA read mapping.
-Preserving the privacy of DNA sequences is crucial, especially when the sequences are used for medical purposes.
-DNA read mapping finds the matching locations of short DNA fragments of the sample genome in the reference genome, which acts as a template genome to construct the complete genome of a given sample.
-Our side-channel attack leverages PiM operations and leaks the private characteristics of a user’s sample genome by observing the memory access patterns of the DNA read mapping application.
-Our evaluation demonstrates that our side-channel attack enables determining the properties of a sample genome at a throughput of 7.5 Mb/s with 96% accuracy.
+We build a PnM-based IMPACT side-channel attack that exploits PiM operations to leak private information belonging to a concurrently running application based on its memory access characteristics.
+We showcase this side-channel attack on read mapping, one of the fundamental applications in DNA sequence analysis where preserving the privacy of the human genomic data is crucial [69, 70, 71, 72, 73].
+Our side-channel attack leverages PiM operations and leaks the private characteristics of a user’s sample genome by observing the memory access patterns of the read mapping application.
+Our evaluation demonstrates that our side-channel attack leaks the properties of a sample genome at a throughput of 7.57 Mb/s with 96% accuracy.
 Defense Mechanisms.
-Based on our security analysis, we discuss three and evaluate two potential defense mechanisms that can mitigate IMPACT attacks by eliminating the timing channel.
-We discuss the performance versus security trade-offs of these countermeasures.
-We demonstrate that the two best-performing defense mechanisms lead to high performance overheads of 25% and 16% on average across five data-intensive workloads.
+Based on our security analysis, we discuss and evaluate potential defense mechanisms that can mitigate IMPACT attacks by eliminating the timing channel.
+We discuss the performance versus security trade-offs of these defenses.
+We demonstrate that the two best-performing defense mechanisms lead to high performance overheads of 26% and 15% on average across five data-intensive workloads.
 This paper makes the following contributions:
-  • We present IMPACT, the first timing attack in PiM systems. We leverage the direct memory access enabled by PiM operations to establish high-throughput and reliable timing covert channels.
-  • We showcase two IMPACT variants by exploiting PnM and PuM techniques: IMPACT-PnM and IMPACT-PuM.
-IMPACT-PnM allows the sender and the receiver to exchange messages by using PiM instructions.
-IMPACT-PuM further leverages the parallelism provided by bulk copy PuM techniques to improve the throughput.
-We extensively evaluate both IMPACT-PnM and IMPACT-PuM and show that they significantly improve communication throughput compared to the state-of-the-art main-memory-based covert channel.
-  • We showcase an IMPACT side-channel attack on DNA read mapping by leveraging PnM instructions.
+  • To the best of our knowledge, this is the first work that analyzes and evaluates the security of emerging PiM architectures against timing covert and side-channel attacks.
+We introduce IMPACT, a new set of high-throughput main memory-based timing attacks that leverage PiM operations to gain fast and reliable access to main memory.
+  • We showcase two IMPACT covert channel variants, IMPACT-PnM and IMPACT-PuM that exploit PnM and PuM techniques, respectively.
+First, IMPACT-PnM allows the sender and the receiver to exchange messages using PiM-enabled instructions.
+Second, IMPACT-PuM further leverages the parallelism provided by bulk copy PuM techniques to improve the communication throughput.
+We show that both attacks outperform the state-of-the-art main-memory-based covert-channel attacks in terms of communication throughput.
+  • We showcase an IMPACT side-channel attack on read mapping by leveraging PnM instructions.
 We show that our side-channel attack enables the attacker to leak sensitive information on the victim’s sample genome with high throughput and a low error rate.
-  • We evaluate two defense mechanisms to protect the system against IMPACT attacks caused by the adoption of PiM solutions.
-We discuss the trade-offs of these mechanisms to aid the adoption of robust and secure PiM architectures.
+  • We evaluate and discuss the trade-offs of three defense mechanisms to protect PiM-enabled systems against IMPACT’s covert- and side-channel attacks.
 2 Background
@@
 Organization.
-Figure 1 illustrates a DRAM module’s hierarchical organization.
-A DRAM chip is organized into multiple DRAM banks, each of which shares a common I/O circuitry in a DRAM bank [71, 72, 73, 74, 75].
-DRAM cells are organized as a two-dimensional array to form a bank and are addressed by its row and column [76, 77, 22, 78].
-The DRAM cell is connected to a bitline that is used to access the data stored in the cell and a wordline that controls access to the cell.
-Each DRAM cell in a DRAM row is connected to a common wordline via its access transistors.
-A bitline connects a column of DRAM cells to a DRAM sense amplifier to read or write data.
-A row of sense amplifiers is called a row buffer [56, 79, 80, 81, 82, 55].
-At the lowest level of the hierarchy, a single DRAM cell comprises 1) a storage capacitor that stores a single bit of data encoded using the charge level in the capacitor and 2) an access transistor that is used to read from and write to the storage capacitor [83, 84, 85, 86, 87].
+Fig.
+1 shows the hierarchical organization of a modern DRAM-based main memory.
+The memory controller connects to a DRAM module over a memory channel.
+A module contains one or more DRAM ranks that time-share the memory channel.
+A rank consists of multiple DRAM chips that operate in lock-step.
+Each DRAM chip contains multiple DRAM banks that can be accessed independently.
+A DRAM bank is organized as a two-dimensional array of DRAM cells, where a row of cells is called a DRAM row.
+A DRAM cell consists of 1) a storage capacitor, which stores one bit of information in the form of electrical charge, and 2) an access transistor, which connects the capacitor to the row buffer through a bitline controlled by a wordline.
 Figure 1: Organization of a typical modern DRAM module.
 Operation.
-The memory controller serves memory access requests.
-To serve memory requests, the memory controller issues DRAM commands, e.g., row activation ($ACT$), bank precharge ($PRE$), data read ($RD$), data write ($WR$), and refresh ($REF$).
-Between consecutive DRAM commands, the memory controller must obey standardized timings to ensure correct DRAM operation.
-For read and write requests, the memory controller first issues an $ACT$ command to the corresponding bank, which opens the corresponding row.
-The data in the row is fetched to the row buffer, where it is temporarily stored.
-The data in the row buffer becomes accessible by a $RD$/$WR$ command when the bitline and cell voltage are restored to a certain level.
-Hence, the memory controller should wait at least the activation latency ($t_{RCD}$) between an $ACT$ and a $RD$/$WR$ command.
-All changes are propagated to the DRAM cells in the open row.
+To perform a read or write operation, the memory controller first needs to open a row, i.e., copy the data of the cells in the row to the row buffer.
+To open a row, the memory controller issues an activation command to a bank by specifying the row address to open.
+After activation is complete, the memory controller issues either read and write commands to read or write a DRAM word within the activated row.
 Subsequent accesses to the same row can be served quickly without fetching the data to the row buffer, creating a row hit [74].
-However, in the case where the target row is not in the row buffer, the memory controller issues a $PRE$ command to close the already open row and issues an $ACT$ command to fetch the target row, creating a row conflict/miss.
-The time interval for issuing 1) a $PRE$ after an $ACT$ must be at least restoration latency ($t_{RAS}$) and 2) an $ACT$ after a $PRE$ must be at least the precharge latency ($t_{RP}$).
+To access data from another DRAM row in the same bank, the memory controller must first close the currently open row by issuing a bank precharge command, creating a row conflict/miss.
 2.2 Processing-in-Memory
-Processing-in-memory (PiM) is a promising paradigm that moves computation closer to where the data resides, alleviating the data movement bottleneck between memory and the processor [13, 23, 8, 24, 25, 26, 11, 27, 19, 28, 1, 2].
-There are two main approaches to PiM: 1) processing-near-memory (PnM) and 2) processing-using-memory (PuM).
-Processing Near Memory (PnM).
-In processing-near-memory (PnM) computation is placed near the memory arrays.
-PnM takes advantage of the computation capability of conventional memory controllers or the logic layer(s) of the 3D-stacked memory technologies [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 21, 39, 40, 41, 42].
-PnM logic can execute portions of applications (e.g., instructions [48, 32], functions [28], whole application threads [20, 49]) depending on the PnM architecture.
-PnM computation units typically employ small caches due to 1) area constraints in the PnM logic and 2) the ineffectiveness of caches in improving the performance of a wide range of memory-intensive applications.
-Therefore, various PnM systems [32, 48, 28] execute the memory-intensive regions of applications on the PnM computation units, while executing the more cache-friendly regions of applications on the host CPU with a deeper cache hierarchy.
-Processing Using Memory (PuM).
-Processing Using Memory (PuM) expands the functionality of the memory arrays with minimal changes so that they can perform computation [50, 51, 52, 53, 54, 55, 56, 49, 57, 58, 59, 60, 61, 62, 63, 64].
-This approach leverages the properties and operational principles of memory cells and arrays to perform computation, without the need for dedicated logic unit added to the memory system.
-PuM architectures take advantage of the high bandwidth of the memory arrays.
-2.3 Microarchitectural Timing Attacks
-Microarchitectural timing attacks exploit the observable state changes in microarchitecture to build covert channel attacks that can propagate messages between processes and leak secrets of concurrently running processes.
-These attacks measure the time to complete an operation that depends on the shared structures in the system (e.g., accessing a memory location by going through the shared caches) to infer the states of these shared structures.
-Main Memory Timing Attacks.
-Existing main memory timing attacks exploit several shared structures in the main memory to leak data, and one such structure is the row buffer.
-The row buffer acts as an internal cache inside DRAM banks.
-Thus, an attacker can understand if a row is recently accessed (i.e., already in the row buffer) or not by measuring access to that row.
-Based on this observation, prior works propose several attacks, revealing the DRAM address mapping and leaking keystrokes.
-DRAMA [70] (i) builds a covert channel that exchanges messages between processes using the row buffer conflicts and hits, and (ii) showcases a side-channel attack that can leak keystrokes in a real system.
-Similarly, many works [70, 89, 90] exploit row buffer-based timing differences to determine DRAM address mappings.
-Other attacks based on the main memory exploit either the memory bus [91], or observations of latencies incurred by other memory-related operations such as memory deduplication [92, 93], and memory compression [94, 95, 96, 97, 98, 99, 100, 101] and decompression [102].
+Processing-in-memory (PiM) moves computation closer to where the data resides, alleviating the data movement bottleneck between memory and the processor [13, 23, 8, 24, 25, 26, 11, 27, 19, 28, 1, 2].
+There are two main PiM approaches: 1) processing-near-memory (PnM) and 2) processing-using-memory (PuM).
+PnM place compute logic near memory arrays (e.g., processing capability in the memory controller and logic layer(s) of 3D-stacked memory [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 21, 39, 40, 41, 42]).
+PnM logic can execute application regions (e.g., instructions [67, 32], functions [28], application threads [20, 54]) depending on the design.
+PnM computation units typically employ small caches due to 1) area constraints and 2) the ineffectiveness of caches in improving the performance of a wide range of memory-intensive applications.
+Therefore, various systems [32, 67, 28] offload the memory-intensive application regions on the PnM units, while executing more cache-friendly application regions on the host CPU with a deeper cache hierarchy.
+PuM uses analog operational properties of the memory circuitry (e.g., DRAM) to enable massively parallel in-memory computation [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62].
+2.3 Main Memory Timing Attacks
+Microarchitectural timing attacks exploit the observable microarchitectural state changes to propagate messages between processes and leak secrets.
+These attacks measure the time to complete an operation that depends on a shared system structure (e.g., accessing a memory location by going through the shared caches) to infer the state of the shared structure.
+Main memory timing attacks exploit several shared structures in the main memory.
+One such structure is the DRAM row buffer that acts as an internal cache inside DRAM banks.
+An attacker can understand if a row is recently accessed (i.e., already in the row buffer) or not by measuring the time it takes to access that row.
+Using this observation, DRAMA [68] builds a side-channel attack to leak keystrokes in a real system.
+Several prior works [68, 75, 76, 77, 78] leak DRAM address mapping functions.
+Other main memory-based attacks exploit the memory bus [79], and memory deduplication [80, 81] and (de)compression [82, 83, 84, 85, 86, 87, 88, 89, 90] latencies.
 3 A Case for PiM-based Timing Attacks
-The adoption of processing-in-memory (PiM) architectures has been gaining momentum in recent years because they provide high performance and low energy consumption by alleviating the data movement bottleneck.
-Therefore, it is timely and critical to analyze and evaluate the security of PiM architectures to prevent any of their vulnerabilities from causing widespread susceptibility as PiM architectures become widely available.
-Unfortunately, no prior work investigates and evaluates the security challenges of new and emerging PiM architectures.
+The adoption of processing-in-memory (PiM) architectures has been gaining momentum in recent years because they provide high performance and low energy consumption by reducing data movement.
+Therefore, it is timely and critical to analyze and evaluate the security of PiM architectures to prevent any vulnerabilities from causing widespread susceptibility as PiM architectures become widely available.
+Unfortunately, no prior work analyzes and evaluates the security of emerging PiM architectures against timing covert and side-channel attacks.
 After a thorough analysis of existing PiM architectures, we show that the adoption of PiM solutions creates opportunities for high-throughput main memory-based timing attacks for two main reasons.
-First, the adoption of PiM solutions in the forms of processing-near-memory (PnM) and processing-using-memory (PuM) exposes the internal states of the main memory (e.g., row buffer states of DRAM banks) by providing direct access to the user space applications.
+First, the adoption of PiM solutions provides user space applications with fast and reliable direct access to main memory.
 Second, timing attacks in PiM architectures can be more difficult to detect and mitigate due to commonly employed mechanisms, like cache monitoring mechanisms that track the cache miss ratio during runtime to detect anomalies [64, 65, 66], being either inapplicable or performance-inefficient in PiM architectures.
@@
 A DRAM row buffer acts as a direct-mapped cache inside a DRAM bank and is shared across applications that can access main memory.
-Accessing a row that is already cached in the row buffer results in a significantly lower latency because the memory controller does not need to issue additional commands, $PRE$ to precharge the open row and $ACT$ to fetch the target row to the row buffer.
-Therefore, it is possible for a user space application to infer whether a row is accessed recently by measuring the time it takes to access the row.
-To demonstrate the latency gap between different states of the row buffer, we evaluate microbenchmarks that create memory accesses that result in row buffer hits and conflicts, and we measure the latencies observed by the user space application.
-Figure 2 shows the breakdown of memory access latencies categorized into two classes: row buffer hits and conflicts.
-Figure 2: The distribution of memory access latencies.
-We observe that the average latency gap between a row buffer conflict and a row buffer hit is 74 CPU cycles.
-Since this latency gap is large enough to detect, it is possible for user space applications that share the main memory to exploit the row buffer as a covert channel to exchange messages.
-Our results align with the observations of prior works [70].
+Accessing a row that is already in the row buffer results in a significantly lower latency because the memory controller does not need to issue additional commands to precharge the open row and fetch the target row to the row buffer.
+Therefore, a user space application can infer whether a row has been accessed recently by measuring the time it takes to access the row.
+To demonstrate the latency gap between different row buffer states, we evaluate microbenchmarks that create memory accesses that result in row buffer hits and conflicts.
+We observe that a row buffer conflict takes 74 CPU cycles more than a hit, which is large enough to detect.
+So, a user space application can exploit the row buffer as a timing covert and side channel.
+Our results align with the observations of prior works [68, 77, 78].
 3.2 Existing Main Memory Attack Primitives
-Main memory timing attacks are powerful as they can be used to gather information on applications that share main memory and establish covert channels to exchange messages between different processes.
-However, these covert-channel attacks are limited in throughput in traditional architectures for two main reasons: (i) accessing a memory location in main memory requires expensive cache bypassing techniques due to frequently accessed data being cached to improve performance, and (ii) even when bypassing the cache hierarchy, deep cache hierarchies employed in modern processors incur an additional high lookup latency on memory requests that are eventually sent to main memory.
-There are several ways to bypass the cache hierarchy in traditional systems by using one or more of four attack primitives: (i) using specialized instructions to flush a cache line (e.g., clflush), (ii) using cache eviction sets (i.e., by creating many conflicting memory requests to replace the cache line with newly-fetched cache lines), (iii) memory access requests generated by direct memory access (DMA) and remote direct memory access (R-DMA) engines, and (iv) non-temporal memory hints (e.g., movnti).
-However, when used in a main memory-based timing covert channel attack, these methods do not provide fast and reliable communication.
+Main memory timing attacks are powerful as they can be used to gather information on all applications that share main memory.
+However, these covert-channel attacks have limited throughput in traditional architectures for two main reasons: (i) accessing a memory location in main memory requires expensive cache bypassing techniques due to frequently accessed data being cached to improve performance, and (ii) even when bypassing the cache hierarchy, deep cache hierarchies employed in modern processors incur an additional high lookup latency on memory requests.
+There are several ways to bypass the cache hierarchy by using one or more of four attack primitives: (i) using specialized instructions to flush a cache line (e.g., clflush), (ii) using cache eviction sets (i.e., by creating many conflicting memory requests to replace the cache line with newly-fetched cache lines), (iii) memory access requests generated by (remote) direct memory access, (R)DMA, engines, and (iv) non-temporal memory hints (e.g., movnti).
+However, no attack primitive provides both fast and reliable communication.
 Table 1: Efficiency and Effectiveness of Attack Primitives
@@
 Table 1 shows a comprehensive comparison of these four attack primitives in terms of these properties.
-First, specialized cache line flush instructions have two main disadvantages that hurt the efficiency and effectiveness of attacks that use them: 1) they put the latency of writing back a cache line to the main memory on the critical path to access a memory location, and 2) in some modern systems, these instructions are privileged and cannot be used by user space applications.
-Second, eviction sets require a large number of load requests proportional to the number of ways in the caches, which scales with modern designs.
+First, specialized cache line flush instructions have two main disadvantages that hurt the efficiency and effectiveness of attacks that use them: 1) they put the write-back (to the main memory) latency of a cache line on the critical path to access a memory location, and 2) in some modern systems, these instructions are privileged and cannot be used by user space applications.
+Second, eviction sets require many load requests proportional to the number of ways in the caches, which scales with modern designs.
 Therefore, the attacks that use them have limited throughput.
-While the ISA guarantees the instructions used to create the eviction sets, the eviction set is not guaranteed to evict the target cache line (due to cache characteristics such as cache replacement policy, cache prefetchers etc.).
+While the instructions used to create the eviction sets are part of the ISA, the eviction set is not guaranteed to evict the target cache line due to the cache replacement policy, cache prefetchers, and other design choices that are not transparent to the attacker.
 Third, DMA and R-DMA engines incur high overheads due to deep software stacks.
 As a result, attacks using this primitive need to utilize states with larger timing differences to be distinguishable.
-An example attack using this primitive[103] showcases a covert channel attack based on memory contention.
+An example attack using this primitive[91] showcases a covert-channel attack based on memory contention.
 However, the basis of this type of covert channel is not as fine-grained as row buffer-based timing differences.
-Therefore, the communication throughput these covert channels provide is limited (in the order of Kb/s).
+Therefore, the communication throughput of these covert channels is limited (in the order of Kb/s).
 Fourth, the implementation of non-temporal memory hints varies across systems, as system designers are not bound to a specific approach.
 Consequently, the ISA does not guarantee that memory accesses related to non-temporal hints will consistently bypass caches.
-In contrast, PiM architectures provide fast and reliable direct memory access to user space applications that can be used to establish a covert-channel attack.
+In contrast, PiM operations provide fast and reliable direct memory access to user space applications, which can be exploited for high-throughput covert- and side-channel attacks.
 3.3 High Throughput Main Memory-based Covert Channels
-To demonstrate the impact of the latencies induced by deep cache hierarchies (i.e., cache access and cache eviction latencies) on communication throughput, we simulate a realistic system (described in Section 5.1) that establishes a state-of-the-art covert channel [70].
+To demonstrate the impact of the deep cache hierarchy latencies (i.e., cache access and eviction latencies) on an attack’s throughput, we simulate a realistic system (described in §5.1), establishing a state-of-the-art covert channel [68].
 Attack Scenarios.
@@
 First, we evaluate a baseline attack that relies on cache bypassing using cache eviction sets.
-To transmit a bit through the row buffer, the baseline attack first evicts the corresponding cache line and then, it fetches the target row into the row buffer using an additional memory request.
+To transmit a bit through the row buffer, the baseline attack first evicts the corresponding cache line, and then it fetches the target row into the row buffer using an additional memory request.
 The baseline attack accomplishes cache line eviction by issuing N memory requests to the target cache set in an N-way cache.
 Note that the actual eviction latency can be much higher due to a larger number of memory accesses required to bypass more complex replacement policies employed in today’s systems or due to address translation overheads.
-Second, we evaluate a direct memory access attack that can directly access the main memory.
-The direct memory access attack issues a memory request to fetch the data to the row buffer and does not need any cache access or eviction steps.
-For both attacks, we measure the time it takes to send a bit through the row buffer including cache access and cache eviction latencies, and with encoding logic-1 values as row buffer hits and logic-0 as row buffer conflicts.
-We calculate the communication throughput based on the average time needed to propagate bits from the sender to receiver.
-We analyze the impact of (i) the LLC size, and (ii) the number of LLC ways on communication throughput to show the limitations in current and future systems with different cache designs.
+Second, we evaluate a direct memory access attack that issues only one memory request to fetch the data to the row buffer without any cache access or evictions.
+For both attacks, we measure the time it takes to send a bit through the row buffer, including cache access and eviction latencies, with encoding logic-1 values as row buffer hits and logic-0 as row buffer conflicts.
+We calculate the communication throughput based on the average time needed to propagate bits from the sender to the receiver.
+We analyze the impact of (i) the LLC size and (ii) the number of LLC ways on communication throughput to show the limitations in current and future systems with different cache designs.
 Impact of LLC Size on Communication Throughput.
-Figure 3 shows the throughput of baseline and direct memory access attacks.
+Fig.
+2 shows the impact of the LLC size on (i) the throughput of baseline and direct memory access attacks and (ii) the latency of a cache line eviction.
 The x-axis shows the cache configuration with increasing last-level cache (LLC) sizes and a fixed number of ways (i.e., 16).
 To calculate the cache access latency with increasing LLC sizes, we followed the same methodology used in prior works [92].
-To better understand the throughput results, Figure 4 shows the latency of a cache line eviction of caches with corresponding sizes.
-Figure 3: Impact of the LLC size on the data leakage throughput of covert channels.
-Figure 4: Latency of a cache line eviction in CPU cycles with increasing LLC size.
 We make two key observations.
 First, the communication throughput of the direct memory access attack is $11.27$ Mb/s across all LLC sizes, which is significantly greater than the throughput of up to $2.29$ Mb/s achieved by the baseline attack.
-Second, the baseline attack’s throughput decreases with increasing LLC sizes due to increasingly expensive cache access and eviction latencies.
-We observe that with increasing LLC size, the eviction latency grows significantly due to the increase in access latency.
-We conclude that with increasing LLC sizes, the cache access and eviction latencies grow significantly and significantly hinder the communication throughput of the baseline attack compared to the direct memory access attack which is independent of the LLC size.
+Second, the baseline attack’s throughput decreases with increasing LLC sizes since the cache access and eviction become progressively more expensive.
+With increasing LLC size, the cache access latency increases and, thus, amplifies the cost of eviction, which requires multiple cache accesses.
+Figure 2: Impact of the LLC size on the data leakage throughput of covert channels and the eviction latency.
 Impact of LLC Associativity on Communication Throughput.
-Figure 5 shows the communication throughput of baseline and direct memory access attacks.
+Fig.
+3 shows the impact of LLC associativity on (i) the communication throughput of baseline and direct memory access attacks and (ii) the eviction latency of a cache line.
 The x-axis shows the increasing number of LLC ways (from 2 to 128) while the total LLC size is fixed at 16 MB (modern processors typically employ 8 cores and a 2MB LLC slice per core).
 Increasing the number of LLC ways increases the number of memory requests required to evict a cache line from the cache.
-To better understand this effect, Figure 6 shows the latency of a cache line eviction across increasing numbers of LLC ways.
-We observe that the baseline attack’s communication throughput decreases significantly with increasing number of LLC ways due to the increased cache eviction latency.
-Figure 5: Impact of the number of LLC ways on the data leakage throughput of covert channels.
-Figure 6: Latency of evicting a set in CPU cycles with increasing number of LLC ways.
-We conclude that attacks that communicate with main memory through caches are greatly limited in throughput due to expensive cache lookup and eviction latencies, which worsen with increasing size and associativity.
+We observe that the baseline attack’s communication throughput decreases significantly with an increasing number of LLC ways due to the higher cache eviction latency.
+Figure 3: Impact of the number of LLC ways on the data leakage throughput of covert channels and the eviction latency.
+We conclude that attacks that communicate with main memory through caches are greatly limited in throughput due to expensive cache access and eviction latencies, which worsen with increasing size and associativity.
 However, a covert channel based on direct memory access can achieve high throughput independent from cache configurations as it does not require bypassing the cache hierarchy.
-Adoption of PiM solutions can fundamentally enable high-throughput timing attacks through main memory by providing direct access to the main memory to user space applications.
-Therefore, it is crucial to carefully consider the security implications of PiM architectures and to develop robust security measures to protect against potential attacks.
-Our goal is to develop a set of high-throughput and hard-to-mitigate timing side and covert-channel attacks in PiM architectures.
-We expect and hope that IMPACT will bring attention to security issues in emerging PiM architectures and result in security awareness in PiM architecture designs.
-4 IMPACT: A High-Throughput Main Memory Timing Attack
-This section introduces IMPACT, a set of high-throughput main memory timing attacks that takes advantage of the direct memory access enabled by the emerging processing-in-memory paradigm.
-IMPACT (i) operates independently of specialized instructions that flush data from the cache hierarchy and (ii) avoids the resource-intensive cache eviction steps typically required for main memory and cache-based timing attacks, achieving higher data leakage throughput than previous memory-based attacks.
-In this work, we present (1) two variants of IMPACT covert channel attack and (2) an IMPACT side-channel attack on DNA read mapping applications enabled by processing-near-memory (PnM) and processing-using-memory (PuM) solutions.
-We evaluate the communication throughput of IMPACT covert channels in Section 5.3 and the leakage throughput of IMPACT side channel in Section 5.4.
-4.1 IMPACT Covert-Channel Attacks
-4.1.1 PnM-based Covert-Channel Attack
+The adoption of PiM solutions can fundamentally enable high-throughput timing attacks through main memory by providing direct access to the main memory to user space applications.
+Therefore, it is crucial to carefully consider the security implications of PiM architectures and develop robust security measures to protect against potential attacks.
+4 IMPACT: A Set of High-Throughput Main Memory Timing Attacks
+This section introduces IMPACT, a set of high-throughput main memory-based timing attacks that take advantage of the direct memory access enabled by the emerging processing-in-memory paradigm.
+IMPACT achieves high communication and information leakage throughput by (i) eliminating expensive cache bypassing steps required by other main memory- and cache-based timing covert- and side-channel attacks and (ii) leveraging the intrinsic parallelism of PiM operations.
+In this work, we present two covert-channel and one side-channel attacks enabled by processing-near-memory (PnM) and processing-using-memory (PuM) solutions.
+4.1 PnM-based IMPACT Covert-Channel Attack
 Baseline PnM Architecture.
-In IMPACT-PnM, we consider PiM-Enabled Instructions (PEI)  [48] as the baseline PnM architecture.
-PEI expands the instruction set architecture (ISA) of the host processor with PIM-enabled instructions (PEIs), while keeping the existing programming model intact.
-A PEI is an instruction that can be executed either on the host processor or on the in-memory computation logic which is integrated near each DRAM bank of an HBM stack.
-PEI can also easily be adopted in other memory technologies since its design does not rely on specific features of the memory technology.
-We chose PEI as the baseline PnM architecture as it introduces a simple and effective model (i.e., offloading computation at the instruction granularity) for PnM, and our attack vector can be generalized for other PnM architectures with similar design components (e.g., FIMDRAM [46]).
-Figure 7 depicts the organization of PEI.
-PEI architecture consists of two key components.
-First, PEI integrates a PEI Computation Unit (PCU) near each DRAM bank  1 and inside the CPU  2.
+PiM-Enabled Instructions (PEI)  [67] is a PnM architecture that expands the instruction set architecture of the host processor with PIM-enabled instructions (PEIs).
+PEIs can be executed either on the host processor or on the in-memory computation logic.
+We choose PEI as the baseline PnM architecture as it introduces a simple and effective model for PnM, and our attack can be generalized for other PnM architectures with similar design components (e.g., FIMDRAM [46]).
+The baseline PEI architecture has two key components.
+First, it integrates a PEI Computation Unit (PCU) near each DRAM bank and inside the CPU.
 In-memory PCUs are shared among all the cores of the host processor and are responsible for executing PEIs near the respective DRAM bank.
-Second, PEI adds a PEI Management Unit (PMU)  3 integrated with the last-level cache of the host system to (i) monitor the data locality of application regions and (ii) map each region to host-side or memory-side PCUs based on their data locality.
-This way a PEI can be selectively executed on the host processor PCU (instead of the in-memory PCU) when large on-chip caches are beneficial for its execution.
-As mentioned in [105], if a locality monitor entry is allocated by a PIM operation, it is often too aggressive to consider it as having high locality on its first hit in the locality monitor.
-Therefore, they propose adding an ignore flag per entry to ignore the first hit to an entry allocated by a PIM operation.
-Such a design allows us to easily bypass the locality monitor even with a small virtual address range in our disposal.
-Figure 7: Overview of PEI’s architecture.
+Second, the PEI architecture has a PEI Management Unit (PMU) that monitors data locality in application regions and decides where to map each region (either to host-side or memory-side PCUs).
+This way, PEI architecture can execute regions with high locality on the host side to benefit from caches.
+PEI adds an ignore flag per locality monitor entry to ignore the first hit in the locality monitor [93].
+This is because if a locality monitor entry is allocated for a PIM operation, it is often too aggressive to consider the operation as having a high locality on its first hit.
+This bit allows IMPACT-PnM to easily bypass the locality monitor even with a small virtual address range at its disposal.
 Attack Overview.
-The PnM-based covert channel (IMPACT-PnM) leverages the direct memory access provided by PiM-enabled instructions, which bypasses the cache hierarchy.
-IMPACT-PnM exploits the timing difference between the latency of a DRAM row buffer hit and a conflict to exchange messages between the sender and the receiver.
-The sender transmits bits of the message to the receiver in multiple batches of M bits, where the value of each bit is encoded as interference (i.e., row buffer hits or conflicts) in the row buffer of a DRAM bank.
-We assume the sender and the receiver co-locate data in the same DRAM banks.
-[^1] If PiM is the main memory, then the sender can co-locate its data with the receiver process’s data by using memory massaging techniques, as used in many prior works [90, 70, 106, 107].
-Two processes can share data itself in various cases, such as databases and websites running on browsers.
-In this attack, we encode logic-1 as interference in the DRAM bank, and logic-0 as no interference.
-Figure 8 and Listing 1 show the end-to-end flow of IMPACT-PnM and all the steps involved in establishing the covert channel between the sender and the receiver.
-Step 1.
-The receiver issues PEI operations to initialize each DRAM bank by activating a predetermined row (Line 18).
-It is necessary that the cache lines of the activated row are not used before in order to bypass the locality monitor (PMU) that is employed in PEI-based systems.
-Then, the sender and the receiver synchronize on a barrier (Lines 5&20) which ensures that they start the communication turn at the same time.
-Step 2.
-The sender starts the communication turn by issuing PEI instructions to activate a row in each DRAM bank (Lines 7-12) or not based on the bit value that will be transmitted.
-To transmit a 0, as shown in Fig.
-8 ( 1) the sender does not activate any row in the corresponding DRAM bank to avoid causing interference in the row buffer (e.g., Row 0xB remains activated in the row buffer  2).
-To transmit a 1, as shown in Fig.
-8 ( 3) the sender activates a row that is different from the one activated by the receiver to cause interference in the corresponding DRAM bank (i.e., a row buffer conflict between Row 0xD and Row 0xE  4).
-At the end of the communication turn (after transmitting M bits), the sender first executes a memory fence (Line 14) to ensure that all the PEI instructions have been executed before the receiver starts probing the DRAM banks and then signals the receiver to start probing the DRAM banks (Line 15).
-Step 3.
-The receiver starts probing the DRAM banks by issuing PEI instructions to activate the same row as the one activated during initialization (Line 25).
-Then, the receiver measures the time it takes to complete the request (Line 27).
-If the latency of executing the PEI exceeds a predetermined hit threshold, the receiver senses the interference in the row buffer and decodes the bit as a 1 (Lines 28-29  5, 6).
-Otherwise, the access is considered a row buffer hit and the receiver decodes the bit as a 0 (Lines 30-31  7, 8).
-The sender and the receiver repeat the same steps for each batch until the entire message is transmitted.
+Before the attack, the sender and the receiver co-locate their data in the same DRAM banks.
+To achieve this, one process uses memory massaging techniques to place its data in the same bank as the other process in main memory, as used in many prior works [76, 68, 94, 95].
+IMPACT-PnM uses PEI operations and exploits the latency difference between a DRAM row buffer hit and a conflict in a shared DRAM bank to exchange messages between the sender and the receiver.
+The sender transmits bits of the message to the receiver in multiple batches of M bits, where the value of each bit is encoded as interference in the row buffer of a DRAM bank.
+In this attack, we encode logic-1 as interference in the DRAM bank (i.e., a row buffer conflict), and logic-0 as no interference (i.e., a row buffer hit).
+Fig. 4 and Listing 1 show the end-to-end flow of IMPACT-PnM and all the steps involved in establishing the covert channel between the sender and the receiver.
+Figure 4: Flow of the PnM-based Covert Channel Attack.
 1 Sender:
-2 message = [0: N-1]; //N banks and N-bit message
+2 message = [0: N-1]; //N banks and N-bit msg.
 3 batch_size = M; //M-bits per batch where M<N
@@
 21 for every batch:
-22 block_on_semaphore();//if semaphore==0, wait
+22 block_on_semaphore();
+23 //if semaphore==0, wait
 24 for each bit in batch_size;
@@
 Listing 1: Pseudocode of PnM-based Covert-Channel Attack.
-Figure 8: Flow of the PnM-based Covert Channel Attack.
+Step 1.
+The receiver issues PEIs to initialize each DRAM bank by activating a predetermined row (Line 18).
+The receiver either uses unused cache lines of the activated row or sets the ignore flag in the corresponding entry in order to bypass the locality monitor.
+Then, the sender and the receiver synchronize on a barrier (Lines 5 and 20), which ensures that they start the communication turn at the same time.
+Step 2.
+The sender starts the communication turn by issuing PEI instructions to activate a row in each DRAM bank (Lines 7-12) or not based on the bit value that will be transmitted.
+To transmit a 0 ( 1), the sender does not activate any row in the corresponding DRAM bank to avoid causing interference in the row buffer (e.g., Row 0xB remains in the row buffer  2).
+To transmit a 1 ( 3), the sender activates a different row to cause interference in the corresponding DRAM bank (i.e., a row buffer conflict between Row 0xD and Row 0xE  4).
+At the end of the communication turn (after transmitting M bits), the sender first executes a memory fence (Line 14) to ensure that all the PEI instructions have been executed and then signals the receiver to start probing the DRAM banks (Line 15).
+Step 3.
+The receiver starts probing the DRAM banks by issuing PEI instructions to activate the same row as the one activated during initialization (Line 25).
+Then, the receiver measures the time it takes to complete the request (Line 27).
+If the latency exceeds a predetermined hit threshold, the receiver senses the interference in the row buffer and decodes the bit as a 1 (Lines 28-29  5, 6).
+Otherwise, the access is considered a row buffer hit and the receiver decodes the bit as a 0 (Lines 30-31  7, 8).
+The sender and the receiver repeat the same steps for each batch until the entire message is transmitted.
 Sender-Receiver Synchronization.
-The sender and the receiver synchronize in order to ensure that the sender’s and receiver’s actions (i) are accurately aligned so that they do not interfere with each other and fail to communicate the message and (ii) to overlap the sender and receiver operations to increase the throughput of the attack.
-To achieve this, the sender and the receiver communicate using semaphores.
-The value of the semaphore indicates the number of batches that the sender has completed transmitting but the receiver has not yet completed probing.
-When the sender completes the transmission of a batch of bits, they increment the semaphore to signal the receiver to start probing the DRAM banks.
-This way, the sender overlaps the transmission of multiple batches of bits with the receiver’s probing operations without the need to wait for the receiver to complete the probing of the DRAM banks.
-Doing so, the sender can freely proceed to future batches without relying on synchronizing with the receiver.
-The receiver blocks on the semaphore until the sender increments it.
-When the semaphore is incremented, the receiver decrements it and starts probing the DRAM banks.
-When the receiver completes the probing of the DRAM banks, they read the semaphore again and the process repeats until the entire message is transmitted.
-In our attack implementation (see §5.1) we use POSIX semaphores to implement the sender-receiver synchronization and use a batch size of 4 bits.
-4.1.2 PuM-based Covert Channel
+The sender and the receiver synchronize to 1) ensure their actions do not interfere with each other and 2) overlap the latencies of their operations to increase the throughput of the attack.
+To do so, the sender and the receiver use a semaphore.
+The value of the semaphore indicates the number of batches that the sender has completed transmitting, but the receiver has not yet probed.
+The receiver blocks on the semaphore until the sender increments it (i.e., does not start probing).
+When the sender completes the transmission of a batch, they increment the semaphore to signal the receiver to start.
+Upon receiving the signal, the receiver decrements the semaphore and starts probing the DRAM banks.
+In parallel, the sender can start transmitting the next batch that utilizes a different set of DRAM banks.
+This way, they can overlap the transmission latency and the probing latency.
+When the receiver completes probing, they read the semaphore again, and the process repeats until the entire message is transmitted.
+In our attack implementation (§5.1), we use POSIX semaphores to implement the sender-receiver synchronization and use a batch size of 4 bits.
+4.2 PuM-based IMPACT Covert-Channel Attack
 Baseline PuM Architecture.
@@
 However, in this attack, we assume a PuM-based RowClone operation.
-For our attack, we focus on the Fast Parallel Mode in RowClone, which copies the data of a row to another row within a DRAM subarray.
-Figure 9 demonstrates how the content of the source row, i.e., src, is copied to the destination row, i.e., dst using RowClone’s Fast Parallel Mode.
-Initially, the rows and the row buffer are closed ( 1).
-RowClone operation consists of two key steps.
-First, the src is activated to store the entire row’s data in the row buffer ( 2).
-Second, the dst is activated, which leads to copying the contents of the row buffer in the dst ( 3).
-RowClone significantly reduces the latency and energy consumption of the bulk data copy operation from one row to another as it enables data movement without transferring data between the DRAM and the processing units.
+RowClone’s Fast Parallel Mode works by issuing back-to-back two activation commands: the first activation copies the content of a source row (i.e., src) into the row buffer.
+The second command connects the destination row (i.e., dst) to the bitlines.
+Since the row buffer already stored the src’s data, by the time the dst is activated, the data in dst is overwritten by src’s data.
 To execute RowClone, we assume that the user space application specifies a source and a destination virtual address range and the corresponding mask to copy only specific parts of the source to the destination.
 This way, the memory controller receives one RowClone request and breaks it into parallel RowClone requests, one for each set bit of the mask.
-Figure 9: Overview of RowClone’s Fast Parallel Mode.
 Attack Overview.
@@
 Source and address ranges correspond to physical pages spanning across all DRAM banks and by using the mask, the sender can encode the transmitted message by causing contention only in selected DRAM banks.
-The receiver repeatedly issues RowClone operations and measures the time it takes to complete these operations.
-This way, it can detect contention in a DRAM bank when another process accesses a different row in the same DRAM bank.
-If the sender issues a RowClone operation to a DRAM bank, it induces contention and the receiver detects a higher latency for its RowClone operation.
-This latency indicates that the corresponding bit is a logic-1.
-Otherwise, if the sender does not induce contention to a DRAM bank, the receiver detects a lower latency for the RowClone operation, indicating that the corresponding bit is a logic-0.
-Figure 10 and Listing 2 show the end-to-end flow of IMPACT-PuM and all the steps involved to establish the covert channel between the sender and the receiver.
+The receiver repeatedly issues RowClone operations and measures the latency of these operations.
+This way, if the sender issues a RowClone operation to a DRAM bank, it induces contention, and the receiver detects a higher latency for its RowClone operation, indicating the transmitted bit is logic-1.
+Otherwise, the receiver detects a lower latency for the RowClone operation, indicating that the corresponding bit is a logic-0.
+Fig. 5 and Listing 2 show the end-to-end flow of IMPACT-PuM and all the steps involved to establish the covert channel between the sender and the receiver.
+Figure 5: PuM covert channel attack flow.
 Step 1.
-First, the receiver initializes all DRAM banks by issuing a RowClone operation to copy the contents of a source row to a destination row (Line 16).
-Then, the sender and the receiver synchronize on a barrier (Lines 10&17) which ensures that they start the communication turn at the same time.
+First, the receiver initializes all DRAM banks by issuing a RowClone operation (Line 16).
+Then, the sender and the receiver synchronize on a barrier (Lines 10 and 17) which ensures that they start the communication turn at the same time.
 Step 2.
-The sender uses a single RowClone operation to transmit in parallel an N-bit message (where N is the number of banks) to the receiver and encodes the message in the mask operand of the RowClone operation (Lines 11-12).
-The sender encodes the message by setting the corresponding bits in the mask operand to 1 or 0 ( 1, 2).
-Doing so, the sender causes contention in the corresponding DRAM banks and activates a different row in the row buffer.
+The sender uses a single RowClone operation to transmit an N-bit message (where N is the number of banks) to the receiver in parallel and uses the mask of the RowClone operation to encode the message (Lines 11-12).
+The sender encodes the message by setting the corresponding bits in the mask to 1 or 0 ( 1, 2).
 To send a 1, the sender triggers RowClone in the corresponding DRAM bank and induces contention by altering the activated row in the row buffer ( 3).
 To send a 0, the sender does not trigger RowClone in the corresponding DRAM bank and does not induce contention ( 4).
-The sender synchronizes with the receiver using a barrier operation (Lines 13&18).
-Figure 10: End-to-End Flow of the Processing-using-Memory Covert Channel Attack.
+The sender synchronizes with the receiver using a barrier operation (Lines 13 and 18).
 1 Memory Controller:
@@
 The receiver continuously issues one RowClone operation for each DRAM bank (in contrast to the sender who issues only a single RowClone operation), to the initial virtual address range (i.e., rows initialized in Step 1) and measures the latency of each operation ( 2, Line 24).
-If the sender induces contention to a DRAM bank, the receiver high latency for the RowClone operation which indicates that the corresponding bit in the message is set to 1 ( 3, Lines 25-26).
+If the sender induces contention in a DRAM bank, the receiver observes high latency for the RowClone operation which indicates that the corresponding bit in the message is set to 1 ( 3, Lines 25-26).
 Otherwise, the receiver detects a fast RowClone operation which indicates that the bit is set to 0 ( 4, Lines 27-28).
-Advantages over PnM-based Covert Channel.
-IMPACT-PuM improves upon the PnM-based covert-channel attack in a major way.
-Using a single RowClone operation, the sender can transmit in parallel an N-bit message to the receiver, where N is the number of DRAM banks which requires less computational resources to transmit the same amount of information compared to the PnM-based covert-channel attack.
-For example, using a single thread, the sender can activate all N DRAM banks and transmit N bits of information in parallel while in the PnM-based covert-channel attack, the sender needs to separately execute a PEI instruction to activate each DRAM bank.
-4.2 PnM-based Side-Channel Attack on DNA Read Mapping
-DNA Read Mapping.
-DNA is one of the most unique identifiers of an individual, such that even a small fraction of DNA can include sufficient information for linking to the owner’s identity[108].
-Therefore, preserving DNA privacy carries importance.
-DNA sequencing of a biological sample produces short DNA fragments called reads.
-Read mapping is a fundamental operation to find the matching locations of reads in the reference genome that acts as a template genome to construct the complete genome of the sample.
-Figure 11 shows a high-level overview of a DNA read mapping (DNA-RM) workflow.
-DNA-RM has two major key steps: (i) seeding and (ii) alignment.
-During the seeding step, we identify the possible locations of the reads on the reference genome based on sequence similarities between the reads and the reference genome.
+Advantage over IMPACT-PnM.
+In IMPACT-PuM, the sender can use a single RowClone operation to transmit an N-bit message to the receiver in parallel, where N is the number of DRAM banks.
+Thus, IMPACT-PuM requires less computational resources to transmit the same amount of information compared to IMPACT-PnM.
+For example, using a single thread, the sender can activate all N DRAM banks and transmit N bits of information in parallel.
+In contrast, in IMPACT-PnM, the sender needs to execute one PEI per bit.
+4.3 PnM-based Side-Channel Attack on Read Mapping
+We build a PnM-based side-channel attack that exploits PiM operations to leak private information of a concurrently running application.
+To demonstrate the IMPACT side-channel attack, we target a DNA sequence analysis application.
+Read Mapping.
+DNA is one of the unique identifiers of an individual [73], such that even a small fraction of DNA can include sufficient information for linking to the owner’s identity [70, 71].
+Therefore, preserving privacy while processing genomic data carries utmost importance to protect sensitive information.
+Sequencing the DNA of a biological sample produces short DNA fragments called reads.
+The reads are then processed by read mapping, which is a fundamental operation to find the matching locations of reads in the reference genome that acts as a template to construct the complete genome of the sample.
+Read mapping (RM) has two key steps: (i) seeding and (ii) alignment.
+During the seeding step, RM identifies the possible locations of the reads on the reference genome based on sequence similarities between the reads and the reference genome.
 Seeding involves hashing small segments of the DNA sequences (i.e., seeds) and probing a hash table that is built from the reference genome.
-After seeding, we align the reads with their candidate regions on the reference (i.e., potential reference segments) using dynamic programming-based methods to find each read’s best matching location among all possible locations in the alignment step.
-Figure 11: A simple overview of DNA read mapping workflow: (a) seeding and (b) alignment
-DNA-RM requires a large amount of data movement to access different parts of the hash table and then corresponding regions on the reference genome.
-Hence, several prior works aim to improve the performance of DNA-RM by offloading seeding and alignment to PiM-enabled systems to reduce the data movement overheads [33, 109, 110, 111, 112, 113, 114, 115].
-In this work, we demonstrate that IMPACT can be used to launch a high-throughput side-channel attack against DNA-RM and quickly leak a large amount of information about the sample genome.
+After seeding, RM aligns the reads with their candidate regions on the reference (i.e., potential reference segments) using dynamic programming-based methods to find each read’s best matching position among all possible locations in the alignment step.
+RM requires a large amount of data movement to access different parts of the hash table and corresponding regions on the reference genome.
+Hence, several prior works aim to improve the performance of RM by offloading seeding and alignment to PiM-enabled systems to reduce the data movement overheads [33, 96, 97, 98, 99, 100, 101, 102].
+In this work, we demonstrate that an attacker can launch a high-throughput side-channel attack against RM by utilizing PiM instructions and quickly leak a large amount of information about the sample genome.
 Attack Overview.
-We assume that the victim is running a DNA-RM application on a PiM-enabled system and the attacker is able to run a malicious application on the same system.
-The attacker and the victim are using the same read mapping tool (e.g., minimap2 [116]), and the hash table is shared between the attacker and the victim.
-[^3] The hash table that contains information about the location of seeds in the reference genome is created once (acts as a database), and each user (attacker and victim in our case) of minimap2 probes it with queries.
-The attacker’s goal is to leak information about the test genome by observing the DNA-RM application’s memory access patterns.
-Figure 12 shows the high-level overview of the attack.
-The victim application hashes each seed of the test genome and offloads the corresponding seeding and alignment step to the PiM-enabled system  1.
+We assume that the victim is running a RM application on a PiM-enabled system and the attacker is able to run a malicious application on the same system.
+The attacker and the victim are using the same read mapping tool (e.g., minimap2 [103]).
+The read mapping tool constructs a hash table that contains information about the seed locations in the reference genome, and allows each user in the system to probe the hash table via queries during read mapping.
+We assume the hash table is distributed across multiple DRAM banks.
+This assumption is realistic as many modern DRAM address mapping schemes (e.g., [104, 105, 106, 107]) aim to interleave consequent memory chunks across different DRAM banks to exploit bank-level parallelism [108, 109].
+Fig.
+6 shows the high-level overview of the attack.
+The victim application extracts the seeds from the reads and offloads the corresponding seeding and alignment step to the PiM-enabled system  1.
 During the seeding step, the corresponding DRAM row gets activated (i.e., Index 0-5 of the hash table)  2.
 The attacker continuously uses PnM instructions to probe the shared hash table in an attempt to identify if the victim application is accessing a specific index of the hash table  3.
-If the attacker observes that the victim application is accessing a specific index, it retrieves partial or exact information about the potential locations of the test genome in the reference genome.
-As shown in previous works on DNA imputation [117, 118], the attacker can use the leaked information to infer properties about some regions of the private sample genome  4.
-Figure 12: High-level overview of IMPACT attack model on PiM-enabled DNA-RM
+If the attacker observes that the victim application is accessing a specific index, it retrieves partial or exact information about the potential locations of the reads in the reference genome.
+As shown in previous works on DNA imputation [110, 111, 112, 113], the attacker can use the leaked information in a completion attack [71] to infer properties about some regions of the private sample genome  4.
+Figure 6: High-level overview of IMPACT attack model on PiM-enabled RM.
 5 Evaluation
@@
 Covert Channel Threat Model.
-Our threat model for the covert-channel attack envisions a scenario where a sender and a receiver operate on the same PiM-enabled system to exchange information.
-PiM instructions are available to user space programs without any restrictions and are not privileged instructions.
-The receiver is assumed to have access to cpuid and rdtscp instructions, enabling high-precision measurement of memory access times.
-For IMPACT-PnM, the sender and the receiver need to colocate data in the same DRAM banks.
-For IMPACT-PuM, (i) the sender allocates two virtual address ranges whose corresponding physical pages span across all DRAM banks, and (ii) RowClone guarantees atomicity - when the sender issues a RowClone operations that span across multiple banks, all individual bank-level RowClone operations need to be completed before another process accesses DRAM.
+We assume a scenario where a sender and a receiver operate on the same PiM-enabled system to exchange information.
+PiM instructions are available to user space applications without any restrictions.
+The receiver has access to cpuid and rdtscp instructions, enabling high-precision measurement of memory access latencies.
+For IMPACT-PnM, the sender and the receiver co-locate data in the same DRAM banks.
+For IMPACT-PuM, (i) the sender allocates two virtual address ranges whose corresponding physical pages span across all DRAM banks, and (ii) RowClone guarantees atomicity: when the sender issues a RowClone operation that spans across multiple banks, all individual bank-level RowClone operations need to be completed before another DRAM operation is executed.
 Side Channel Threat Model.
-Our threat model for the side-channel attack envisions a scenario where an attacker process operates on the same PiM-enabled system as a victim process and leaks information about the victim process.
-We use a DNA-RM implementation based on minimap2 [116] and we modify the implementation to offload the seeding and alignment step to the PiM-enabled system.
+We assume a scenario where an attacker process operates on the same PiM-enabled system as a victim process and leaks information about the victim process.
+We use an RM implementation based on minimap2 [103], and we modify the implementation to offload the seeding and alignment step to the PiM-enabled system.
 We assume that the alignment step includes chaining.
 We use the human genome as the reference and compare it against synthetic sample genomes.
-We assume that the hash table is distributed across each bank (i.e., DRAM employs a page interleaving mapping).
-We experiment with multiple different seed sizes, and we report the best-performing selection.
+We experiment with multiple seed sizes and report the best-performing selection.
 Proof-of-Concept.
 We construct a proof-of-concept of all IMPACT variants using the Sniper Multicore Simulator [114].
-We choose to construct the PoC using Sniper for two reasons.
-First, it is validated against modern Intel processors [120].
-Second, there is no commercially available system that follows a similar architecture as PEI or any modern system that enables end-to-end RowClone operations.
-We extend Sniper to (i) accurately model all internal DRAM operations, (ii) emulate and simulate the PiM interface and PiM operations, and (iii) emulate the functionality of the cpuid and the rdtscp instructions in order to measure execution cycles from the user space.
-Table 2 shows the configuration our simulation setup.
-Before launching the attack, the sender and the receiver threads warm up by executing both attacks in order to avoid compulsory DRAM accesses for instructions or page table walks (i.e., after an L2 TLB miss, the page table walker might issue requests to DRAM) and reduce the noise of the attack.
-For IMPACT-PnM, we assume that both the attacker and the receiver use PEIs to perform an addition operation that takes $\sim 3$ cycles to execute (in addition to the actual DRAM access).
-Comparison Points for Covert-Channel Attacks. To showcase the effectiveness of IMPACT, we evaluate the leakage throughput of five different covert-channel attacks.
-(i) DRAMA-clflush: The main memory covert-channel attack proposed in [70] which uses clflush instructions to send memory requests directly to main memory.
+[^3] There is no commercially available system that follows a similar architecture as PEI or any modern system that enables end-to-end RowClone operations.
+We extend Sniper to (i) accurately model all internal DRAM operations, (ii) emulate and simulate the PiM interface and PiM operations, and (iii) emulate the functionality of the cpuid and the rdtscp instructions to measure execution cycles from the user space.
+Table 2 shows the configuration of our simulation setup.
+Before launching the attack, the sender and the receiver threads warm up by executing both attacks to avoid compulsory DRAM accesses for instructions or page table walks.
+For IMPACT-PnM, we assume that both the attacker and the receiver use PEIs to perform an operation that takes $\sim 3$ cycles to execute (in addition to the actual DRAM access).
+Comparison Points for Covert-Channel Attacks. To showcase the effectiveness of IMPACT, we evaluate the leakage throughput of seven different covert-channel attacks.
+(i) DRAMA-clflush: The row buffer-based covert-channel attack proposed in [68] which uses clflush instructions to send memory requests directly to main memory.
 clflush only probes the LLC to flush the cache line.
-(ii) DRAMA-Eviction: The main memory covert channel attack proposed in [70] which uses cache eviction sets to send memory requests directly to main memory.
-In both DRAMA attacks, the sender creates contention in the DRAM banks by causing row buffer conflicts and the receiver detects the contention to decode the message.
-(iii) Streamline: A state-of-the-art cache covert channel proposed in [121] which enables a flush-less collusion between the sender and the receiver.
-(iv) DMA Engine: A main memory attack that uses the DMA engine to bypass the caches and send memory requests directly to main memory.
-In this scenario, we assume a powerful attacker can infer the DRAM access time directly from the user space by disregarding context switch and OS-related latencies.
-(v) PnM-OffChip: A PnM-based attack that probes the locality monitor and a state-of-the-art off-chip predictor [122] before offloading PEI instructions to compute units near DRAM banks.
-If the off-chip predictor predicts that data experience high locality, the computation is performed by the CPU instead of the PnM, leading to reduced throughput.
-(iv) IMPACT-PnM: Our proposed attack that bypasses the locality monitor and offloads PEI instructions directly to compute units near DRAM banks in order to establish a covert channel between the sender and the receiver.
-(v) IMPACT-PuM: Our proposed attack that makes use of RowClone to offload bulk-copy operations directly to main memory and establish a covert channel between the sender and the receiver.
+(ii) DRAMA-Eviction: The row buffer-based covert channel attack proposed in [68] which uses cache eviction sets to send memory requests directly to main memory.
+(iii) Streamline: State-of-the-art cache covert channel proposed in [115] which enables a flushless communication.
+(iv) DMA Engine: A row buffer-based attack that uses the DMA engine to bypass the caches.
+We assume a powerful attacker can infer the DRAM access time directly from the user space by ignoring context switch and OS-related latencies.
+(v) PnM-OffChip: A PnM-based attack in a PnM architecture with locality monitor and a state-of-the-art off-chip predictor [116].
+If the data is predicted to have high locality, the computation is performed by the CPU instead of the near-memory compute units, leading to reduced attack throughput.
+(vi) IMPACT-PnM (§4.1): An attack that bypasses the locality monitor and offloads PEI instructions directly to compute units near DRAM banks in order to establish a covert channel.
+(vii) IMPACT-PuM (§4.2): An attack that leverages RowClone to offload bulk-copy operations directly to main memory and establish a covert channel.
 For a fair comparison against DRAMA and Streamline, we showcase the upper bound of the communication throughput achieved by each attack.
-To evaluate the throughput of the DRAMA and Streamline-based attacks, we use our simulation infrastructure to extract parameters such as the LLC hit latency, average LLC miss latency, cache lookup latency, cache hit/miss ratio, and feed them in an analytical model that calculates the maximum throughput of the attack.
-We validated the analytical model against the throughput achieved by DRAMA and Streamline in a real system.
-For example, in Streamline, the authors report a maximum communication throughput of 1.8M Mb/s and our model estimates that the upper bound of the throughput is 2.7 Mb/s for the smallest LLC size.
+To calculate their throughput, we use our simulation infrastructure to extract parameters such as the LLC hit latency, average LLC miss latency, cache lookup latency, cache hit/miss ratio, and feed them in an analytical model that calculates the maximum throughput of the attack.
+We validate the model against the real system throughput of DRAMA and Streamline (e.g., Streamline [115] reports a maximum communication throughput of 1.8 Mb/s and we model the upper bound as 2.7 Mb/s for the smallest LLC).
 Noise Sources in Simulation Setup and Error Rate.
-There are two main microarchitectural structures that generate noise in the simulated system: 1) hardware prefetchers and 2) page table walkers.
-To observe the impact of application-induced noise on the throughput of IMPACT attacks, we run applications from GraphBIG [123] benchmark suites concurrently with attacker and victim (or sender and receiver) processes.
-We measure the throughput of each attack by incorporating the error rate in the measurement: we measure the throughput only based on successful leakage.
+We simulate hardware prefetchers and page table walkers to induce noise in the simulated system.
+We measure the throughput of each attack only based on the successfully leaked data (i.e., incorporating the error rate in our measurements).
 Table 2: Simulation Configurations
 CPU 4-core, 4-way, OoO x86 core, 2.6GHz
-MMU L1 I-TLB: 128-entry, L1 DTLB (4KB): 64-entry, 4-way assoc, 1-cycle access, L1 DTLB (2MB): 32-entry, 4-way assoc, 1-cycle access, L2 TLB: 1536-entry, 12-way assoc, 12-cycle access, 3 Split PWCs: 32-entry, 2-cycle access
+MMU L1 DTLB (4KB): 64-entry, 4-way assoc, 1-cycle access, L1 DTLB (2MB): 32-entry, 4-way assoc, 1-cycle access, L2 TLB: 1536-entry, 12-way assoc, 12-cycle access
 L1 Cache I-cache: 32 KB; L1D Cache: 32 KB, 8-way assoc, 4-cycle access latency, LRU, IP-stride prefetcher [117]
@@
 We first demonstrate the PoC results and how we successfully mount the attack in the simulated system.
-Figures 13 and 14 show the latency (in cycles) measured by the receiver after executing a PiM instruction, one for each of the 16 banks (we assume that the message is 16-bits large), in IMPACT-PnM and IMPACT-PuM, respectively.
-As we observe, the receiver is able to successfully distinguish whether the sender causes contention in the bank in both IMPACT-PnM and IMPACT-PuM.
-As shown in the figure, given a latency threshold of 150 cycles, the receiver can decode the complete message: if the latency is less than 150 cycles (i.e., row buffer hit), the receiver decodes the message as a logic-0, otherwise (i.e., row buffer conflict) it decodes the message as a logic-1.
-Figure 13: Measured latency (in cycles) by the receiver to execute a PEI instruction for each one of the 16 banks in order to decode a 16-bit message.
-Figure 14: Measured latency (in cycles) by the receiver to execute a RowClone instruction for each one of the 16 banks in order to decode a 16-bit message.
-5.3 Throughput Analysis of IMPACT Covert-Channel Attacks
-We examine the throughput of all different attacks to understand the potential of IMPACT in establishing a high throughput covert channel between the sender and the receiver.
-Figure 15 shows the throughput achieved by all different attacks when one sender and one receiver thread run concurrently across systems with increasing LLC sizes.
-We make four key observations.
+Fig.
+7 shows the latency (in cycles) measured by the receiver after executing (a) a PEI instruction and (b) a RowClone operation, one for each of the 16 banks assuming the message is 16-bits large, in IMPACT-PnM and IMPACT-PuM, respectively.
+We observe that the receiver can successfully distinguish whether the sender causes contention in the bank in both IMPACT-PnM and IMPACT-PuM.
+The receiver can decode the complete message using a threshold value of 150 cycles as follows.
+If the latency is less than 150 cycles (i.e., row buffer hit), the receiver decodes a bit as logic-0, otherwise (i.e., row buffer conflict) it decodes the bit as logic-1.
+Figure 7: Latency (in cycles) measured by the receiver to execute (a) a PEI instruction and (b) a RowClone instruction for each bank in order to decode a 16-bit message.
+5.3 Covert-Channel Attack Throughput Analysis
+Fig.
+8 shows the throughput achieved by different covert-channel attacks when one sender and one receiver thread run concurrently across systems with increasing LLC sizes.
+We make five key observations.
 First, IMPACT-PnM and IMPACT-PuM achieve significantly higher throughput than all other attack vectors.
-For example, IMPACT-PnM achieves a throughput of 12.87 Mb/s irrespective of the cache size, which is 2.29$\times$ (up to 4.91$\times$ with an LLC size of 128MB) higher than the throughput achieved by the second-best performing attack, DRAMA-clflush.
-Second, DRAMA and Streamline achieve decreasingly lower throughput as the LLC size increases.
+IMPACT-PnM and IMPACT-PuM achieve communication throughputs of 12.87 Mb/s and 14.16 Mb/s, respectively, irrespective of the cache size.
+They support up to 4.91$\times$ and 5.41$\times$ of the throughput achieved by the state-of-the-art main memory-based attack, DRAMA-clflush.
+Second, IMPACT-PuM achieves 10% higher throughput than that of IMPACT-PnM since the sender can issue all RowClone requests in parallel, whereas in IMPACT-PnM, the sender has to send the PiM instructions in a sequential manner.
+Third, as the LLC size increases, the throughput of DRAMA-eviction, DRAMA-clflush, and Streamline decreases significantly.
 This is because the larger the LLC size, the higher the latency it takes to access the LLC.
-Third, IMPACT-PuM achieves 10% higher throughput than IMPACT-PnM since the sender can issue all RowClone requests in parallel, whereas in IMPACT-PnM the sender has to send the PiM instructions in a sequential manner.
-Fourth, the DMA engine attack achieves a throughput of 5.27 Mb/s irrespective of the cache size, but is almost 2.4x slower than the PnM-based one due to the OS-related overheads associated with it (e.g., context switch overhead, OS instructions, etc.).
-Fifth, the PnM-OffChip attack leads to high throughput, up to 12.64 Mb/s, but the throughput gets reduced down to 10.64 Mb/s as the cache size increases since the off-chip predictor consults more in favor of caching data when the LLC is large, essentially reducing the attack throughput.
-We conclude that IMPACT-based attacks can achieve significantly higher throughput than the existing state-of-the-art cache covert channels.
-Figure 15: Communication throughput comparison across increasing LLC sizes.
+Fourth, the DMA engine attack achieves a throughput of 5.27 Mb/s irrespective of the cache size but is almost 2.4$\times$ slower than IMPACT-PnM due to its OS-related overheads (e.g., context switch overhead, OS instructions).
+Fifth, the PnM-OffChip attack achieves a maximum throughput of 12.64 Mb/s.
+As the LLC size increases, the throughput is reduced to 10.64 Mb/s because the off-chip predictor decides to cache more data when the LLC is large.
+We conclude that IMPACT achieves significantly higher throughput than the existing state-of-the-art main memory-, cache- and DMA engine-based covert-channel attacks.
+Figure 8: Throughput across increasing LLC sizes.
 Breakdown of Throughput Improvements.
-The impact of the fundamental difference between PiM architectures and traditional architectures can be seen as the difference between PnM-OffChip and DRAMA-clflush attack throughputs in Fig.
-15.
-The impact of bypassing the locality monitor can be seen as the difference between IMPACT-PnM and PnM-OffChip attack throughputs.
-Finally, the impact of the parallelism provided by RowClone can be seen as the difference between IMPACT-PuM and IMPACT-PnM attack throughputs.
-To better understand why IMPACT-PuM achieves higher throughput than IMPACT-PnM, we evaluate the time it takes for the sender and the receiver to perform their respective operations in both attacks.
-Figure 16 shows the time (in cycles) it takes for the sender to send the message and for the receiver to read the message in IMPACT-PnM and IMPACT-PuM.
-As we observe, in IMPACT-PuM, the sender can send the message 14$\times$ faster than the sender in IMPACT-PuM since it can issue all RowClone requests in parallel.
-In IMPACT-PnM, even though the sender is significantly slower than the sender of IMPACT-PuM, the sender and the receiver overlap, at the cost of synchronization, in order to hide the latency of the sender.
+Fig.
+9 shows the time (in CPU cycles) it takes for the sender to send the message and for the receiver to read the message in IMPACT-PnM and IMPACT-PuM.
+We observe that in IMPACT-PuM, the sender can send the message 14$\times$ faster than the sender in IMPACT-PuM since it can issue all RowClone requests in parallel.
+In IMPACT-PnM, even though the sender is significantly slower than the sender of IMPACT-PuM, the sender and the receiver overlap, at the cost of synchronization, to hide the latency of the sender.
 This is why IMPACT-PnM achieves only 10% lower throughput than IMPACT-PuM.
 Figure 9: Time (in cycles) it takes (i) for the sender to send the message and (ii) for the receiver to read the message in IMPACT-PnM and IMPACT-PuM.
-5.4 Throughput Analysis of IMPACT Side-Channel Attack
-Figure 17 shows the throughput (taking into account only correct identification of activated hash table entries), left y-axis, and the error rate (i.e., wrong identification of activated hash table entries), right y-axis, of the side-channel attack against DNA-RM across different number of DRAM banks.
-We make two key observations with respect to the throughput of the attack.
-First, when using a PiM-enabled DRAM device with 1024 banks, the attacker is able to leak information about the sample genome at a throughput of 7.57 Mb/s.
-Second, as the number of banks increases, even though the throughput of the attack decreases down to 2.56 Mb/s (since the attacker has to probe more banks and the error rate also increases), the leaked information provides more exact information about the sample genome since the amount of hash table data per bank decreases, reducing the number of potential locations of the reads in the reference genome.
-We make two key observations with respect to the error rate of the attack.
-First, for the 1024 bank configuration, the error rate of the attack is less than 5%.
-Second, we observe that the error rate of the attack is less than 15% across all different DRAM bank configurations.
-As the number of banks increases, the error rate of the attack increases since the attacker has to probe more banks, which makes the attack more prone to noise and reduces its bandwidth(i.e., DRAM command delay, memory controller scheduling, etc.).
-Even though the error rate of the attack increases as the number of banks increases, the attacker is able to leak information about the test genome with higher accuracy.
-For example, when using a PiM-enabled DRAM device with 1024 banks, the victim activates a row with 16 hash table entries, and the attacker is able to identify if the sample read is present in any of the locations pointed by the 20 hash table entries with an accuracy of 96%.
-When using a PiM-enabled DRAM device with 2048 banks, the victim activates a row with 8 hash table entries, and the attacker is able to identify if the sample read is present in any of the locations pointed by the 8 hash table entries with an accuracy of 91%.
+5.4 Side-Channel Attack Throughput Analysis
+Fig.
+10 shows the throughput and the error rate of the side-channel attack against read mapping (RM) across different numbers of DRAM banks in the left and right y-axis, respectively.
+We measure the throughput of the attack based on the correct guesses and calculate the error rate based on the number of incorrect guesses.
 The exact end-to-end accuracy of identifying the properties of the sample genome depends on the chaining mechanism and the imputation algorithm [110, 111] used by the attacker, and it is out of the scope of this work.
-Figure 17: Leakage throughput and error rate of the side-channel attack against DNA-RM across different numbers of DRAM banks.
-6 Mitigations for IMPACT
-This section discusses the effectiveness of three different techniques to mitigate IMPACT: (i) An address mapping scheme that partitions memory banks across different processes, (ii) a memory controller policy that enforces closing the row buffer after every memory access, and (iii) a memory controller policy that enforces constant-time DRAM accesses to hide the timing differences from the attackers.
+We make two key observations.
+First, with a PiM-enabled DRAM with 1024 banks, the attacker can leak information about the sample genome at a throughput of 7.57 Mb/s and an error rate of $<5$%.
+Second, as the number of banks increases to 8196, the throughput of the attack decreases to 2.56 Mb/s and the error rate increases to $<15$%.
+This is because the attacker has to probe more banks, which makes the attack more prone to noise and reduces its bandwidth.
+However, at the same time, the leaked information provides more exact information about the sample genome since the number of hash table entries per bank decreases, reducing the number of potential locations of the reads in the reference genome per bank.
+For example, in a PiM-enabled DRAM device with 1024 banks, the victim activates a row with 16 hash table entries.
+Assuming the attacker identifies the activated bank correctly, they determine that one of the 16 hash table entries was targeted by the victim.
+In contrast, in a PiM-enabled device with 2048 banks, the victim activates a row with 8 hash table entries.
+When the attacker identifies the activated bank correctly, they determine that one of the 8 hash table entries was targeted by the victim, which leaks more precise information on the sample genome.
+Figure 10: Leakage throughput and error rate of the RM side-channel attack across different numbers of DRAM banks.
+6 Defenses against IMPACT
+This section discusses the effectiveness of three defense mechanisms that eliminates the row buffer-based timing channel.
 Memory Partitioning (MPR).
-Memory bank partitioning can serve as an effective countermeasure against main-memory-based covert channel attacks.
-By dividing the memory into separate banks, each with its own dedicated resources and access control, memory bank partitioning restricts the interactions between different memory regions.
-This partitioning prevents unauthorized communication channels from crossing the boundaries of the designated banks.
-In the context of a IMPACT, memory bank partitioning limits the sender’s ability to transmit information across different memory banks and the receiver’s ability to observe contention in different memory banks.
-By segregating the memory into distinct banks, the covert channel’s ability to exploit timing differences is significantly constrained.
-However, statically partitioning DRAM resources across applications disables data sharing, which can significantly (i) increase data duplication, (ii) increase memory underutilization, and (iii) lead to frequent swapping from the disk when the memory space is full.
+Bank-level memory partitioning enforces that each DRAM bank can only be accessed by one process that it is allocated to.
+In IMPACT, this limits the sender’s ability to transmit information across different DRAM banks and the receiver’s ability to observe contention created by the sender.
+However, this approach has three major drawbacks.
+First, it can lead to limiting the number of concurrently running applications due to the fixed number of DRAM banks available.
+Second, it can cause memory underutilization due to allocating DRAM bank-sized memory to applications.
+Third, it disables data sharing, which can significantly increase data duplication and memory underutilization, and lead to frequent swapping from the disk when the memory space is full.
 Closed Row Policy (CRP).
-Enforcing a memory controller policy that closes a DRAM row after each operation can be an effective measure to mitigate and eventually stop timing covert channel attacks.
-By closing the DRAM row after each operation, the memory controller ensures that subsequent memory accesses, whether legitimate or covert, cannot access the same row without incurring the overhead of row activation.
-In the context of IMPACT, which heavily relies on the difference in memory access latency to transmit information, closing the DRAM row after each operation disrupts the timing patterns exploited by the attacker.
-However, as we show in Section 6, enforcing a closed row policy leads to significant performance overheads, even for workloads with low data locality, as every DRAM access will lead to a DRAM row buffer miss.
+Enforcing a memory controller policy that closes a DRAM row after each operation can be an effective defense against IMPACT.
+By closing the DRAM row after each operation, the memory controller ensures that subsequent memory accesses cannot access the same row without requiring a row activation.
+Closed row policy disrupts the timing patterns exploited by the attacker in IMPACT.
+However, it also leads to significant performance overheads, even for workloads with low data locality, as every DRAM access leads to a row buffer miss.
 Constant-Time DRAM Accesses (CTD).
-One way to mitigate main-memory-based covert channel attacks is to enforce a constant-time DRAM access policy.
+Another defense mechanism is to enforce a constant-time DRAM access (CTD) policy.
 When a memory access is issued, the memory controller waits for a fixed amount of time before returning the data to the processor.
 This fixed amount of time is equal to the worst-case DRAM access latency.
-By enforcing a constant-time DRAM access policy, the memory controller ensures that all memory accesses take the same amount of time, regardless of the accessed memory location.
-In the context of IMPACT, where the communication relies on the timing or side effects of memory accesses, enforcing a constant-time DRAM access policy obfuscates the timing patterns exploited by the attacker.
-However, as we show in Section 6, enforcing a constant-time DRAM access policy leads to significant performance overheads as every DRAM access costs the worst-case DRAM access latency.
+By enforcing CTD, the memory controller ensures that all memory accesses take the same amount of time, regardless of the memory location.
+Thus, obfuscates the timing patterns exploited by the attacker in IMPACT.
+However, this leads to significant performance overheads as every DRAM access has the worst-case latency.
 Performance Evaluation.
-To better understand the effectiveness of the countermeasures, we evaluate the performance of CRP and CTD compared to the baseline open-row policy for 5 multi-programmed graph applications which (i) use and share the input graph (i.e., accesses to the same DRAM bank) and (ii) experience a mix of low and high data locality, i.e., the number of last-level cache misses per kilo instructions.
+We evaluate the performance of CRP and CTD compared to the baseline open-row policy for 5 multi-programmed graph applications which (i) share the input graph (i.e., access to the same DRAM bank) and (ii) experience a mix of low and high data locality (i.e., the number of last-level cache misses per kilo instructions).
 We use a 2-core system where each core runs a different instance of the same application (e.g., core 0 runs an instance of PageRank and core 1 the same instance of PageRank using the same input).
 We evaluate five workloads from GraphBIG benchmark suite [120]: Betweenness Centrality (BC; Data set size: 7.8 GB; LLC MPKI: 0.57), Breadth-First search (BFS; 7.6 GB; 38.59), Connected components (CC; 7.9 GB; 45.2), Triangle counting (TC; 8.2 GB; 5.08), PageRank (PR; 7.0 GB; 1.86).
-Figure 18 shows the performance overhead of CRP and CTD compared to the open-row policy for all 5 workloads.
+Fig.
+11 shows the performance overhead of CRP and CTD compared to the open-row policy.
 We make two key observations.
-First, CTD leads to the highest performance degration, 26% on average, compared to the open row policy.
-Especially, in workloads like PR and BC, CTD leads to a performance degradation of 37% and 45% respectively.
-Second, CRP leads to a performance degradation of 15% on average compared to the open row policy but leads to low performance overheads in workloads that do not benefit from the open row policy like TC, CC and BFS.
-We conclude that existing countermeasures lead to high performance overheads and there is a dire need to come up with techniques that improve security with low performance overhead.
+First, CTD leads to the highest performance degradation, 26% on average, compared to the open row policy.
+Second, CRP leads to an average performance degradation of 15% compared to the open row policy but leads to low performance overheads in workloads that do not benefit from the open row policy such as TC, CC and BFS.
+We conclude that fundamental defense mechanisms that eliminate the timing channel lead to high performance overheads, and there is a dire need for high-performance defense mechanisms.
 Figure 11: Performance overhead of Closed-Row policy and Constant-Time policy compared to Open-Row policy.
 7 Related Work
-To the best of our knowledge, this is the first work that evaluates the implications of PiM on timing attacks.
+To the best of our knowledge, this is the first work that evaluates the timing attack vulnerabilities in PiM systems.
 Security Implications of PiM Systems.
@@
 Arafin et al.
-[127] survey and theoretically analyze security vulnerabilities in PiM systems, with a focus on five specific attack scenarios.
+[121] survey and theoretically analyze security vulnerabilities, focusing on five specific attacks.
 In this work, we quantitatively and qualitatively analyze and evaluate main memory-based covert and side channels in detail.
@@
 [123] extracts neural network architectural information from an RRAM accelerator using power side channels.
-Similarly, Ensan et al.
-[128] proposes an attack revealing implemented intellectual property in RRAM-based PiM architectures using power/timing side channels.
-In contrast, in this work, we demonstrate how to establish high-throughput timing covert channels using PiM operations through software.
-Timing Attacks.
-A significant body of work has focused on exploring timing attacks [130, 131].
-These attacks involve exploiting variations in the execution time of memory and processor core operations to infer sensitive information.
-Researchers have investigated various timing attack techniques, such as cache-based attacks [132, 133, 134, 121, 135, 136, 137, 138, 139] to uncover vulnerabilities in traditional systems.
-There is an extensive body of research related to traditional covert-channel attacks, including Flush+Reload [136], where the sender flushes a shared cache line and monitors the access time by the receiver to infer transmitted data.
-Flush+Flush [135] utilizes cache flush operations to probe for the presence of specific data patterns.
-Prime+Probe [140] exploits cache set conflicts to communicate information between sender and receiver.
-Streamline [121] is a flush-less cache covert channel attack that enables asynchronous communication.
-We compare IMPACT to Streamline in Section 5.3 and show that IMPACT achieves significantly higher throughput.
-Other Main Memory-based Timing Attack Mitigations.
-We already extensively discuss and evaluate mitigations that fundamentally eliminate the timing channel in Section 6.
-We also discuss practical solutions that restrict access to the timing channel through cache management methods and show that they are not applicable to PiM-based attacks.
-Another mitigation method is to restrict the usage of fine-grained timers to prevent attackers from exploiting the timing channel.
-A set of modern processors limits access to fine-grained timers such as Apple M1 cores [141].
-Fine-grained timers are used in a wide range of user space applications, such as performance and power consumption monitoring, multithreading, synchronization and locking, optimizations on protocols such as TCP.
-Restricting access to fine-grained timers can also disable these applications or degrade their effects.
-DAGguise [142] is a defense mechanism against memory timing side-channel attacks.
-DAGguise utilizes a memory access pattern representation obtained through offline profiling and shapes an application’s memory access pattern to eliminate information leakage.
-However, DAGguise does not mitigate row buffer-based timing channels as is and requires the memory controller to employ a closed row policy (evaluated in §6).
+Ensan et al.
+[122] proposes an attack on RRAM-based PiM architectures using power/timing side channels.
+In contrast, we demonstrate how to establish high-throughput timing covert and side channels using PiM operations.
+Other Main Memory-based Timing Attack Defenses.
+We already extensively discuss defenses that fundamentally eliminate the timing channel (§6) and defenses based on cache management methods that are inapplicable.
+Another approach is to restrict the fine-grained timers’ usage to prevent attackers from measure memory requests’ latencies, which is applied in some modern processors (e.g., Apple M1 cores [124]).
+However, these timers are used in many user space applications (e.g., performance and power consumption monitoring, multithreading, and synchronization and locking).
+Restricting access to these timers can disable these applications or degrade their performance.
+DAGguise [125] is a defense mechanism against memory timing side-channel attacks that obfuscates an application’s memory access pattern.
+DAGguise does not mitigate row buffer-based timing channels and requires a closed row policy.
+Trusted execution environments (TEEs) (e.g., Intel SGX [126, 127]) can isolate applications from the rest of the system.
+However, prior works already showcase multiple side-channel attacks on TEEs leveraging the main memory [128], and other structures [129, 130, 131, 128, 132, 133, 134, 135, 136, 137].
 Enhancing System Security using PiM.
-A set of prior works proposes leveraging PIM to enhance the security of modern computing systems.
-SecNDP [143] proposes an encryption and verification mechanism for untrusted PiM devices that enables computation over encrypted data.
-SecNDP only focuses on protecting confidentiality and does not consider or analyze potential timing channels in contrast to IMPACT.
-Nelson et al.[144] propose offloading security critical tasks to near-memory compute units to avoid storing sensitive information in caches and eliminate cache-based side-channel attacks leaking them.
-In contrast, in this work, we demonstrate and evaluate the implications of PiM on main memory-based timing attacks.
-InvisiMem.
-[145] and ObfusMem [146] propose leveraging 3D-stacked memory to design secure processors that provide oblivious RAM equivalent guarantees.
+Several prior works enhance the security by leveraging PiM.
+SecNDP [138] proposes an encryption and verification mechanism for untrusted PiM devices.
+It only focuses on protecting confidentiality and does not consider or analyze potential timing channels.
+Nelson et al.
+[139] propose offloading security critical tasks to near-memory compute units to avoid storing sensitive information in caches and eliminate cache-based side-channel attacks.
+In this work, we demonstrate and evaluate the implications of PiM on main memory-based timing attacks.
+InvisiMem [140] and ObfusMem [141] propose leveraging 3D-stacked memory to design secure processors that provide oblivious RAM equivalent guarantees.
 8 Conclusion
-The adoption of processing-in-memory (PiM) architectures has gained significant attention due to their ability to provide high performance and low energy consumption by reducing data movement bottlenecks.
-In this work, we introduce IMPACT, a new high-throughput main memory timing attack that takes advantage of the direct memory access provided by PiM architectures.
+We introduce IMPACT, a set of high-throughput main memory timing attacks that leverage the direct memory access enabled by PiM architectures.
 IMPACT exploits the shared DRAM row buffer to 1) establish covert channels to communicate with other processes at high throughput and 2) leak security-critical information of victim processes based on their memory accesses.
 IMPACT covert channels leverage parallelism and eliminate cache bypassing steps, resulting in increased throughput compared to existing main memory-based covert channels.
-IMPACT side-channel attack targets a genomics application (i.e., DNA read mapping) and leaks the private characteristics of a user’s sample genome by leveraging PiM operations.
-The proposed countermeasures for IMPACT and the discussion on secure and robust PiM architecture designs highlight the need for addressing security concerns in PiM systems.
+IMPACT side-channel attack targets a genomics application and leaks the private characteristics of a user’s sample genome by leveraging PiM operations.
+We evaluate and discuss three fundamental defense mechanisms against IMPACT.

### v4

@@
-Amplifying Main Memory-Based Timing Covert and Side Channels using Processing-in-Memory Operations
+Revisiting Main Memory-Based Covert and Side Channel Attacks in the Context of Processing-in-Memory
 Abstract
-The adoption of processing-in-memory (PiM) architectures has been gaining momentum because they provide high performance and low energy consumption by alleviating the data movement bottleneck.
-Yet, the security of such architectures has not been thoroughly explored.
-The adoption of PiM solutions provides a new way to directly access main memory, which malicious user applications can potentially exploit.
-We show that this new way to access main memory opens opportunities for high-throughput timing attack vectors that are hard-to-mitigate without significant performance overhead.
-We introduce IMPACT, a set of high-throughput main memory-based timing attacks that leverage characteristics of PiM architectures to establish covert and side channels.
+We introduce IMPACT, a set of high-throughput main memory-based timing attacks that leverage characteristics of processing-in-memory (PiM) architectures to establish covert and side channels.
 IMPACT enables high-throughput communication and private information leakage by exploiting the shared DRAM row buffer.
-To achieve high throughput, IMPACT (i) eliminates expensive cache bypassing steps required by processor-centric main memory and cache-based timing covert and side-channel attacks and (ii) leverages the intrinsic parallelism of PiM operations.
+To achieve high throughput, IMPACT (i) eliminates expensive cache bypassing steps required by processor-centric memory-based timing attacks and (ii) leverages the intrinsic parallelism of PiM operations.
 We showcase two applications of IMPACT.
-First, we build two covert-channel attack variants that run on the host CPU and leverage different PiM approaches (i.e., processing-near-memory and processing-using-memory) to gain direct and fast access to main memory and establish high-throughput communication covert channels.
-Second, we showcase a side-channel attack that leaks private information of concurrently running victim applications that are accelerated with PiM.
-Our results demonstrate that (i) our covert channels achieve 12.87 Mb/s and 14.16 Mb/s communication throughput, respectively, which is up to 4.91$\times$ and 5.41$\times$ faster than the state-of-the-art main memory-based covert channels, and (ii) our side-channel attack allows the attacker to leak secrets with a low error rate.
-To avoid such covert and side channels in emerging PiM systems, we propose and evaluate three defenses that fundamentally eliminate the timing channel.
-11footnotetext: K. Kanellopoulos and F. N. Bostancı are co-primary authors.
+First, we build two covert channels that leverage different PiM approaches (i.e., processing-near-memory and processing-using-memory) to establish high-throughput covert communication channels.
+Our covert channels achieve 8.2 Mb/s and 14.8 Mb/s communication throughput, respectively, which is 3.6$\times$ and 6.5$\times$ higher than the state-of-the-art main memory-based covert channel.
+Second, we showcase a side-channel attack that leaks private information of concurrently-running victim applications with a low error rate.
+Our source-code is openly and freely available at https://github.com/CMU-SAFARI/IMPACT.
+11footnotetext: F. N. Bostancı and K. Kanellopoulos are co-primary authors.
 1 Introduction
-Data movement between the main memory and the processor is a critical bottleneck present in today’s computing systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
-Hence, both industry and academia have invested significant efforts to bring computation closer to where data resides, creating a paradigm called processing-in-memory (PiM)  [13, 23, 8, 24, 25, 26, 11, 27, 19, 28, 1, 2].
-PiM solutions provide significant performance improvement and energy efficiency by reducing data movement.
-PiM solutions take either of two approaches: (i) processing-near-memory (PnM), where the computation happens in a computation unit placed on the same chip as the memory array [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 21, 39, 40, 41, 42], which also appears in real-world products (e.g., UPMEM [43], AxDIMM [41], AiM [44], HB-PNM [45] and FIMDRAM [46]) and (ii) processing-using-memory (PuM) where the memory array performs the computation by leveraging the analog computation capability of DRAM cells [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62].
-Prior works on PnM and PuM clearly demonstrate that these emerging architectures promise great performance and energy efficiency benefits.
-In addition, the existence of real products that implement such techniques [43, 41, 44, 45, 46] indicates that these techniques are likely to be used in many real systems in the near future.
-Therefore, it is important to investigate the security aspect of such techniques and implement countermeasures to avoid potential widespread security vulnerabilities that their deployment can cause.
-Unfortunately, no prior work analyzes and evaluates the security of emerging PiM architectures against timing covert and side-channel attacks.
-In this work, we analyze existing PiM architectures and show that the adoption of PiM solutions creates opportunities for high-throughput main memory-based attacks that are hard-to-mitigate without significant performance overheads.
-This is possible due to two main reasons: (i) PiM architectures provide direct access to main memory, which is essential for their performance and energy efficiency benefits, and (ii) a large set of existing practical defense mechanisms are not applicable in these PiM architectures.
+Data movement between computation units (e.g., CPUs, GPUs, ASICs) and main memory (e.g., DRAM) is a major performance and energy bottleneck in current processor-centric computing systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], and is expected to worsen due to the increasing data intensiveness of modern applications, e.g., machine learning [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] and genomics [34, 35, 36, 37, 38, 39, 40, 41, 42].
+To mitigate the overheads caused by data movement, various works propose Processing-in-Memory (PiM) architectures [13, 43, 8, 44, 45, 46, 11, 47, 19, 48, 1, 2].
+There are two main approaches to PiM [49, 10, 8]: (i) Processing-near-Memory (PnM) [50, 51, 52, 53, 37, 54, 55, 38, 56, 57, 21, 58, 59, 60, 26], where computation logic is added near the memory arrays (e.g., in a DRAM chip, next to each bank, or at the logic layer of a 3D-stacked memory [61, 62, 63, 64, 65, 66, 20, 67, 21, 11, 68, 25, 69, 35, 26]); and (ii) Processing-using-Memory (PuM) [70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85], where computation is performed by exploiting the analog operational properties of the memory circuitry.
+Prior works from industry and academia demonstrate the large performance and energy benefits of PiM architectures.
+A set of PiM techniques are already implemented in real products [86, 60, 87, 88, 89] and many more are expected to be adopted in the near future.
+Therefore, it is important to investigate the security aspects of such techniques and implement countermeasures to avoid potential widespread security vulnerabilities that their deployment can cause.
+Unfortunately, no prior work analyzes and evaluates the security of emerging PiM architectures against timing covert- and side-channel attacks.
+In this work, we analyze PiM architectures and show that the adoption of PiM architectures creates opportunities for critical main memory-based timing attacks due to two reasons.
+First, to eliminate data movement, PiM architectures provide direct access to main memory, which is a key building block for high-throughput main memory-based timing attacks.
+Second, defenses against these attacks either incur high performance overheads or are not applicable to PiM architectures.
 1.
 Direct Access to Main Memory.
-High-throughput main memory-based timing side- and covert-channel attacks need to access main memory with high frequency to observe or change the states of shared main memory structures (e.g., DRAM row buffers).
-In today’s systems, it is difficult to access main memory with high frequency due to the deep cache hierarchies that filter the memory requests of attacker processes.
-[^1] We provide analysis as to why prior attack primitives that bypass caches do not provide fast and reliable direct main memory access in §3.2.
-The adoption of PiM techniques changes how applications and computation units access main memory: (i) due to area and thermal dissipation limitations, compute units in PnM architectures have small or no caches that allow the user space applications to bypass caches more easily (i.e., with fewer memory accesses), and (ii) PuM architectures provide the ability to offload specialized PiM instructions (e.g., bulk data copy and initialization, and bitwise/arithmetic operations) to the main memory that completely bypass the deep cache hierarchies of modern processors.
+Main memory-based timing attacks establish covert communication channels or leak secrets by observing and changing the shared main memory states (e.g., DRAM row buffer states).
+To achieve high throughput, timing attacks require fast and reliable access to main memory.
+In today’s systems, it is difficult to access main memory with high frequency due to the deep cache hierarchies that filter the memory requests and incur additional latency overheads [90].
+[^1] We provide analysis as to why prior attack primitives that bypass caches do not provide fast and reliable enough direct main memory access in §3.2.
+The adoption of PiM techniques changes how applications and computation units access main memory: (i) due to area and thermal dissipation limitations, compute units in PnM architectures have small or no caches that allow the userspace applications to bypass caches more easily (i.e., with fewer instructions and lower latency), and (ii) PuM architectures provide the ability to offload to main memory specialized PiM instructions (e.g., bulk data copy and initialization [75, 91, 73, 92, 85], and bitwise/arithmetic operations [76, 93, 73, 79, 78, 58, 85, 81, 80, 82]) that completely bypass the deep cache hierarchies of modern processors.
 As a result, PiM architectures provide userspace applications with fast and reliable direct main memory access that can be exploited by attackers for high-throughput timing covert- and side-channel attacks.
@@
 Hard-to-Mitigate with Practical Defenses.
-A fundamental solution to memory-based timing attacks is to eliminate the timing channel itself.
-Prior defenses explore fundamental solutions such as memory partitioning and enforcing constant-time memory accesses.
-However, these defenses incur significantly high performance overheads (as shown in Section 6).
-Therefore, several prior works [63, 64, 65, 66] explore practical defenses and detection mechanisms that restrict access to timing channels with cache management-based methods, such as restricting the access to cache management instructions (e.g., clflush instruction in x86 systems [63]) and detecting timing attacks based on cache access patterns (e.g., a high number of cache misses can indicate a timing attack) [64, 65, 66].
-These mechanisms are inapplicable to PiM-based attacks because these attacks completely bypass the cache hierarchy.
-Based on our analysis, we introduce IMPACT, a set of high-throughput In-Memory Processing-based timing Attacks that leverage the direct and fast main memory accesses enabled by PiM architectures.
-Using the direct access via PiM operations, IMPACT observes and exploits the shared DRAM row buffer.
-We demonstrate that IMPACT achieves high communication and information leakage throughput by (i) eliminating expensive cache bypassing steps required by other main memory- and cache-based timing covert- and side-channel attacks and (ii) leveraging the intrinsic parallelism of PiM operations.
+A fundamental solution to memory-based timing attacks is to eliminate the timing channel completely (e.g., by enforcing constant-time memory accesses or introducing intrusive and restrictive memory partitioning between different security contexts).
+However, eliminating the timing channel incurs very high performance overheads (as shown in Section 7).
+Therefore, several prior works [94, 95, 96, 97] propose defenses and detection mechanisms that restrict the use of timing channels with cache management-based methods, such as restricting the userspace applications’ access to cache management instructions (e.g., clflush instruction in x86 systems [94]) and detecting timing attacks based on cache access patterns (e.g., a high number of cache misses can indicate a timing attack) [95, 96, 97].
+These defenses are inapplicable to PiM architectures because these architectures provide direct access to memory by bypassing the cache hierarchy.
+Based on our analysis, we introduce IMPACT, a set of high-throughput In-Memory Processing-based timing Attacks that leverage direct and fast main memory accesses enabled by PiM architectures.
+Using direct access to DRAM via PiM operations, IMPACT observes and exploits the shared DRAM row buffer states.
+We demonstrate that IMPACT achieves high communication and information leakage throughput by (i) eliminating expensive cache bypassing steps required by other main memory- and cache-based timing covert- and side-channel attacks, and (ii) leveraging the intrinsic parallelism of PiM operations.
 We showcase this finding on two covert- and one side-channel attacks.
-To protect future systems against IMPACT covert and side channels, we propose three defense mechanisms that fundamentally eliminate the channel itself.
+To protect future systems against IMPACT, we describe and analyze four defense mechanisms.
 IMPACT-PnM Covert Channel Variant.
-We build a PnM-based covert channel that uses PiM-enabled instructions [67] to directly access the main memory.
-The sender bypasses the locality monitor employed by the system and offloads PiM-enabled instructions for computation directly inside the main memory.
-The receiver measures the time it takes to access a specific set of rows to retrieve the message encoded with row buffer hits and conflicts.
-IMPACT-PnM achieves 12.87 Mb/s communication throughput, which is up to 4.91$\times$ faster than the state-of-the-art main memory-based attack [68].
+We build a covert channel exploiting a PnM architecture called PiM-enabled instructions [67] that enables executing instructions in compute units near DRAM.
+In this variant, the receiver uses PiM-enabled instructions to measure the time it takes to execute a simple arithmetic operation on a set of DRAM rows and detect row buffer conflicts.
+The sender transmits a message by encoding the message with row buffer conflicts in selected addresses and carefully creating conflicts with the receiver’s accesses using PiM-enabled instructions.
+By executing instructions near memory, IMPACT-PnM achieves 8.2 Mb/s communication throughput, which is 3.6$\times$ faster than the state-of-the-art main memory-based covert-channel attack [98].
 IMPACT-PuM Covert Channel Variant.
-We build a PuM-based covert channel that leverages the inherent parallelism in main memory using bulk copy and initialization techniques, specifically RowClone [52], to quickly transmit messages through the shared row buffer states.
-This way, IMPACT-PuM transmits 16-bit large messages 14$\times$ faster than IMPACT-PnM and achieves communication throughput of 14.16 Mb/s (up to 5.41$\times$ faster than the state-of-the-art [68]).
+We build a covert channel exploiting a PuM technique called RowClone [75] that enables bulk data copy and initialization.
+In this variant, the sender transmits a message with in-memory row-copy operations by copying rows in different DRAM banks in parallel for different bits of the message.
+The receiver decodes the message by issuing row-copy operations and measuring their latency.
+By exploiting the parallelism in memory, IMPACT-PuM achieves a communication throughput of 14.8 Mb/s, which is 6.5$\times$ faster than the state-of-the-art [98].
 IMPACT Side-Channel Attack.
-We build a PnM-based IMPACT side-channel attack that exploits PiM operations to leak private information belonging to a concurrently running application based on its memory access characteristics.
-We showcase this side-channel attack on read mapping, one of the fundamental applications in DNA sequence analysis where preserving the privacy of the human genomic data is crucial [69, 70, 71, 72, 73].
-Our side-channel attack leverages PiM operations and leaks the private characteristics of a user’s sample genome by observing the memory access patterns of the read mapping application.
-Our evaluation demonstrates that our side-channel attack leaks the properties of a sample genome at a throughput of 7.57 Mb/s with 96% accuracy.
+We build a PnM-based IMPACT side-channel attack that exploits PiM-enabled instructions [67] to leak private information of a concurrently-running application.
+We showcase this side-channel attack on genomic read mapping, one of the fundamental applications in DNA sequence analysis where preserving the privacy of the human genomic data is crucial [34, 99, 36, 100, 101, 102, 103].
+Our side-channel attack leverages PiM-enabled instructions [67] to leak the private characteristics of a user’s query genome by observing the memory access patterns of the read mapping application.
+Our evaluation demonstrates that our side-channel attack leaks the properties of a query genome at a throughput of 7.6 Mb/s with 96% accuracy.
 Defense Mechanisms.
-Based on our security analysis, we discuss and evaluate potential defense mechanisms that can mitigate IMPACT attacks by eliminating the timing channel.
-We discuss the performance versus security trade-offs of these defenses.
-We demonstrate that the two best-performing defense mechanisms lead to high performance overheads of 26% and 15% on average across five data-intensive workloads.
+Based on our security analysis, we discuss and evaluate potential defense mechanisms to mitigate IMPACT attacks by eliminating the timing channel and reducing the attack throughput.
+We describe the performance versus security trade-offs of these defenses.
+We eventually conclude that mitigating IMPACT incurs high performance overheads and more research is needed to find low-overhead solutions.
 This paper makes the following contributions:
-  • To the best of our knowledge, this is the first work that analyzes and evaluates the security of emerging PiM architectures against timing covert and side-channel attacks.
+  • To our knowledge, this is the first work that analyzes and evaluates the security of emerging PiM architectures against timing covert-channel and side-channel attacks.
   • We introduce IMPACT, a new set of high-throughput main memory-based timing attacks that leverage PiM operations to gain fast and reliable access to main memory.
-  • We showcase two IMPACT covert channel variants, IMPACT-PnM and IMPACT-PuM that exploit PnM and PuM techniques, respectively.
-First, IMPACT-PnM allows the sender and the receiver to exchange messages using PiM-enabled instructions.
-Second, IMPACT-PuM further leverages the parallelism provided by bulk copy PuM techniques to improve the communication throughput.
+  • We demonstrate two IMPACT covert channel variants, IMPACT-PnM and IMPACT-PuM that exploit PnM and PuM techniques, respectively.
+IMPACT-PnM allows the sender and the receiver to exchange messages using PiM-enabled instructions.
+IMPACT-PuM further leverages the parallelism provided by bulk in-memory copy operations like RowClone.
 We show that both attacks outperform the state-of-the-art main-memory-based covert-channel attacks in terms of communication throughput.
-  • We showcase an IMPACT side-channel attack on read mapping by leveraging PnM instructions.
-We show that our side-channel attack enables the attacker to leak sensitive information on the victim’s sample genome with high throughput and a low error rate.
-  • We evaluate and discuss the trade-offs of three defense mechanisms to protect PiM-enabled systems against IMPACT’s covert- and side-channel attacks.
+  • We demonstrate an IMPACT side-channel attack on genomic read mapping by leveraging PnM, using PiM-enabled instructions.
+We show that our side-channel attack leaks sensitive information with high throughput and a low error rate.
+  • We describe and evaluate the trade-offs of four defense mechanisms to protect PiM-enabled systems against IMPACT’s covert- and side-channel attacks.
+We call for future research to find low-overhead solutions.
 2 Background
+This section provides a concise overview of 1) DRAM organization and operation, 2) processing-in-memory, and 3) main memory-based timing attacks.
 2.1 DRAM Organization and Operation
@@
 1 shows the hierarchical organization of a modern DRAM-based main memory.
-The memory controller connects to a DRAM module over a memory channel.
-A module contains one or more DRAM ranks that time-share the memory channel.
-A rank consists of multiple DRAM chips that operate in lock-step.
+A memory channel connects the processor (CPU) to a DRAM module, where a module consists of multiple DRAM ranks.
+A rank consists of multiple DRAM chips that operate in lockstep.
 Each DRAM chip contains multiple DRAM banks that can be accessed independently.
@@
 To open a row, the memory controller issues an activation command to a bank by specifying the row address to open.
-After activation is complete, the memory controller issues either read and write commands to read or write a DRAM word within the activated row.
+After activation is complete, the memory controller issues either read or write commands to read or write a DRAM word within the activated row.
 Subsequent accesses to the same row can be served quickly without fetching the data to the row buffer, creating a row hit [104].
@@
 Processing-in-Memory (PiM) moves computation closer to where the data resides, alleviating the data movement bottleneck between memory and the processor [13, 43, 8, 44, 45, 46, 11, 47, 19, 48, 1, 2].
-There are two main PiM approaches: 1) processing-near-memory (PnM) and 2) processing-using-memory (PuM).
-PnM place compute logic near memory arrays (e.g., processing capability in the memory controller and logic layer(s) of 3D-stacked memory [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 21, 39, 40, 41, 42]).
-PnM logic can execute application regions (e.g., instructions [67, 32], functions [28], application threads [20, 54]) depending on the design.
+There are two main PiM approaches [49, 10, 8]: 1) Processing-near-Memory (PnM) and 2) Processing-using-Memory (PuM).
+PnM places computation logic near memory arrays (e.g., in a DRAM chip, next to each bank, or at the logic layer of 3D-stacked memory [50, 51, 52, 53, 37, 54, 55, 38, 56, 57, 21, 58, 59, 60, 26]).
+The PnM logic can execute application regions (e.g., instructions [67, 53], functions [48], application threads [20, 77]) depending on the design.
 PnM computation units typically employ small caches due to 1) area constraints and 2) the ineffectiveness of caches in improving the performance of a wide range of memory-intensive applications.
-Therefore, various systems [32, 67, 28] offload the memory-intensive application regions on the PnM units, while executing more cache-friendly application regions on the host CPU with a deeper cache hierarchy.
+Therefore, various systems [53, 67, 48] offload the memory-intensive application regions on PnM units, while executing more cache-friendly application regions on the host CPU with a deeper cache hierarchy.
 PuM uses analog operational properties of the memory circuitry (e.g., DRAM) to enable massively parallel in-memory computation [70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85].
@@
 One such structure is the DRAM row buffer that acts as an internal cache inside DRAM banks.
-An attacker can understand if a row is recently accessed (i.e., already in the row buffer) or not by measuring the time it takes to access that row.
+An attacker can understand whether a row is recently accessed (i.e., already in the row buffer) by measuring the time it takes to access that row.
 Using this observation, DRAMA [98] builds a side-channel attack to leak keystrokes in a real system.
-Several prior works [68, 75, 76, 77, 78] leak DRAM address mapping functions.
-Other main memory-based attacks exploit the memory bus [79], and memory deduplication [80, 81] and (de)compression [82, 83, 84, 85, 86, 87, 88, 89, 90] latencies.
+Several prior works (e.g., [98, 105, 106, 107, 108]) leak DRAM address mapping functions.
+Other main memory-based attacks exploit the memory bus contention [109] and memory deduplication [110, 111] and (de)compression [112, 113, 114, 115, 116, 117, 118, 119, 120] latencies.
 3 A Case for PiM-based Timing Attacks
-The adoption of processing-in-memory (PiM) architectures has been gaining momentum in recent years because they provide high performance and low energy consumption by reducing data movement.
+The adoption of PiM architectures has been gaining momentum in recent years because they provide high performance and low energy consumption by reducing data movement.
 Therefore, it is timely and critical to analyze and evaluate the security of PiM architectures to prevent any vulnerabilities from causing widespread susceptibility as PiM architectures become widely available.
-Unfortunately, no prior work analyzes and evaluates the security of emerging PiM architectures against timing covert and side-channel attacks.
-After a thorough analysis of existing PiM architectures, we show that the adoption of PiM solutions creates opportunities for high-throughput main memory-based timing attacks for two main reasons.
-First, the adoption of PiM solutions provides user space applications with fast and reliable direct access to main memory.
-Second, timing attacks in PiM architectures can be more difficult to detect and mitigate due to commonly employed mechanisms, like cache monitoring mechanisms that track the cache miss ratio during runtime to detect anomalies [64, 65, 66], being either inapplicable or performance-inefficient in PiM architectures.
+Unfortunately, no prior work analyzes and evaluates the security of emerging PiM architectures against timing covert- and side-channel attacks.
+After a thorough analysis of existing PiM architectures, we show that adopting PiM solutions creates opportunities for high-throughput main memory-based timing attacks for two main reasons.
+First, the adoption of PiM solutions provides userspace applications with fast and reliable direct access to main memory to reduce data movement.
+Second, defenses against these attacks either incur high performance overheads or are not applicable to PiM architectures.
+In this work, we study the impact of PiM architectures on row buffer-based timing channels and their throughput.
 3.1 Row Buffer Timing Channel
 A DRAM row buffer acts as a direct-mapped cache inside a DRAM bank and is shared across applications that can access main memory.
-Accessing a row that is already in the row buffer results in a significantly lower latency because the memory controller does not need to issue additional commands to precharge the open row and fetch the target row to the row buffer.
+Accessing a row that is already in the row buffer results in a significantly lower latency because the memory controller does not need to issue additional commands (i.e., precharge the bank and fetch the target row).
 Therefore, a userspace application can infer whether a row has been accessed recently by measuring the time it takes to access the row.
-To demonstrate the latency gap between different row buffer states, we evaluate microbenchmarks that create memory accesses that result in row buffer hits and conflicts.
-We observe that a row buffer conflict takes 74 CPU cycles more than a hit, which is large enough to detect.
-So, a user space application can exploit the row buffer as a timing covert and side channel.
-Our results align with the observations of prior works [68, 77, 78].
+To demonstrate the latency difference between a row buffer hit and a row buffer conflict in a PiM-enabled system, we simulate such a system (described in detail in §5.1).
+We run a microbenchmark that 1) creates memory accesses that result in row buffer hits and conflicts, and 2) measures the memory access latency for each memory request.
+We observe that a row buffer conflict takes 74 CPU cycles more than a hit for a CPU operating at 2.6 GHz, which is large enough to detect, which aligns with the observations of prior works [98, 107, 108].
+Based on this observation, we conclude that a userspace application can exploit the row buffer as a timing channel in a PiM-enabled system.
 3.2 Existing Main Memory Attack Primitives
 Main memory timing attacks are powerful as they can be used to gather information on all applications that share main memory.
-However, these covert-channel attacks have limited throughput in traditional architectures for two main reasons: (i) accessing a memory location in main memory requires expensive cache bypassing techniques due to frequently accessed data being cached to improve performance, and (ii) even when bypassing the cache hierarchy, deep cache hierarchies employed in modern processors incur an additional high lookup latency on memory requests.
-There are several ways to bypass the cache hierarchy by using one or more of four attack primitives: (i) using specialized instructions to flush a cache line (e.g., clflush), (ii) using cache eviction sets (i.e., by creating many conflicting memory requests to replace the cache line with newly-fetched cache lines), (iii) memory access requests generated by (remote) direct memory access, (R)DMA, engines, and (iv) non-temporal memory hints (e.g., movnti).
-However, no attack primitive provides both fast and reliable communication.
+However, these attacks have limited throughput in traditional architectures because accessing main memory directly is difficult due to deep cache hierarchies.
+There are four main attack primitives that enables attackers to bypass the cache hierarchy and directly access main memory: 1) using specialized instructions to flush a target cache line (e.g., clflush [121]) from the cache deterministically, 2) using cache eviction sets [122] to replace a target cache line in the cache (i.e., by creating many conflicting memory requests to replace the cache line with newly-fetched cache lines), 3) using a (remote) direct memory access engine (i.e., (R)DMA engine) [123, 124], and 4) using non-temporal memory hints (e.g., movnti instruction [121]).
+We analyze the attack primitives in terms of efficiency and effectiveness.
+To do so, we identify four properties of an efficient (i.e., low-latency) and effective attack primitive: 1) avoiding cache lookup overhead, 2) avoiding high-latency memory accesses, 3) creating an easily detectable timing difference, and 4) reliable functionality (e.g., guaranteed to work by the ISA).
+Table 1 shows a comprehensive comparison of these four attack primitives and PiM operations in terms of these properties.
 Table 1: Efficiency and Effectiveness of Attack Primitives
@@
 Eviction Sets
-DMA/R-DMA N/A
+DMA/RDMA N/A
 Non-temporal
@@
 PiM Operations
-We identify four properties of an efficient (i.e., low-latency) and effective attack primitive: 1) avoiding cache lookup overhead, 2) avoiding high-latency memory accesses, 3) creating an easily detectable timing difference, and 4) reliable functionality (e.g., guaranteed to work by the ISA).
-Table 1 shows a comprehensive comparison of these four attack primitives in terms of these properties.
-First, specialized cache line flush instructions have two main disadvantages that hurt the efficiency and effectiveness of attacks that use them: 1) they put the write-back (to the main memory) latency of a cache line on the critical path to access a memory location, and 2) in some modern systems, these instructions are privileged and cannot be used by user space applications.
-Second, eviction sets require many load requests proportional to the number of ways in the caches, which scales with modern designs.
+First, specialized cache flush instructions have two main disadvantages that limit their effectiveness at directly accessing main memory: 1) they put the write-back (to the main memory) latency of a cache line on the critical path to access a memory location [121], and 2) in some modern systems, these instructions are privileged and cannot be used by userspace applications (e.g., ARM processors [125]) (not shown in the table).
+Second, to directly access main memory, cache eviction sets require many load requests proportionally to the number of ways in the caches, which increases with modern designs.
 Therefore, the attacks that use them have limited throughput.
-While the instructions used to create the eviction sets are part of the ISA, the eviction set is not guaranteed to evict the target cache line due to the cache replacement policy, cache prefetchers, and other design choices that are not transparent to the attacker.
-Third, DMA and R-DMA engines incur high overheads due to deep software stacks.
+Even though the instructions used to create the eviction sets are part of the ISA, the eviction set is not guaranteed to evict the target cache line due to the cache replacement policy, cache prefetchers, and other memory system design choices that are not transparent to the attacker.
+Third, DMA and RDMA engines [123, 124] incur high overheads due to deep software stacks.
 As a result, attacks using this primitive need to utilize states with larger timing differences to be distinguishable.
-An example attack using this primitive[91] showcases a covert-channel attack based on memory contention.
-However, the basis of this type of covert channel is not as fine-grained as row buffer-based timing differences.
+An example work using this primitive [126] showcases a covert-channel attack based on memory contention.
+However, memory contention-based timing differences is not as fine-grained as row buffer-based timing differences.
 Therefore, the communication throughput of these covert channels is limited (in the order of Kb/s).
-Fourth, the implementation of non-temporal memory hints varies across systems, as system designers are not bound to a specific approach.
-Consequently, the ISA does not guarantee that memory accesses related to non-temporal hints will consistently bypass caches.
-In contrast, PiM operations provide fast and reliable direct memory access to user space applications, which can be exploited for high-throughput covert- and side-channel attacks.
+Fourth, the implementation of non-temporal memory hints varies across systems, as system designers are not bound by the instruction set architecture.
+These instructions may prefetch data from the last level cache (prefetchnta), or can be cached in intermediate buffers (e.g., movnti instruction) [127].
+Consequently, the ISA does not guarantee that non-temporal hints will consistently and directly access main memory.
+In contrast, PiM operations provide fast and reliable direct memory access to userspace applications, which can be exploited for high-throughput covert- and side-channel attacks, as we show in this paper.
 3.3 High Throughput Main Memory-based Covert Channels
-To demonstrate the impact of the deep cache hierarchy latencies (i.e., cache access and eviction latencies) on an attack’s throughput, we simulate a realistic system (described in §5.1), establishing a state-of-the-art covert channel [68].
-Attack Scenarios.
+To demonstrate the impact of the deep cache hierarchy latencies (i.e., cache access and eviction latencies) on an attack’s throughput, we simulate a PiM-enabled system (described in §5.1), establishing a state-of-the-art row buffer-based covert channel [98].
+3.3.1 Attack Scenarios
 We simulate two attacks.
-First, we evaluate a baseline attack that relies on cache bypassing using cache eviction sets.
-To transmit a bit through the row buffer, the baseline attack first evicts the corresponding cache line, and then it fetches the target row into the row buffer using an additional memory request.
-The baseline attack accomplishes cache line eviction by issuing N memory requests to the target cache set in an N-way cache.
-Note that the actual eviction latency can be much higher due to a larger number of memory accesses required to bypass more complex replacement policies employed in today’s systems or due to address translation overheads.
+First, we evaluate a baseline main memory attack that relies on cache bypassing using cache eviction sets.
+To transmit messages through the row buffer, the sender and receiver processes 1) evict the corresponding cache line to directly access main memory and 2) fetch the target row into the row buffer using an additional memory request.
+The baseline attack evicts a cache line by issuing N memory requests to the target cache set in an N-way cache.
+The actual eviction latency in a real system can be much higher due to a larger number of memory accesses required to bypass more complex replacement policies employed in today’s systems [128, 129] or due to address translation overheads [130, 131, 132, 133, 134, 135, 136, 137].
 Second, we evaluate a direct memory access attack that issues only one memory request to fetch the data to the row buffer without any cache access or evictions.
-For both attacks, we measure the time it takes to send a bit through the row buffer, including cache access and eviction latencies, with encoding logic-1 values as row buffer hits and logic-0 as row buffer conflicts.
-We calculate the communication throughput based on the average time needed to propagate bits from the sender to the receiver.
+For both attacks, we measure the time it takes to send a bit through the row buffer, including cache access and eviction latencies, encoding logic-1 values as row buffer hits and logic-0 as row buffer conflicts.
+We calculate the communication throughput based on the average time to propagate bits from the sender to the receiver.
 We analyze the impact of (i) the LLC size and (ii) the number of LLC ways on communication throughput to show the limitations in current and future systems with different cache designs.
-Impact of LLC Size on Communication Throughput.
+3.3.2 Impact of LLC Size on Communication Throughput
 Fig.
-2 shows the impact of the LLC size on (i) the throughput of baseline and direct memory access attacks and (ii) the latency of a cache line eviction.
+2 shows the impact of the LLC size on (i) the throughput of the baseline and direct memory access attacks on the primary y-axis (left) and (ii) the latency of a cache line eviction on the secondary y-axis (right).
 The x-axis shows the cache configuration with increasing last-level cache (LLC) sizes and a fixed number of ways (i.e., 16).
-To calculate the cache access latency with increasing LLC sizes, we followed the same methodology used in prior works [92].
+We calculate the cache access latency with increasing LLC sizes using CACTI 6.0 [138].
+Figure 2: Impact of the LLC size on data leakage throughput of covert channels and LLC eviction latency.
 We make two key observations.
-First, the communication throughput of the direct memory access attack is $11.27$ Mb/s across all LLC sizes, which is significantly greater than the throughput of up to $2.29$ Mb/s achieved by the baseline attack.
+First, the communication throughput of the direct memory access attack is $11.27$ Mb/s across all LLC sizes, which is significantly greater than the throughput of up to $2.29$ Mb/s achieved by the baseline cache eviction sets based attack.
 Second, the baseline attack’s throughput decreases with increasing LLC sizes since the cache access and eviction become progressively more expensive.
-With increasing LLC size, the cache access latency increases and, thus, amplifies the cost of eviction, which requires multiple cache accesses.
-Figure 2: Impact of the LLC size on the data leakage throughput of covert channels and the eviction latency.
-Impact of LLC Associativity on Communication Throughput.
+With increasing LLC size, the cache access latency increases, amplifying the latency of eviction, which requires multiple cache accesses.
+Therefore, we conclude that with increasing LLC sizes, cache access and eviction latencies increase, hindering the baseline attack’s communication throughput, compared to the direct memory access attack, which is independent of the LLC size.
+3.3.3 Impact of LLC Associativity on Communication Throughput
 Fig.
-3 shows the impact of LLC associativity on (i) the communication throughput of baseline and direct memory access attacks and (ii) the eviction latency of a cache line.
-The x-axis shows the increasing number of LLC ways (from 2 to 128) while the total LLC size is fixed at 16 MB (modern processors typically employ 8 cores and a 2MB LLC slice per core).
-Increasing the number of LLC ways increases the number of memory requests required to evict a cache line from the cache.
-We observe that the baseline attack’s communication throughput decreases significantly with an increasing number of LLC ways due to the higher cache eviction latency.
-Figure 3: Impact of the number of LLC ways on the data leakage throughput of covert channels and the eviction latency.
-We conclude that attacks that communicate with main memory through caches are greatly limited in throughput due to expensive cache access and eviction latencies, which worsen with increasing size and associativity.
-However, a covert channel based on direct memory access can achieve high throughput independent from cache configurations as it does not require bypassing the cache hierarchy.
-The adoption of PiM solutions can fundamentally enable high-throughput timing attacks through main memory by providing direct access to the main memory to user space applications.
-Therefore, it is crucial to carefully consider the security implications of PiM architectures and develop robust security measures to protect against potential attacks.
+3 shows the impact of LLC associativity on (i) the communication throughput of baseline and direct memory access attacks on the primary y-axis (left) and (ii) the eviction latency of a cache line on the secondary y-axis (right).
+The x-axis shows the increasing number of LLC ways (from 2 to 128) with a fixed total LLC size (e.g., 16 MB).
+We observe that as the number of LLC ways increases, the number of memory requests required to evict a cache line increases, greatly reducing the baseline attack’s communication throughput.
+Figure 3: Impact of LLC associativity on the data leakage throughput of covert channels and the eviction latency.
+Based on these two experiments, we conclude that attacks that communicate with main memory through caches are greatly limited in throughput due to expensive cache access and eviction latencies, which worsen with increasing size and associativity.
+In contrast, a covert channel based on direct memory access can achieve high throughput independently of the cache configuration, as it does not require bypassing the cache hierarchy.
 4 IMPACT: A Set of High-Throughput Main Memory Timing Attacks
-This section introduces IMPACT, a set of high-throughput main memory-based timing attacks that take advantage of the direct memory access enabled by the emerging processing-in-memory paradigm.
+We introduce IMPACT, a set of high-throughput main memory-based timing attacks that take advantage of the direct memory access enabled by emerging PiM architectures.
 IMPACT achieves high communication and information leakage throughput by (i) eliminating expensive cache bypassing steps required by other main memory- and cache-based timing covert- and side-channel attacks and (ii) leveraging the intrinsic parallelism of PiM operations.
-In this work, we present two covert-channel and one side-channel attacks enabled by processing-near-memory (PnM) and processing-using-memory (PuM) solutions.
+In this work, we present two covert-channel and one side-channel attacks enabled by PnM and PuM solutions.
 4.1 PnM-based IMPACT Covert-Channel Attack
 Baseline PnM Architecture.
-PiM-Enabled Instructions (PEI)  [67] is a PnM architecture that expands the instruction set architecture of the host processor with PIM-enabled instructions (PEIs).
-PEIs can be executed either on the host processor or on the in-memory computation logic.
+PiM-Enabled Instructions (PEI)  [67] is a PnM architecture that expands the instruction set architecture with instructions that can be executed in compute units in the processor core or near memory based on a dynamic decision made considering the locality of the required memory accesses.
 We choose PEI as the baseline PnM architecture as it introduces a simple and effective model for PnM, and our attack can be generalized for other PnM architectures with similar design components (e.g., FIMDRAM [89]).
@@
 First, it integrates a PEI Computation Unit (PCU) near each DRAM bank and inside the CPU.
-In-memory PCUs are shared among all the cores of the host processor and are responsible for executing PEIs near the respective DRAM bank.
-Second, the PEI architecture has a PEI Management Unit (PMU) that monitors data locality in application regions and decides where to map each region (either to host-side or memory-side PCUs).
-This way, PEI architecture can execute regions with high locality on the host side to benefit from caches.
-PEI adds an ignore flag per locality monitor entry to ignore the first hit in the locality monitor [93].
-This is because if a locality monitor entry is allocated for a PIM operation, it is often too aggressive to consider the operation as having a high locality on its first hit.
-This bit allows IMPACT-PnM to easily bypass the locality monitor even with a small virtual address range at its disposal.
+In-memory PCUs are shared among all cores of the host processor and are responsible for executing PEIs near the respective DRAM bank.
+Second, there is a PEI Management Unit (PMU) that monitors data locality in application regions with a locality monitor, and decides where to map each region (either to host-side or memory-side PCUs).
+This way, the PEI architecture can execute application regions with high data locality on the host side to benefit from the cache hierarchy.
+In contrast, if an application region does not benefit from the cache hierarchy, it is executed on the memory-side PCU to benefit from the lower memory access latency.
 Attack Overview.
-Before the attack, the sender and the receiver co-locate their data in the same DRAM banks.
-To achieve this, one process uses memory massaging techniques to place its data in the same bank as the other process in main memory, as used in many prior works [76, 68, 94, 95].
+Before the attack, the sender and the receiver co-locate their data in the same set of DRAM banks.
+To achieve this, one process uses memory massaging techniques to place its data in a page located in the same DRAM bank as the other process, as used in many prior works (e.g., [106, 98, 139, 140]).
+To incentivize the PnM architecture to execute the instruction in a memory-side PCU and directly access DRAM, the attack processes access different cache lines in the allocated rows, and allocate and use multiple rows in each bank.
 IMPACT-PnM uses PEI operations and exploits the latency difference between a DRAM row buffer hit and a conflict in a shared DRAM bank to exchange messages between the sender and the receiver.
 The sender transmits bits of the message to the receiver in multiple batches of M bits, where the value of each bit is encoded as interference in the row buffer of a DRAM bank.
-In this attack, we encode logic-1 as interference in the DRAM bank (i.e., a row buffer conflict), and logic-0 as no interference (i.e., a row buffer hit).
+In this attack, we encode logic-1 as interference (i.e., a row buffer conflict), and logic-0 as no interference (i.e., a row buffer hit).
 Fig. 4 and Listing 1 show the end-to-end flow of IMPACT-PnM and all the steps involved in establishing the covert channel between the sender and the receiver.
 Figure 4: Flow of the PnM-based Covert Channel Attack.
-1 Sender:
-2 message = [0: N-1]; //N banks and N-bit msg.
-3 batch_size = M; //M-bits per batch where M<N
-4 init_semaphore(0);
-5 barrier(); // synchronize with receiver
+1 Receiver (Step 1):
+2 init_row = initialize_DRAM_rows_with_PEIs();
+3
+4 Sender (Step 2):
+5 message[0: N-1]; // N-bit message
+6 batch_size = M; // M-bit batch, where M<N
+7 for every batch:
 8 bank = 0;
-7 for every batch:
 9 for each bit in batch_size;
-9 if message[bank] == 1:
-10 pim_add(row[bank]);
-11 else:
+10 if message[bank] == 1: // activate a row
+11 pim_add(sender_row[bank]);
+12 else: // do not interfere with the receiver
 13 NOP();
 14 bank++;
-14 memory_fence(); // signal receiver
-15 increment_semaphore(); // to start probing
-16
-17 Receiver:
-18 init_row = initialize_DRAM_rows_with_PEIs();
+15 // ensure all memory requests are served
+16 // before starting the next batch
+17 memory_fence();
+18
+19 Receiver (Step 3):
+20 for every batch:
 21 bank = 0;
-20 barrier(); // synchronize with sender
-21 for every batch:
-22 block_on_semaphore();
-23 //if semaphore==0, wait
 22 for each bit in batch_size;
@@
 26 latency = end - start;
-29 if latency > threshold:
+27 if latency > THRESHOLD: // row buffer conflict
 28 result = ’1’;
-31 else:
+29 else: // row buffer hit
 30 result = ’0’;
 31 bank++;
+32 memory_fence(); // complete the batch
 Listing 1: Pseudocode of PnM-based Covert-Channel Attack.
-Step 1.
-The receiver issues PEIs to initialize each DRAM bank by activating a predetermined row (Line 18).
-The receiver either uses unused cache lines of the activated row or sets the ignore flag in the corresponding entry in order to bypass the locality monitor.
-Then, the sender and the receiver synchronize on a barrier (Lines 5 and 20), which ensures that they start the communication turn at the same time.
+Step 1. The receiver issues PEIs to initialize each DRAM bank by activating a predetermined row (Line 2, 1). Initialization copies the receiver’s data to row buffers.
 Step 2.
-The sender starts the communication turn by issuing PEI instructions to activate a row in each DRAM bank (Lines 7-12) or not based on the bit value that will be transmitted.
-To transmit a 0 ( 1), the sender does not activate any row in the corresponding DRAM bank to avoid causing interference in the row buffer (e.g., Row 0xB remains in the row buffer  2).
-To transmit a 1 ( 3), the sender activates a different row to cause interference in the corresponding DRAM bank (i.e., a row buffer conflict between Row 0xD and Row 0xE  4).
-At the end of the communication turn (after transmitting M bits), the sender first executes a memory fence (Line 14) to ensure that all the PEI instructions have been executed and then signals the receiver to start probing the DRAM banks (Line 15).
+The sender transmits a batch of the message by mapping each bit in the batch to a separate DRAM bank ( 2).
+Then, the sender checks the bit value (Line 10).
+If the bit is 1, the sender issues a PEI to activate a row (Line 11, 3) and thus, copies its content to the row buffer ( 4).
+Otherwise, the sender issues a NOP instruction (Lines 12-13, 5) and does not copy any new data to the row buffer ( 6).
 Step 3.
-The receiver starts probing the DRAM banks by issuing PEI instructions to activate the same row as the one activated during initialization (Line 25).
-Then, the receiver measures the time it takes to complete the request (Line 27).
-If the latency exceeds a predetermined hit threshold, the receiver senses the interference in the row buffer and decodes the bit as a 1 (Lines 28-29  5, 6).
-Otherwise, the access is considered a row buffer hit and the receiver decodes the bit as a 0 (Lines 30-31  7, 8).
+The receiver probes the DRAM banks by issuing PEIs to the initialized rows and measuring the PEI latencies (Lines 23-25, 7).
+If the latency of a PEI exceeds a predetermined row buffer hit threshold, the receiver detects interference and decodes the bit as a 1 (Lines 27-28, 8).
+Otherwise, the access is considered a row buffer hit, and the receiver decodes the bit as a 0 (Lines 29-30, 9).
+After the transmission of a batch of the message (i.e., M bits), the sender and the receiver execute a memory fence (Lines 17 and 32) to ensure that all the PEIs are executed before moving on to the next batch.
+The receiver accesses the next cache line in the initialized row to ensure its PEI is executed near memory by bypassing the locality monitor.
+If the number of batches is greater than the number of cache lines in a row, it uses another row in the same bank to measure access latency.
 The sender and the receiver repeat the same steps for each batch until the entire message is transmitted.
@@
 The sender and the receiver synchronize to 1) ensure their actions do not interfere with each other and 2) overlap the latencies of their operations to increase the throughput of the attack.
-To do so, the sender and the receiver use a semaphore.
-The value of the semaphore indicates the number of batches that the sender has completed transmitting, but the receiver has not yet probed.
+In our proof-of-concept implementation, the sender and the receiver use a semaphore.
+The semaphore’s value indicates the number of batches that the sender has completed transmitting, but the receiver has not yet probed.
 The receiver blocks on the semaphore until the sender increments it (i.e., does not start probing).
@@
 Upon receiving the signal, the receiver decrements the semaphore and starts probing the DRAM banks.
-In parallel, the sender can start transmitting the next batch that utilizes a different set of DRAM banks.
-This way, they can overlap the transmission latency and the probing latency.
 When the receiver completes probing, they read the semaphore again, and the process repeats until the entire message is transmitted.
-In our attack implementation (§5.1), we use POSIX semaphores to implement the sender-receiver synchronization and use a batch size of 4 bits.
 4.2 PuM-based IMPACT Covert-Channel Attack
 Baseline PuM Architecture.
-We assume a PuM architecture that provides user applications with RowClone [52] that enables bulk data movement operations (e.g., bulk data copy and initialization) within DRAM’s subarray boundaries by exploiting the internal operations of DRAM.
-[^2] There could be near-memory RowClone variants that move data without using operational principles of DRAM as well.
-However, in this attack, we assume a PuM-based RowClone operation.
-RowClone’s Fast Parallel Mode works by issuing back-to-back two activation commands: the first activation copies the content of a source row (i.e., src) into the row buffer.
-The second command connects the destination row (i.e., dst) to the bitlines.
-Since the row buffer already stored the src’s data, by the time the dst is activated, the data in dst is overwritten by src’s data.
-To execute RowClone, we assume that the user space application specifies a source and a destination virtual address range and the corresponding mask to copy only specific parts of the source to the destination.
+We assume a PuM architecture that provides user applications with RowClone [75] which enables bulk data copy and initialization operations within DRAM’s subarray boundaries by exploiting the internal operations of DRAM.
+RowClone’s Fast Parallel Mode works by issuing two activation commands back-to-back: the first activation copies the content of a source row (i.e., src) into the row buffer.
+The second connects the destination row (i.e., dst) to the bitlines.
+Since the row buffer stores the src’s data after the first activation, by the time the dst is activated, the data in dst is overwritten by src’s data.
+To execute RowClone, we assume that the userspace application specifies 1) a source virtual address range, 2) a destination virtual address range, and 3) a mask to copy specific parts of the source range to the destination range.
 This way, the memory controller receives one RowClone request and breaks it into parallel RowClone requests, one for each set bit of the mask.
 Attack Overview.
-PuM-based covert channel attack (IMPACT-PuM) exploits in-DRAM bulk copy operations to establish a covert channel between a sender and a receiver.
-The sender transmits a message by executing a RowClone operation with three operands: (i) a source address range, (ii) a destination address range, and (iii) a mask to select which parts of the ranges will be copied.
-Source and address ranges correspond to physical pages spanning across all DRAM banks and by using the mask, the sender can encode the transmitted message by causing contention only in selected DRAM banks.
-The receiver repeatedly issues RowClone operations and measures the latency of these operations.
-This way, if the sender issues a RowClone operation to a DRAM bank, it induces contention, and the receiver detects a higher latency for its RowClone operation, indicating the transmitted bit is logic-1.
-Otherwise, the receiver detects a lower latency for the RowClone operation, indicating that the corresponding bit is a logic-0.
-Fig. 5 and Listing 2 show the end-to-end flow of IMPACT-PuM and all the steps involved to establish the covert channel between the sender and the receiver.
-Figure 5: PuM covert channel attack flow.
-Step 1.
-First, the receiver initializes all DRAM banks by issuing a RowClone operation (Line 16).
-Then, the sender and the receiver synchronize on a barrier (Lines 10 and 17) which ensures that they start the communication turn at the same time.
+Our PuM-based covert channel attack (IMPACT-PuM) exploits in-DRAM bulk copy operations to establish a covert channel between a sender and a receiver.
+Before the attack, the sender and the receiver co-locate their data in the same set of DRAM banks.
+The sender transmits a message to the receiver in multiple batches of M bits.
+To do so, the sender issues RowClone operations by carefully selecting 1) source and destination address ranges corresponding to physical pages spanning the selected set of DRAM banks, and 2) a mask to cause interference only in selected DRAM banks.
+The sender and receiver processes synchronize with barriers similar to IMPACT-PnM (§4.1)
+Fig. 5 and Listing 2 show the end-to-end flow of IMPACT-PuM and the steps for establishing the covert channel between the sender and the receiver.
+Figure 5: PuM covert-channel attack flow.
+Step 1. First, the receiver initializes the DRAM banks by issuing a RowClone operation (Line 25, 1).
 Step 2.
-The sender uses a single RowClone operation to transmit an N-bit message (where N is the number of banks) to the receiver in parallel and uses the mask of the RowClone operation to encode the message (Lines 11-12).
-The sender encodes the message by setting the corresponding bits in the mask to 1 or 0 ( 1, 2).
-To send a 1, the sender triggers RowClone in the corresponding DRAM bank and induces contention by altering the activated row in the row buffer ( 3).
-To send a 0, the sender does not trigger RowClone in the corresponding DRAM bank and does not induce contention ( 4).
-The sender synchronizes with the receiver using a barrier operation (Lines 13 and 18).
+The sender transmits the message by dividing the message into M-bit batches.
+Then, the sender maps each bit of a batch to a separate DRAM bank and creates a mask for the batch (Line 18), determining which banks will be executing a RowClone operation (Line 18).
+By using a mask, the sender uses only a single RowClone operation to transmit an M-bit batch of the message (where M is the number of banks) (Line 20, 2).
+The memory controller issues individual RowClone operations to each DRAM bank based on the specified mask (Lines 1-9).
+To send a 1, the sender triggers RowClone in the corresponding DRAM bank and activates a different row from the receiver’s row in the bank (Lines 4-7, 3).
+To send a 0, the sender does not trigger RowClone in the corresponding DRAM bank and does not interfere with the receiver’s activated row (Lines 8-9, 4).
+Step 3.
+The receiver detects interference in DRAM banks using RowClone operations as follows.
+The receiver issues one RowClone operation for each DRAM bank to precisely measure the time it takes to complete the operation in each bank (Lines 29-33).
+To detect if the sender activated another row, the receiver uses its initial destination address range as the source (i.e., swapping the direction of the copy operation) (Line 31).
+If the sender induces interference in a DRAM bank, the receiver observes high latency for the RowClone operation, indicating the corresponding bit in the message is set to 1 (Lines 34-35, 5).
+Otherwise, the receiver detects a fast RowClone operation, which indicates that the bit is set to 0 (Lines 36-37, 6).
 1 Memory Controller:
 2 row_clone(src, dst, mask):
-3 for each DRAM bank do:
-4 if mask[bank] == 1 then:
-5 row_clone(bank, src, dst);
+3 for each DRAM bank:
+4 if mask[bank] == 1:
+5 // copy the corresponding src row
+6 // to the destination row
+7 row_clone(src, dst, bank);
 8 else:
-7 NOP();
-8
+9 NOP(); // do not interfere with the receiver
+10
 11 Sender:
-10 barrier_1();
-11 mask = init(message);
+12 message[0:N-1]; // N-bit message
+13 batch_size = M; // M-bit batch, where M<N
+14 start = 0;
+15 for every batch:
+16 end = start + batch_size;
+17 // create mask for the batch
+18 mask = message[start:end];
+19 // issue a RowClone operation using the mask
 20 row_clone(src, dst, mask);
-13 barrier_2();
-14
+21 start = end; // update start for the next batch
+22 memory_fence(); // complete the batch
+23
 24 Receiver:
 25 init_DRAM_rows_with_RowClone();
-17 barrier_1();
-18 barrier_2();
-19 for each DRAM bank do:
+26 for every batch:
+27 for each DRAM bank:
 28 mask = 1 << bank;
 29 timer(start);
-22 row_clone(src, dst, mask);
+30 // issue RowClone to one bank at a time
+31 row_clone(dst, src, mask);
 32 timer(end);
 33 latency = end - start;
-25 if latency > threshold:
+34 if latency > THRESHOLD: // row buffer conflict
 35 result = ’1’;
-27 else:
+36 else: // row buffer hit
 37 result = ’0’;
+38 memory_fence(); // complete the batch
 Listing 2: Pseudocode of PuM-based Covert-Channel Attack.
-Step 3.
-The receiver in IMPACT-PuM detects contention in DRAM banks by issuing RowClone operations and measuring their latency.
-The receiver continuously issues one RowClone operation for each DRAM bank (in contrast to the sender who issues only a single RowClone operation), to the initial virtual address range (i.e., rows initialized in Step 1) and measures the latency of each operation ( 2, Line 24).
-If the sender induces contention in a DRAM bank, the receiver observes high latency for the RowClone operation which indicates that the corresponding bit in the message is set to 1 ( 3, Lines 25-26).
-Otherwise, the receiver detects a fast RowClone operation which indicates that the bit is set to 0 ( 4, Lines 27-28).
-Advantage over IMPACT-PnM.
-In IMPACT-PuM, the sender can use a single RowClone operation to transmit an N-bit message to the receiver in parallel, where N is the number of DRAM banks.
-Thus, IMPACT-PuM requires less computational resources to transmit the same amount of information compared to IMPACT-PnM.
-For example, using a single thread, the sender can activate all N DRAM banks and transmit N bits of information in parallel.
-In contrast, in IMPACT-PnM, the sender needs to execute one PEI per bit.
-4.3 PnM-based Side-Channel Attack on Read Mapping
+Advantage over IMPACT-PnM (§4.1).
+In IMPACT-PuM, the sender can use a single RowClone operation to transmit an M-bit message to the receiver in parallel, where M is the number of DRAM banks.
+Thus, IMPACT-PuM requires fewer computational resources to transmit the same amount of information compared to IMPACT-PnM.
+For example, using a single thread, the sender can activate all M DRAM banks and transmit M bits of information in parallel.
+In contrast, in IMPACT-PnM (§4.1), the sender needs to execute one PEI for every bit.
+4.3 PnM-based Side-Channel Attack on Genomic Read Mapping
 We build a PnM-based side-channel attack that exploits PiM operations to leak private information of a concurrently running application.
 To demonstrate the IMPACT side-channel attack, we target a DNA sequence analysis application.
-Read Mapping.
-DNA is one of the unique identifiers of an individual [73], such that even a small fraction of DNA can include sufficient information for linking to the owner’s identity [70, 71].
+Genomic Read Mapping.
+DNA is a unique identifier of an individual [103], such that even a small fraction of DNA can include sufficient information for linking a DNA sample to the owner’s identity [100, 101].
 Therefore, preserving privacy while processing genomic data carries utmost importance to protect sensitive information.
-Sequencing the DNA of a biological sample produces short DNA fragments called reads.
-The reads are then processed by read mapping, which is a fundamental operation to find the matching locations of reads in the reference genome that acts as a template to construct the complete genome of the sample.
-Read mapping (RM) has two key steps: (i) seeding and (ii) alignment.
-During the seeding step, RM identifies the possible locations of the reads on the reference genome based on sequence similarities between the reads and the reference genome.
-Seeding involves hashing small segments of the DNA sequences (i.e., seeds) and probing a hash table that is built from the reference genome.
-After seeding, RM aligns the reads with their candidate regions on the reference (i.e., potential reference segments) using dynamic programming-based methods to find each read’s best matching position among all possible locations in the alignment step.
+Sequencing the DNA of a biological sample produces short DNA fragments called reads [141].
+The reads are then processed by read mapping [34, 99, 36, 100, 101, 102, 103], which is a fundamental task performed to find the matching locations of reads in the reference genome that acts as a template to construct the complete genome of the sample.
+Read mapping (RM) has two key steps: (i) seeding [142] and (ii) alignment [143, 144].
+Fig.
+6 shows a high-level overview of RM.
+During the seeding step, RM identifies the possible locations of the reads in the reference genome based on sequence similarities between the reads and the reference genome.
+Seeding involves hashing small segments of the DNA sequences (i.e., seeds) ( 1) and probing a hash table that is built from the reference genome ( 2).
+The hash table provides potential locations for the seed in the reference genome ( 3).
+In the alignment step, RM aligns the reads with their candidate regions in the reference (i.e., potential reference segments) using dynamic programming methods [99, 100] to find each read’s best matching position among all possible locations ( 4).
+Figure 6: DNA Read Mapping Operation: Seeding and Sequence Alignment.
 RM requires a large amount of data movement to access different parts of the hash table and corresponding regions on the reference genome.
-Hence, several prior works aim to improve the performance of RM by offloading seeding and alignment to PiM-enabled systems to reduce the data movement overheads [33, 96, 97, 98, 99, 100, 101, 102].
-In this work, we demonstrate that an attacker can launch a high-throughput side-channel attack against RM by utilizing PiM instructions and quickly leak a large amount of information about the sample genome.
+Hence, several prior works aim to improve the performance of RM by offloading seeding and alignment to PiM-enabled systems to reduce the data movement overheads [37, 145, 146, 40, 147, 148, 149, 150, 35, 151, 152].
+In this work, we demonstrate that an attacker can launch a high-throughput side-channel attack against RM by utilizing PiM instructions and quickly leak a large amount of information about the query genome.
 Attack Overview.
-We assume that the victim is running a RM application on a PiM-enabled system and the attacker is able to run a malicious application on the same system.
+We assume that the victim is running an RM application on a PiM-enabled system and the attacker is able to run a malicious application on the same system.
 The attacker and the victim are using the same read mapping tool (e.g., minimap2 [153]).
@@
 We assume the hash table is distributed across multiple DRAM banks.
-This assumption is realistic as many modern DRAM address mapping schemes (e.g., [104, 105, 106, 107]) aim to interleave consequent memory chunks across different DRAM banks to exploit bank-level parallelism [108, 109].
+This assumption is realistic, as many modern DRAM address mapping schemes (e.g., [154, 155, 156, 157]) aim to interleave consequent memory chunks across different DRAM banks to exploit bank-level parallelism [158, 159].
 Fig.
 7 shows the high-level overview of the attack.
-The victim application extracts the seeds from the reads and offloads the corresponding seeding and alignment step to the PiM-enabled system  1.
-During the seeding step, the corresponding DRAM row gets activated (i.e., Index 0-5 of the hash table)  2.
-The attacker continuously uses PnM instructions to probe the shared hash table in an attempt to identify if the victim application is accessing a specific index of the hash table  3.
+The hash table is distributed across banks as explained above ( 1).
+During the attack, the victim application extracts seeds from the reads and offloads the corresponding seeding and alignment steps to the PiM-enabled system ( 2).
+During the seeding step, the corresponding DRAM row gets activated (i.e., Index 0-M of the hash table) ( 3).
+The attacker continuously uses PnM instructions to probe the shared hash table in an attempt to identify if the victim application is accessing a specific index of the hash table ( 4).
 If the attacker observes that the victim application is accessing a specific index, it retrieves partial or exact information about the potential locations of the reads in the reference genome.
-As shown in previous works on DNA imputation [110, 111, 112, 113], the attacker can use the leaked information in a completion attack [71] to infer properties about some regions of the private sample genome  4.
-Figure 6: High-level overview of IMPACT attack model on PiM-enabled RM.
-5 Evaluation
-5.1 Methodology
+As shown in previous works on DNA imputation [160, 161, 162, 163], the attacker can use the leaked information in a completion attack [101] to infer properties about some regions of the private query genome ( 5).
+Figure 7: High-level overview of the IMPACT attack model on PiM-enabled genomic read mapping.
+5 Methodology
+5.1 Threat Models
 Covert Channel Threat Model.
@@
 PiM instructions are available to userspace applications without any restrictions.
-The receiver has access to cpuid and rdtscp instructions, enabling high-precision measurement of memory access latencies.
-For IMPACT-PnM, the sender and the receiver co-locate data in the same DRAM banks.
-For IMPACT-PuM, (i) the sender allocates two virtual address ranges whose corresponding physical pages span across all DRAM banks, and (ii) RowClone guarantees atomicity: when the sender issues a RowClone operation that spans across multiple banks, all individual bank-level RowClone operations need to be completed before another DRAM operation is executed.
+The receiver has access to cpuid and rdtscp instructions [121], enabling high-precision measurement of memory access latencies.
+For IMPACT-PnM, the sender and the receiver co-locate data in the same set of DRAM banks.
+For IMPACT-PuM, the sender allocates two virtual address ranges whose corresponding physical pages span across the same set of DRAM banks.
 Side Channel Threat Model.
 We assume a scenario where an attacker process operates on the same PiM-enabled system as a victim process and leaks information about the victim process.
-We use an RM implementation based on minimap2 [103], and we modify the implementation to offload the seeding and alignment step to the PiM-enabled system.
-We assume that the alignment step includes chaining.
-We use the human genome as the reference and compare it against synthetic sample genomes.
+We use a genomic read mapping implementation based on minimap2 [153], and we modify the implementation to offload the seeding and alignment steps to the PiM-enabled system [67].
+We assume that the alignment step includes chaining [99, 153].
+We use the human genome as the reference and compare it against synthetic query genomes.
 We experiment with multiple seed sizes and report the best-performing selection.
-Proof-of-Concept.
+5.2 Evaluation Methodology
+5.2.1 Proof-of-Concept
 We construct a proof-of-concept of all IMPACT variants using the Sniper Multicore Simulator [164].
 [^2] There is no commercially available system that follows a similar architecture as PEI or any modern system that enables end-to-end RowClone operations.
-We extend Sniper to (i) accurately model all internal DRAM operations, (ii) emulate and simulate the PiM interface and PiM operations, and (iii) emulate the functionality of the cpuid and the rdtscp instructions to measure execution cycles from the user space.
+We extend Sniper to (i) accurately model all internal DRAM operations, (ii) emulate and simulate the PiM interface and PiM operations, and (iii) emulate the functionality of the cpuid and the rdtscp instructions [121] to measure execution time.
 Table 2 shows the configuration of our simulation setup.
-Before launching the attack, the sender and the receiver threads warm up by executing both attacks to avoid compulsory DRAM accesses for instructions or page table walks.
-For IMPACT-PnM, we assume that both the attacker and the receiver use PEIs to perform an operation that takes $\sim 3$ cycles to execute (in addition to the actual DRAM access).
-Comparison Points for Covert-Channel Attacks. To showcase the effectiveness of IMPACT, we evaluate the leakage throughput of seven different covert-channel attacks.
-(i) DRAMA-clflush: The row buffer-based covert-channel attack proposed in [68] which uses clflush instructions to send memory requests directly to main memory.
+Before launching the attack, the sender and the receiver threads warm up to avoid compulsory DRAM accesses for instructions or page table walks.
+For the PnM system [67], we model the additional latency of a PEI (e.g., latency of accessing additional system structures) as $3$ cycles [67].
+Table 2: Simulated System Configuration
+CPU 4-core, 4-way issue, OoO x86 core, 2.6GHz
+MMU L1 DTLB (4KB): 64-entry, 4-way, 1-cycle, L1 DTLB (2MB): 32-entry, 4-way, 1-cycle, L2 TLB: 1536-entry, 12-way, 12-cycle
+L1 Cache I-cache: 32 KB; L1D Cache: 32 KB, 8-way, 4-cycle, LRU, IP-stride prefetcher [165]
+L2 Cache 2 MB, 16-way, 16-cycle, SRRIP [166], Streamer prefetcher [167]
+L3 Cache 2 MB/core, 16-way, 50-cycle, SRRIP [166]
+Main Memory DDR4-2400, 16 Banks, 4 bank groups, 1 Rank, 1 Channel, Row size = 8192 bytes, $t_{RCD}$ = 13.5 ns, $t_{RP}$ = 13.5 ns , $t_{RC}$ = 13.5 ns, Open Row policy, Row Timeout = 100 ns
+5.2.2 Comparison Points for Covert-Channel Attacks
+To showcase the effectiveness of IMPACT, we evaluate the leakage throughput of five different covert-channel attacks.
+(i) DRAMA-clflush: The row buffer-based covert-channel attack proposed in [98] which uses clflush instructions [121] to send memory requests directly to main memory.
 clflush only probes the LLC to flush the cache line.
-(ii) DRAMA-Eviction: The row buffer-based covert channel attack proposed in [68] which uses cache eviction sets to send memory requests directly to main memory.
-(iii) Streamline: State-of-the-art cache covert channel proposed in [115] which enables a flushless communication.
-(iv) DMA Engine: A row buffer-based attack that uses the DMA engine to bypass the caches.
-We assume a powerful attacker can infer the DRAM access time directly from the user space by ignoring context switch and OS-related latencies.
-(v) PnM-OffChip: A PnM-based attack in a PnM architecture with locality monitor and a state-of-the-art off-chip predictor [116].
-If the data is predicted to have high locality, the computation is performed by the CPU instead of the near-memory compute units, leading to reduced attack throughput.
-(vi) IMPACT-PnM (§4.1): An attack that bypasses the locality monitor and offloads PEI instructions directly to compute units near DRAM banks in order to establish a covert channel.
-(vii) IMPACT-PuM (§4.2): An attack that leverages RowClone to offload bulk-copy operations directly to main memory and establish a covert channel.
-For a fair comparison against DRAMA and Streamline, we showcase the upper bound of the communication throughput achieved by each attack.
-To calculate their throughput, we use our simulation infrastructure to extract parameters such as the LLC hit latency, average LLC miss latency, cache lookup latency, cache hit/miss ratio, and feed them in an analytical model that calculates the maximum throughput of the attack.
-We validate the model against the real system throughput of DRAMA and Streamline (e.g., Streamline [115] reports a maximum communication throughput of 1.8 Mb/s and we model the upper bound as 2.7 Mb/s for the smallest LLC).
-Noise Sources in Simulation Setup and Error Rate.
-We simulate hardware prefetchers and page table walkers to induce noise in the simulated system.
-We measure the throughput of each attack only based on the successfully leaked data (i.e., incorporating the error rate in our measurements).
-Table 2: Simulation Configurations
-CPU 4-core, 4-way, OoO x86 core, 2.6GHz
-MMU L1 DTLB (4KB): 64-entry, 4-way assoc, 1-cycle access, L1 DTLB (2MB): 32-entry, 4-way assoc, 1-cycle access, L2 TLB: 1536-entry, 12-way assoc, 12-cycle access
-L1 Cache I-cache: 32 KB; L1D Cache: 32 KB, 8-way assoc, 4-cycle access latency, LRU, IP-stride prefetcher [117]
-L2 Cache 1 MB, 16-way assoc, 12-cycle latency, SRRIP [118], Streamer [119]
-L3 Cache 2 MB/core, 16-way assoc, 32-cycle latency, SRRIP [118]
-Main Memory DDR4-2400, 16 Banks, 4 Ranks, 1 Channel, Row size = 8192 bytes, $t_{RCD}$ = 13.5 ns, $t_{RP}$ = 13.5 ns , $t_{RC}$ = 13.5 ns, Open Row policy, Row Timeout = 100 ns
+(ii) DRAMA-Eviction: The row buffer-based covert channel attack proposed in [98], which uses cache eviction sets [122] to send memory requests directly to main memory.
+(iii) DMA Engine: A row buffer-based attack that uses the DMA engine [123, 124] to bypass the cache hierarchy.
+(iv) IMPACT-PnM (§4.1): An attack that bypasses the locality monitor and offloads PEI instructions directly to compute units near DRAM banks in order to establish a covert channel.
+(v) IMPACT-PuM (§4.2): An attack that leverages RowClone to offload bulk-copy operations directly to main memory and establish a covert channel.
+5.2.3 Noise Sources in Simulation Setup and Error Rate
+We simulate hardware prefetchers [167, 165] and page table walkers [121] to induce noise in the simulated system.
+We measure the throughput of each attack only based on the successfully leaked data (i.e., incorporating error rate in our measurements).
+6 Evaluation
 6.1 Validation of Proof-of-Concept
-We first demonstrate the PoC results and how we successfully mount the attack in the simulated system.
+We first demonstrate proof-of-concept attacks for IMPACT-PnM and IMPACT-PuM.
+We send 16-bit messages with a batch size of 16 (i.e., using 16 banks in parallel).
 Fig.
-7 shows the latency (in cycles) measured by the receiver after executing (a) a PEI instruction and (b) a RowClone operation, one for each of the 16 banks assuming the message is 16-bits large, in IMPACT-PnM and IMPACT-PuM, respectively.
-We observe that the receiver can successfully distinguish whether the sender causes contention in the bank in both IMPACT-PnM and IMPACT-PuM.
-The receiver can decode the complete message using a threshold value of 150 cycles as follows.
-If the latency is less than 150 cycles (i.e., row buffer hit), the receiver decodes a bit as logic-0, otherwise (i.e., row buffer conflict) it decodes the bit as logic-1.
-Figure 7: Latency (in cycles) measured by the receiver to execute (a) a PEI instruction and (b) a RowClone instruction for each bank in order to decode a 16-bit message.
-5.3 Covert-Channel Attack Throughput Analysis
+8 shows the latency (in cycles) measured by the receiver during transmission of example messages.
+The y-axis shows the measured latency for each (a) PEI and (b) RowClone operation for IMPACT-PnM and IMPACT-PuM, respectively.
+The x-axis shows the transmitted bit values where each bit is transmitted using a different bank.
+We make two key observations.
+First, for both attacks, the receiver successfully determines the transmitted bit values by detecting the row buffer conflicts caused by the sender.
+Second, the receiver decodes the complete message using a threshold value of 150 cycles.
+If the measured latency is less than 150 cycles (i.e., row buffer hit), the receiver decodes a bit as logic-0; otherwise (i.e., row buffer conflict), it decodes the bit as logic-1.
+Figure 8: Latency (in cycles) measured by the receiver to execute (a) a PEI and (b) a RowClone operation for each bank to decode a 16-bit message.
+6.2 Covert Channel Throughput Analysis
 Fig.
-8 shows the throughput achieved by different covert-channel attacks when one sender and one receiver thread run concurrently across systems with increasing LLC sizes.
-We make five key observations.
+9 shows the throughput achieved by IMPACT-PnM, IMPACT-PuM and state-of-the-art main memory-based covert-channel attacks in modern systems with increasing LLC sizes.
+The x-axis shows the increasing LLC size (from 1 MB to 128 MB), and the y-axis shows the leakage throughput of the covert channels.
+For all attacks, one sender and one receiver process run concurrently and transmit a fixed-length message.
+We make four key observations.
 First, IMPACT-PnM and IMPACT-PuM achieve significantly higher throughput than all other attack vectors.
-IMPACT-PnM and IMPACT-PuM achieve communication throughputs of 12.87 Mb/s and 14.16 Mb/s, respectively, irrespective of the cache size.
-They support up to 4.91$\times$ and 5.41$\times$ of the throughput achieved by the state-of-the-art main memory-based attack, DRAMA-clflush.
-Second, IMPACT-PuM achieves 10% higher throughput than that of IMPACT-PnM since the sender can issue all RowClone requests in parallel, whereas in IMPACT-PnM, the sender has to send the PiM instructions in a sequential manner.
-Third, as the LLC size increases, the throughput of DRAMA-eviction, DRAMA-clflush, and Streamline decreases significantly.
-This is because the larger the LLC size, the higher the latency it takes to access the LLC.
-Fourth, the DMA engine attack achieves a throughput of 5.27 Mb/s irrespective of the cache size but is almost 2.4$\times$ slower than IMPACT-PnM due to its OS-related overheads (e.g., context switch overhead, OS instructions).
-Fifth, the PnM-OffChip attack achieves a maximum throughput of 12.64 Mb/s.
-As the LLC size increases, the throughput is reduced to 10.64 Mb/s because the off-chip predictor decides to cache more data when the LLC is large.
-We conclude that IMPACT achieves significantly higher throughput than the existing state-of-the-art main memory-, cache- and DMA engine-based covert-channel attacks.
-Figure 8: Throughput across increasing LLC sizes.
+IMPACT-PnM and IMPACT-PuM provide communication throughputs of 8.2 Mb/s and 14.8 Mb/s, respectively, irrespective of the cache size.
+They support up to 3.6$\times$ and 6.5$\times$ of the throughput achieved by the state-of-the-art main memory-based attack, DRAMA-clflush.
+Second, IMPACT-PuM provides 82% higher throughput than IMPACT-PnM.
+This is because the sender in IMPACT-PuM can transmit a message with lower latency using fewer operations (§4.2).
+Third, as the LLC size increases, the throughputs of DRAMA-eviction and DRAMA-clflush decrease significantly due to the increasing LLC access latency.
+Fourth, the DMA engine attack provides a throughput of 0.81 Mb/s irrespective of the LLC size, which is 10.1$\times$ slower than IMPACT-PnM due to its OS-related overheads (e.g., context switch overhead, OS instructions).
+We conclude that IMPACT-PnM and IMPACT-PuM achieve significantly higher throughput than the existing state-of-the-art main memory-based covert-channel attacks.
+Figure 9: Leakage throughput of IMPACT-PnM, IMPACT-PuM, and state-of-the-art main memory-based covert channel attacks.
 Breakdown of Throughput Improvements.
+We analyze the throughput breakdown of IMPACT-PnM and IMPACT-PuM by calculating the time spent on sender and receiver routines to transmit a fixed-length message.
 Fig.
-9 shows the time (in CPU cycles) it takes for the sender to send the message and for the receiver to read the message in IMPACT-PnM and IMPACT-PuM.
-We observe that in IMPACT-PuM, the sender can send the message 14$\times$ faster than the sender in IMPACT-PuM since it can issue all RowClone requests in parallel.
-In IMPACT-PnM, even though the sender is significantly slower than the sender of IMPACT-PuM, the sender and the receiver overlap, at the cost of synchronization, to hide the latency of the sender.
-This is why IMPACT-PnM achieves only 10% lower throughput than IMPACT-PuM.
+10 shows the time (in CPU cycles) it takes for the sender to send the message and for the receiver to decode it in IMPACT-PnM and IMPACT-PuM.
+We make two key observations.
+First, in IMPACT-PuM, the sender routine takes 11.1$\times$ less time than the sender in IMPACT-PnM.
+This is because IMPACT-PuM can issue only one RowClone request to transmit N bits, where N is the number of banks.
+In contrast, IMPACT-PnM transmits the same message with potentially multiple PEIs.
+Second, the receiver routines of IMPACT-PnM and IMPACT-PuM spend similar amounts of time decoding the message as they both need to use distinct PiM operations to detect which DRAM banks were accessed by the sender.
+Based on these observations, we conclude that IMPACT-PuM has a higher throughput than IMPACT-PnM due to the differences in the sender routines.
 Figure 10: Time (in cycles) it takes (i) for the sender to send the message and (ii) for the receiver to read the message in IMPACT-PnM and IMPACT-PuM.
 6.3 Side-Channel Attack Throughput Analysis
+We evaluate the side-channel attack against genomic read mapping (RM) (§4.3) in a PiM-enabled system [67], sweeping the number of DRAM banks that store the hash table used in the seeding step.
 Fig.
-10 shows the throughput and the error rate of the side-channel attack against read mapping (RM) across different numbers of DRAM banks in the left and right y-axis, respectively.
-We measure the throughput of the attack based on the correct guesses and calculate the error rate based on the number of incorrect guesses.
-The exact end-to-end accuracy of identifying the properties of the sample genome depends on the chaining mechanism and the imputation algorithm [110, 111] used by the attacker, and it is out of the scope of this work.
+11 shows the throughput and the error rate of the side-channel attack in the primary y-axis and secondary y-axis, respectively.
+The x-axis shows the total number of DRAM banks in the system.
+We measure the throughput of the attack based on the correct guesses of the hash table entries accessed and calculate the error rate based on the number of incorrect guesses.
+[^3] The exact end-to-end accuracy of identifying the properties of the query genome depends on the chaining mechanism and the imputation algorithm [160, 161] used by the attacker, and it is out of the scope of this work.
+Figure 11: Leakage throughput and error rate of the genomic read mapping side-channel attack.
 We make two key observations.
-First, with a PiM-enabled DRAM with 1024 banks, the attacker can leak information about the sample genome at a throughput of 7.57 Mb/s and an error rate of $<5$%.
-Second, as the number of banks increases to 8196, the throughput of the attack decreases to 2.56 Mb/s and the error rate increases to $<15$%.
-This is because the attacker has to probe more banks, which makes the attack more prone to noise and reduces its bandwidth.
-However, at the same time, the leaked information provides more exact information about the sample genome since the number of hash table entries per bank decreases, reducing the number of potential locations of the reads in the reference genome per bank.
+First, the attacker can leak information about the query genome at a throughput of 7.57 Mb/s and an error rate of $<5$% in a system with 1024 banks.
+Second, as the number of banks increases, the throughput of the attack decreases to 2.56 Mb/s and the error rate increases to $<15$%.
+This is because the attacker has to probe more banks, making the attack more prone to noise and reducing its bandwidth.
+However, at the same time, the leaked information provides more exact information about the query genome since the number of hash table entries per bank decreases, reducing the number of potential locations of the reads in the reference genome per bank.
 For example, in a PiM-enabled DRAM device with 1024 banks, the victim activates a row with 16 hash table entries.
@@
 In contrast, in a PiM-enabled device with 2048 banks, the victim activates a row with 8 hash table entries.
-When the attacker identifies the activated bank correctly, they determine that one of the 8 hash table entries was targeted by the victim, which leaks more precise information on the sample genome.
-Figure 10: Leakage throughput and error rate of the RM side-channel attack across different numbers of DRAM banks.
+When the attacker identifies the activated bank correctly, they determine that one of the 8 hash table entries was targeted by the victim, which leaks more precise information on the query genome.
+Based on these observations, we conclude that an attacker can successfully leak query genome characteristics in a PiM-enabled system by leveraging PiM operations.
 7 Defenses Against IMPACT
-This section discusses the effectiveness of three defense mechanisms that eliminates the row buffer-based timing channel.
-Memory Partitioning (MPR).
-Bank-level memory partitioning enforces that each DRAM bank can only be accessed by one process that it is allocated to.
-In IMPACT, this limits the sender’s ability to transmit information across different DRAM banks and the receiver’s ability to observe contention created by the sender.
+We demonstrate high-throughput attack vectors that can emerge with the adoption of PiM architectures.
+Our results provide insights into new side and covert channel attack vectors that PiM systems might face, and show that these attack vectors should be considered while designing such PiM systems.
+In this section, we discuss the effectiveness of four defense mechanisms that mitigate the row buffer-based timing channel we introduce.
+7.1 Memory Partitioning (MPR)
+Bank-level memory partitioning (similarly to [168, 169]) allocates each DRAM bank to only one process and enforces that only that process accesses the DRAM bank.
+This way, none of the processes can observe latency differences due to other processes’ memory accesses.
+Thus, the attacker processes cannot build communication channels or leak information.
 However, this approach has three major drawbacks.
-First, it can lead to limiting the number of concurrently running applications due to the fixed number of DRAM banks available.
+First, it limits the number of applications that can run concurrently due to the fixed number of DRAM banks available in the system.
 Second, it can cause memory underutilization due to allocating DRAM bank-sized memory to applications.
-Third, it disables data sharing, which can significantly increase data duplication and memory underutilization, and lead to frequent swapping from the disk when the memory space is full.
-Closed Row Policy (CRP).
-Enforcing a memory controller policy that closes a DRAM row after each operation can be an effective defense against IMPACT.
+Third, it disables data sharing across processes, which can significantly increase data duplication and memory underutilization, and lead to frequent swapping from the disk when the memory space is full.
+7.2 Closed Row Policy (CRP)
+Enforcing a memory controller policy that closes a DRAM row after each operation, a closed row policy (similar to as described in [155, 170, 171]), can be an effective defense against IMPACT.
 By closing the DRAM row after each operation, the memory controller ensures that subsequent memory accesses cannot access the same row without requiring a row activation.
-Closed row policy disrupts the timing patterns exploited by the attacker in IMPACT.
-However, it also leads to significant performance overheads, even for workloads with low data locality, as every DRAM access leads to a row buffer miss.
-Constant-Time DRAM Accesses (CTD).
+The closed row policy disrupts the timing patterns exploited by IMPACT.
+However, it also leads to significant performance overheads [172, 170], even for workloads with low data locality, as every DRAM access leads to a row buffer miss.
+7.3 Constant-Time DRAM Accesses (CTD)
 Another defense mechanism is to enforce a constant-time DRAM access (CTD) policy.
@@
 This fixed amount of time is equal to the worst-case DRAM access latency.
-By enforcing CTD, the memory controller ensures that all memory accesses take the same amount of time, regardless of the memory location.
-Thus, obfuscates the timing patterns exploited by the attacker in IMPACT.
-However, this leads to significant performance overheads as every DRAM access has the worst-case latency.
-Performance Evaluation.
-We evaluate the performance of CRP and CTD compared to the baseline open-row policy for 5 multi-programmed graph applications which (i) share the input graph (i.e., access to the same DRAM bank) and (ii) experience a mix of low and high data locality (i.e., the number of last-level cache misses per kilo instructions).
-We use a 2-core system where each core runs a different instance of the same application (e.g., core 0 runs an instance of PageRank and core 1 the same instance of PageRank using the same input).
-We evaluate five workloads from GraphBIG benchmark suite [120]: Betweenness Centrality (BC; Data set size: 7.8 GB; LLC MPKI: 0.57), Breadth-First search (BFS; 7.6 GB; 38.59), Connected components (CC; 7.9 GB; 45.2), Triangle counting (TC; 8.2 GB; 5.08), PageRank (PR; 7.0 GB; 1.86).
-Fig.
-11 shows the performance overhead of CRP and CTD compared to the open-row policy.
+By enforcing CTD, the memory controller ensures that all memory accesses take the same amount of time, which eliminates the timing differences exploited by IMPACT.
+However, this leads to significant performance overheads (§7.4) as every DRAM access has the worst-case latency.
+7.4 Reducing the Attack Throughput with Adaptive Constant-Time DRAM (ACT)
+Another approach is to reduce the leakage throughput (without completely getting rid of the covert-channel attack) while incuring lower performance overheads than existing defenses.
+To understand the effectiveness of such a solution, we implement a memory controller-based countermeasure that adaptively enforces constant-time memory access latency per DRAM bank based on the row buffer contention observed.
+We call this countermeasure Adaptive Constant-Time (ACT).
+ACT counts the row buffer conflicts for each DRAM bank during an epoch and decides which latency policy (i.e., constant-time or default latency) to employ in the next one.
+It switches to the constant-time policy for a bank when there are more row buffer conflicts than a predetermined threshold in the last epoch.
+This way, in the next epoch, all memory requests being served by the bank will exhibit the same latency (i.e., the worst case latency), and the attacker always observes interference.
+While employing constant-time latency, ACT still keeps track of the row buffer conflicts.
+It switches back to the default policy only when the number of conflicts is lower than the threshold in the last epoch.
+This can reduce the performance overheads caused by CTD.
+ACT does not fundamentally solve the row buffer-based timing channel, and an attacker can still mount the attack by transmitting messages only when ACT is employing default latency and then waiting while ACT employs constant-time latency.
+However, this limits the usable epochs during execution and reduces the attack throughput.
+We implement ACT and evaluate its impact on system performance and the attack throughput.
+We evaluate three different configurations an epoch is set to 1000ns: (i) ACT-Aggressive: employing constant-time latency for the next 4000 epochs after the 1st conflict occurs in the bank, (ii) ACT-Mild: employing constant-time latency for the next 2 epochs after the 1st conflict occurs in the bank, (iii) ACT-Conservative: employing constant-time latency for the next 2 epochs after 5 conflicts occur in the bank.
+To analyze ACT’s impact on system performance, we evaluate four workloads from the GraphBIG benchmark suite [173]: Betweenness Centrality, Breadth-First search (BFS), Connected components (CC), Triangle counting (TC) and, XSBench (XS) [174].
+To analyze ACT’s impact on attack throughput, we calculate the reduction in IMPACT-PnM’s throughput.
+Figure 12 shows the performance comparison of ACT-Conservative, ACT-Mild, ACT-Aggressive, and CTD normalized to a system with no IMPACT defense.
+Figure 12: Performance comparison of Constant-Time policy and Adaptive Constant-Time defense policies over the baseline system with no IMPACT defense.
 We make two key observations.
-First, CTD leads to the highest performance degradation, 26% on average, compared to the open row policy.
-Second, CRP leads to an average performance degradation of 15% compared to the open row policy but leads to low performance overheads in workloads that do not benefit from the open row policy such as TC, CC and BFS.
-We conclude that fundamental defense mechanisms that eliminate the timing channel lead to high performance overheads, and there is a dire need for high-performance defense mechanisms.
-Figure 11: Performance overhead of Closed-Row policy and Constant-Time policy compared to Open-Row policy.
+First, ACT-Aggressive causes similar performance overhead as CTD, while reducing the throughput of IMPACT-PnM by 72% on average (not shown in the figure), in contrast to CTD, which prevents the attack completely.
+Second, the two less aggressive variants of ACT (ACT-Mild and ACT-Conservative) lead to only 10% performance overhead.
+However, they cannot reduce the throughput of IMPACT-PnM.
+This is because the attack utilizes all available banks during a batch of transmission, which increases the idle time between each transmission that uses a specific DRAM bank (i.e., the attack uses the same bank only after a time period where it issues requests to the other banks).
+This requires significantly high delay times for the mitigation to reduce the throughput.
+Our findings highlight the necessity for a new approach in defending against main-memory timing attacks.
+8 Discussion
+8.1 Other Potential PiM-Based Attack Vectors
+In this work, we show that emerging PiM architectures can enable a set of high-throughput main memory-based timing attacks based on the DRAM row buffer.
+Adopting PiM architectures can lead to other attack vectors or exacerbate existing ones, requiring further exploration and analysis.
+One such example is the impact of PiM on read disturbance vulnerability [175, 172].
+A prior PiM proposal in 2021 [82] states that the proposal may lead to increased vulnerability to RowHammer [175] attacks.
+A new work in 2025, PuDHammer [176], studies the read disturbance effects of Processing-using-DRAM (PuD) operations and shows that these operations greatly increase read disturbance vulnerabilities of DRAM modules (up to 158.6$\times$).
+Thus, it is critical to employ robust and scalable RowHammer [175] and RowPress [172] solutions to prevent read disturbance bitflips in PiM-enabled systems.
+Other potential attack vectors include 1) thermal attacks that leverage PiM operations to increase the temperature in DRAM to make DRAM unavailable and restrict its service time, and degrade reliability, and 2) timing attacks that leverage coarse-grained function offloading.
+Exploring these directions requires special attention and research, which we leave for future work.
+8.2 Applicability of IMPACT and its Defenses
+to Complex PiM Architectures
+Our observations apply to other PiM architectures as long as they satisfy the key properties that enable the attack: 1) allowing co-locating data from different applications in a DRAM bank, and 2) enabling subsequent accesses from different applications targeting the same DRAM bank.
+Complex PnM architectures (e.g., systems like Tesseract [177, 178, 179] that employ sophisticated prefetching and network functionality) can degrade the attack throughput by introducing additional noise.
+The performance overhead of an IMPACT defense heavily depends on how the defense mechanism interacts with the underlying system.
+The defense mechanisms we discuss in §7 are not specifically tailored for a given underlying system (e.g., constant-time policy).
+The new defense mechanism that we introduce in §7.4 enables the constant-time policy only for the bank under contention, reducing the performance overhead, but does not provide strong security guarantees.
+8.3 Restricting Access to PIM Operations
+A key contribution of our study is to show that unrestricted access to PiM operations enables potentially harmful high-throughput side and covert channel attack vectors.
+None of the current PiM proposals we are aware of discuss restricting access to PiM operations to ensure security.
+We hope and believe that our study will guide system designers in balancing the performance, flexibility, and security trade-offs while implementing effective access control mechanisms to PiM.
+8.4 Applicability to Future DRAM Devices
+The row buffer timing channel would persist in future DRAM-based systems, enabling IMPACT.
+With newer generations, we observe that DRAM modules consist of increasingly more banks, which increases IMPACT’s covert channel throughput as the attack would benefit from more parallelism.
+Another change is the introduction of RowHammer mitigations, PRAC [180, 181, 182] and RFM [183], in recent DDR specifications that block DRAM accesses when necessary to prevent RowHammer bitflips.
+These mitigations incur at least 350ns (at most 1400ns) latency per preventive action.
+This is significantly higher than the row buffer conflict latency and can be filtered out by the receiver.
+We believe that increasing DRAM’s complexity will increase attack vectors as manufacturers try to combat the challenges of technology scaling.
+For example, recent preprints show that PRAC and RFM introduce timing channels that can be exploited for side-channel [184] and covert-channel [184, 185] attacks.
 9 Related Work
-To the best of our knowledge, this is the first work that evaluates the timing attack vulnerabilities in PiM systems.
-Security Implications of PiM Systems.
+To our knowledge, this is the first work that evaluates timing attack vulnerabilities in PiM systems.
+9.1 Security Implications of PiM Systems
 A set of prior works [186, 187, 188] discuss specific security vulnerabilities in PiM systems.
@@
 Wang et al.
-[123] extracts neural network architectural information from an RRAM accelerator using power side channels.
+[188] extract neural network architectural information from an RRAM accelerator using power side channels.
 Ensan et al.
-[122] proposes an attack on RRAM-based PiM architectures using power/timing side channels.
+[187] propose an attack on RRAM-based PiM architectures using power/timing side channels.
 In contrast, we demonstrate how to establish high-throughput timing covert and side channels using PiM operations.
-Other Main Memory-based Timing Attack Defenses.
-We already extensively discuss defenses that fundamentally eliminate the timing channel (§6) and defenses based on cache management methods that are inapplicable.
-Another approach is to restrict the fine-grained timers’ usage to prevent attackers from measure memory requests’ latencies, which is applied in some modern processors (e.g., Apple M1 cores [124]).
+9.2 Other Timing Attacks
+We already discuss main memory-based timing attacks in §2.3.
+In addition, prior works showcase multiple cache-based timing side and covert-channel attacks (e.g., [189, 190, 191, 192, 193, 194, 195, 196, 197]).
+Among these works, Streamline [192] has the highest throughput as it avoids explicit cache flushes and utilizes a large transmission set to evict cache lines implicitly by thrashing the cache.
+IMPACT achieves a leakage throughput within the same order of magnitude of Streamline and does not require a transmission set larger than the cache size.
+9.3 Other Main Memory Timing Attack Defenses
+We already extensively discuss defenses that fundamentally eliminate the timing channel (§7).
+Another approach is to restrict the high-resolution timers’ usage to prevent attackers from measuring memory requests’ latencies, which is applied in some modern processors (e.g., Apple M1 cores [198]).
 However, these timers are used in many userspace applications (e.g., performance and power consumption monitoring, multithreading, and synchronization and locking).
@@
 Trusted execution environments (TEEs) (e.g., Intel SGX [200, 201]) can isolate applications from the rest of the system.
-However, prior works already showcase multiple side-channel attacks on TEEs leveraging the main memory [128], and other structures [129, 130, 131, 128, 132, 133, 134, 135, 136, 137].
-Enhancing System Security using PiM.
-Several prior works enhance the security by leveraging PiM.
+Prior works [202, 203, 204, 205, 206, 207, 208, 209, 210, 211] already demonstrate multiple side-channel attacks on TEEs leveraging main memory [205], cache hierarchy [202, 203, 204, 205, 206], page table entries [207, 208, 209, 205], and branch predictors [210, 211].
+9.4 Enhancing System Security using PiM
+Several prior works enhance system security by leveraging PiM.
 SecNDP [212] proposes an encryption and verification mechanism for untrusted PiM devices.
@@
 Nelson et al.
-[139] propose offloading security critical tasks to near-memory compute units to avoid storing sensitive information in caches and eliminate cache-based side-channel attacks.
+[213] propose offloading security-critical tasks to PnM units to avoid storing sensitive information in caches and eliminate cache-based side-channel attacks.
 In this work, we demonstrate and evaluate the implications of PiM on main memory-based timing attacks.
@@
 10 Conclusion
-We introduce IMPACT, a set of high-throughput main memory timing attacks that leverage the direct memory access enabled by PiM architectures.
+We introduce IMPACT, a set of high-throughput main memory timing attacks that leverage the direct memory access enabled by Processing-in-Memory (PiM) architectures.
 IMPACT exploits the shared DRAM row buffer to 1) establish covert channels to communicate with other processes at high throughput and 2) leak security-critical information of victim processes based on their memory accesses.
 IMPACT covert channels leverage parallelism and eliminate cache bypassing steps, resulting in increased throughput compared to existing main memory-based covert channels.
-IMPACT side-channel attack targets a genomics application and leaks the private characteristics of a user’s sample genome by leveraging PiM operations.
-We evaluate and discuss three fundamental defense mechanisms against IMPACT.
+IMPACT side-channel attack targets a genomics application and leaks the private information with a low error rate.
+We evaluate and discuss four potential defense mechanisms against IMPACT and eventually conclude that mitigating IMPACT incurs high performance overhead and more research is needed to find low-overhead solutions.
+Acknowledgments
+We thank the anonymous reviewers of USENIX Security’24, ISCA’24, MICRO’24, ASPLOS’25, and DSN’25 for their constructive feedback.
+We thank the SAFARI Research Group members for providing a stimulating intellectual environment.
+We acknowledge the generous gifts from our industrial partners, including Google, Huawei, Intel, and Microsoft.
+This work, along with our broader work in Processing-in-Memory and memory systems, is supported in part by the Semiconductor Research Corporation (SRC), the ETH Future Computing Laboratory (EFCL), AI Chip Center for Emerging Smart Systems (ACCESS), sponsored by InnoHK funding, Hong Kong SAR, European Union’s Horizon programme for research and innovation [101047160 - BioPIM], a Google Security and Privacy Research Award, and the Microsoft Swiss Joint Research Center.
+Appendix A Artifact Appendix
+A.1 Abstract
+This artifact provides the source code and scripts to reproduce the experiments and results presented in our DSN 2025 paper.
+It includes implementations of various covert channel attacks (PNM, DRAMA-clflush, DRAMA-evict, PUM, DMA) and a defense mechanism, along with scripts to automate the setup, execution, and result analysis.
+A.2 Artifact Checklist (Meta-information)
+  • Program: C++ programs, Python scripts, shell scripts.
+  • Compilation: GNU Make.
+  • Run-time environment: Linux (tested on Ubuntu 20.04 and 22.04), Python 3.
+  • Execution: Bash scripts, Python scripts.
+  • Metrics: Timing results, performance metrics of attacks and defenses.
+  • Output: Figures in PDF format and related data in plaintext files.
+  • How much disk space required (approximately)?: 20GB.
+  • How much time is needed to prepare workflow (approximately)?: $\sim 1$ hour.
+  • How much time is needed to complete experiments (approximately)?: 2-4 hours.
+  • Publicly available?: Yes
+  • Archived (provide DOI)?: Yes, DOI: 10.5281/zenodo.15116851
+A.3 Description
+A.3.1 How to Access
+The source code and scripts can be downloaded from Zenodo (https://zenodo.org/records/15116851) and GitHub (https://github.com/CMU-SAFARI/IMPACT).
+A.3.2 Hardware dependencies
+  • x86-64 system.
+  • At least 8GB of RAM (128GB recommended for parallel defense experiments).
+  • $>64$ cores are suggested for parallel defense experiments.
+A.3.3 Software dependencies
+  • Linux Operating System (tested on Ubuntu 20.04 and 22.04).
+  • Bash.
+  • Python (version 3.12.9)
+  • GNU Make (version 4.3).
+  • wget, tar.
+  • GCC (11.4.0), g++ (11.4.0).
+  • zlib1g-dev.
+A.3.4 Data sets
+The defense mechanism experiments may require trace files. The script includes commented-out lines to download example traces. You may need to download these or provide your own traces.
+A.4 Installation
+A.4.1 Dependency Installation
+To install all dependencies, run the dependency installation script: install_dependencies.sh.
+{mdframed}
+[backgroundcolor=darkgray, linecolor=black, linewidth=0.5pt] \$ cd IMPACT
+\$ sh install_dependencies.sh
+\$ bash
+\$ conda activate impact
+A.4.2 Conda Environment Setup
+If you would like to install Miniconda manually, follow the steps below.
+  1. 1.
+  2. 2.
+    Run the Miniconda installer.
+  3. 3.
+    Open a new terminal.
+A.5 Experiment Workflow
+To run all experiments, run the main script: artifact.sh. This script 1) builds the attacks and the simulator and 2) runs all experiments required to reproduce the key results.
+{mdframed}
+[backgroundcolor=darkgray, linecolor=black, linewidth=0.5pt] \$ sh artifact.sh
+The artifact.sh script will execute the following steps.
+A.5.1 Building the Attacks and Simulator
+  1. 1.
+    The script will navigate to the impact/covert_channel_attack/ directory.
+  2. 2.
+    It will clean and build the attack code using make clean and make all.
+  3. 3.
+    It will navigate to the simulator/sniper directory.
+  4. 4.
+    It will clean and build the simulator using make distclean and make -j\$(nproc).
+A.5.2 Running Examples and Attacks
+A.5.3 Basic Example Run
+The script includes a commented-out line to run a basic example using the Sniper simulator. Uncomment the line.
+A.5.4 Running Attacks for Figure 9
+The script executes the following attacks:
+  • PNM: python3 \$home_dir/impact/covert_channel_attack/script/run_pnm.py \$home_dir 16 \$home_dir/results_attacks/pnm
+  • DRAMA (clflush): python3 \$home_dir/impact/covert_channel_attack/script/run_drama.py \$home_dir 16 \$home_dir/results_attacks/drama
+  • DRAMA (eviction sets): python3 \$home_dir/impact/covert_channel_attack/script/run_drama_ev.py \$home_dir 16 \$home_dir/results_attacks/drama_ev
+  • PUM: python3 \$home_dir/impact/covert_channel_attack/script/run_pum.py \$home_dir 16 \$home_dir/results_attacks/pum
+  • DMA: python3 \$home_dir/impact/covert_channel_attack/script/run_dma.py \$home_dir 16 \$home_dir/results_attacks/dma
+A.5.5 Running Attacks for Figure 8 and 10 (PNM POC)
+python3 \$home_dir/impact/covert_channel_attack/script/run_pnm_poc.py \$home_dir 16 \$home_dir/results_attacks/pnm_poc
+A.5.6 Running Attacks for Figure 8 and 10 (PUM POC)
+python3 \$home_dir/impact/covert_channel_attack/script/run_pum_poc.py \$home_dir 16 \$home_dir/results_attacks/pum_poc
+A.5.7 Running Defense Mechanism Experiments for Figure 12
+  • Download traces (if needed).
+  • Create the jobfile: python3 \$home_dir/scripts/create_jobfile_impact.py \$home_dir/simulator/sniper \$home_dir/scripts/defenses.jobfile
+  • Execute the jobfile: sh \$home_dir/scripts/defenses.jobfile
+A.6 Plotting the Results
+Once the experiments are completed, the reader can plot the figures using the plotting scripts we provide.
+To reproduce the figures, 1) install plotting dependencies:
+{mdframed}
+[backgroundcolor=darkgray, linecolor=black, linewidth=0.5pt] \$ pip install numpy matplotlib pandas
+2) run the plotting script: plot_figures.sh
+{mdframed}
+[backgroundcolor=darkgray, linecolor=black, linewidth=0.5pt] \$ sh plot_figures.sh
+This script will create Figures 8, 9, 10, and 12 under the figures/ directory.
+A.7 Evaluation & Expected Results
+A.7.1 Post-Execution Steps
+Review Figures 8, 9, 10, and 12 in the figures/ directories.
+Note that the attack codes are compiled during the setup, and the results presented in Figures 8, 9, and 10 might vary slightly based on your system configurations (e.g., compiler version).
+A.7.2 Troubleshooting
+Refer to the troubleshooting section in the README.md for potential issues and solutions.
+A.8 Methodology
+Submission, reviewing and badging methodology: